### v1

# Tackling Long-Horizon Tasks with Model-based Offline Reinforcement Learning

Abstract
Model-based offline reinforcement learning (RL) is a compelling approach that addresses the challenge of learning from limited, static data by generating imaginary trajectories using learned models. However, it falls short in solving long-horizon tasks due to high bias in value estimation from model rollouts. In this paper, we introduce a novel model-based offline RL method, Lower Expectile Q-learning (LEQ), which enhances long-horizon task performance by mitigating the high bias in model-based value estimation via expectile regression of $\lambda$-returns. Our empirical results show that LEQ significantly outperforms previous model-based offline RL methods on long-horizon tasks, such as the D4RL AntMaze tasks, matching or surpassing the performance of model-free approaches. Our experiments demonstrate that expectile regression, $\lambda$-returns, and critic training on offline data are all crucial for addressing long-horizon tasks. Additionally, LEQ achieves performance comparable to the state-of-the-art model-based and model-free offline RL methods on the NeoRL benchmark and the D4RL MuJoCo Gym tasks.
1 Introduction
One of the major challenges in offline reinforcement learning (RL) is the overestimation of values for out-of-distribution actions due to the lack of environment interactions [21, 19]. Model-based offline RL addresses this issue by generating additional (imaginary) training data using a learned model, thereby augmenting the given offline data with synthetic experiences that cover out-of-distribution states and actions [34, 16, 35, 2, 30]. While these approaches have demonstrated strong performance in simple, short-horizon tasks, they struggle with noisy model predictions and value estimations, particularly in long-horizon tasks [23]. This challenge is evident in their poor performances (i.e. near zero) on the D4RL AntMaze tasks [6, 15].
Typical model-based offline RL methods alleviate the inaccurate value estimation problem (mostly overestimation) by penalizing Q-values estimated from model rollouts with uncertainties in model predictions [34, 16] or value predictions [30, 14]. While these penalization terms prevent a policy from exploiting erroneous value estimations, the policy now does not maximize the true value, but maximizes the value penalized by heuristically estimated uncertainties, which can lead to sub-optimal behaviors. This is especially problematic in long-horizon, sparse-reward tasks, where Q-values are similar across nearby states [23].
Another way to reduce bias in value estimates is by using multi-step returns [31, 12]. CBOP [14] constructs an explicit distribution of multi-step Q-values from thousands of model rollouts and uses this value as a target for training the Q-function. However, CBOP is computationally expensive for estimating a target value and uses multi-step returns solely for Q-learning, which provides insufficient learning signals for obtaining long-horizon behaviors.
To tackle long-horizon tasks with model-based offline RL, we introduce a simple yet effective model-based offline RL algorithm, Lower Expectile Q-learning (LEQ). As illustrated in Figure 1, LEQ uses expectile regression with a small $\tau$ for both policy and Q-function training, providing an efficient and elegant way to achieve conservative Q-value estimates. Moreover, to better handle long-horizon tasks, we propose to optimize a policy and Q-function using $\lambda$-returns (i.e. TD($\lambda$) targets) of long ($15$-step) model rollouts, allowing the policy to directly learn from low-bias multi-step returns [28].
The experiments on the D4RL AntMaze and MuJoCo Gym tasks [6], as well as the NeoRL benchmark [26], demonstrate that our proposed conservative policy optimization with $\lambda$-return and critic training on offline data significantly improves offline RL policies in long-horizon tasks while achieving comparable performance in short-horizon, dense-reward tasks. Specifically, to the best of our knowledge, LEQ is the first model-based offline RL algorithm capable of matching or outperforming the performance of model-free offline RL algorithms on the long-horizon AntMaze tasks [6, 15].
Figure 1: Lower Expectile Q-learning (LEQ). (left) In offline model-based RL, an agent can generate imaginary trajectories using a world model. (right) For conservative Q-evaluation of the policy, LEQ learns the lower expectile of the target $Q$-distribution from a sampled individual rollout $\mathcal{T}_{i}$, without estimating the entire Q-distribution with exhaustive rollouts.
2 Related Work
Offline RL [21] aims to solve a reinforcement learning problem only with pre-collected datasets, better than behavioral cloning policies [25]. One can simply apply off-policy RL algorithms on top of the fixed dataset. However, off-policy RL methods suffer from the overestimation of Q-values for actions unseen in the offline dataset [8, 18, 19], since an overestimated value function cannot get corrected through online environment interactions in offline RL.
Model-free offline RL algorithms have addressed this value overestimation problem on out-of-distribution actions by (1) regularizing a policy to only output actions in the offline data [24, 17, 7] or (2) adopting a conservative value estimation for executing actions different from the dataset [19, 1]. Despite their strong performances on the standard offline RL benchmarks, model-free offline RL policies tend to be constrained to the support of the data (i.e. state-action pairs in the offline dataset), which may lead to limited generalization capability.
Model-based offline RL approaches have tried to overcome this limitation by suggesting a better use of the limited offline data – learning a world model and generating imaginary data with the learned model that covers out-of-distribution actions. Similar to Dyna-style online model-based RL [32, 9, 10, 11], an offline model-based RL policy can be trained on both offline data and model rollouts. But, again, learned models may be inaccurate on states and actions outside the data support, making a policy easily exploit the learned models.
Recent model-based offline RL algorithms have adopted the conservatism idea from model-free offline RL, penalizing policies incurring (1) uncertain transition dynamics [34, 16, 35] or (2) uncertain value estimation [30, 14]. This conservative use of model-generated data enables model-based offline RL to outperform model-free offline RL in widely used offline RL benchmarks [30]. However, uncertainty estimation is difficult and often inaccurate [35]. Instead of relying on such heuristic [34, 16, 30] or expensive [14] uncertainty estimation, we propose to learn a conservative value function via expectile regression with a small $\tau$, which is simple, efficient, yet effective.
3 Preliminaries
Problem setup.
[…]
In this paper, we consider the offline RL setup [21], where a policy $\pi$ is trained with a fixed given offline dataset, $\mathcal{D}_{\text{env}}=\{\tau_{1},\tau_{2},...,\tau_{N}\}$, without any additional online interactions.
Model-based offline RL.
As an offline RL policy is trained from a fixed dataset, one of the major challenges in offline RL is the limited data support; thus, lack of generalization to out-of-distribution states and actions. Model-based offline RL [16, 34, 35, 27, 30, 14] tackles this problem by augmenting the training data with imaginary training data (i.e. model rollouts) generated from the learned transition dynamics and reward model, $p_{\psi}(\mathbf{s}_{t+1},r\mid\mathbf{s}_{t},\mathbf{a}_{t})$.
The typical process of model-based offline RL is as follows: (1) pretrain a model (or an ensemble of models) and an initial policy from the offline data, (2) generate short imaginary rollouts $\{\tau\}$ using the pretrained model and add them to the training dataset $\mathcal{D}_{\text{model}}\leftarrow\mathcal{D}_{\text{model}}\cup\{\tau\}$, (3) perform an offline RL algorithm on the augmented dataset $\mathcal{D}_{\text{model}}\cup\mathcal{D}_{\text{env}}$, and repeat (2) and (3).
Expectile regression.
Expectile is a generalization of the expectation of a distribution $X$. While the expectation of $X$, $\mathbb{E}[X]$, can be viewed as a minimizer of the least-square objective, $L_{2}(y)=\mathbb{E}_{x\sim X}[(y-x)^{2}]$, $\tau$-expectile of $X$, $\mathbb{E}^{\tau}[X]$, can be defined as a minimizer of the asymmetric least-square objective:
$L_{2}^{\tau}(y)=\mathbb{E}_{x\sim X}\left[\lvert\tau-\mathbbm{1}(y>x)\rvert\cdot(y-x)^{2}\right],$ (1)
where $\lvert\tau-\mathbbm{1}(y>x)\rvert$ is an asymmetric weighting of $L_{2}$ and $0\leq\tau\leq 1$.
We refer a $\tau$-expectile with $\tau<0.5$ as a lower expectile of $X$. When $\tau<0.5$, the objective assigns a high weight $1-\tau$ for smaller $x$ and a low weight $\tau$ for bigger $x$. Thus, minimizing the objective with $\tau<0.5$ leads to a conservative statistical estimate compared to the expectation.
4 Approach
The primary limitation for model-based offline RL in solving long-horizon tasks is inherent errors in a world model and critic outside the offline data support. Conservative value estimation can effectively handle such (falsely optimistic) errors. Prior approaches estimate conservative values through diverse uncertainty penalties; but they are either unreliable [35] or computationally expensive [14].
In this paper, we introduce Lower Expectile Q-learning (LEQ), an efficient model-based offline RL method that achieves conservative value estimation via expectile regression of Q-values with lower expectiles when learning from model-generated data (Section 4.1). Additionally, we address the noisy value estimation problem in long-horizon tasks [23] using $\lambda$-returns on $10$-step imaginary rollouts (Section 4.2). Finally, we train a deterministic policy conservatively by maximizing the lower expectile of $\lambda$-returns (Section 4.3). The overview of LEQ is described in Algorithm 1.
Algorithm 1 LEQ: Lower Expectile Q-learning with $\lambda$-returns
0:  Offline dataset $\mathcal{D}_{\text{env}}$, expectile $\tau\leq 0.5$, imagination length $H$, dataset expansion length $R$.
1:  Initialize world models $\{p_{\psi_{1}},\cdots,p_{\psi_{M}}\}$, policy $\pi_{\theta}$, and Q-function $Q_{\phi}$
2:  Pretrain $\{p_{\psi_{1}},\cdots,p_{\psi_{M}}\}$ on $\mathcal{D}_{\text{env}}$triangleright $\mathcal{L}_{\text{wm}}(\psi)=-\mathbb{E}_{(\mathbf{s},\mathbf{a},r,\mathbf{s}^{\prime})\in\mathcal{D}_{\text{env}}}\log p_{\psi}(\mathbf{s}^{\prime},r\mid\mathbf{s},\mathbf{a})$
3:  Pretrain $\pi_{\theta}$ and $Q_{\phi}$ on $\mathcal{D}_{\text{env}}$triangleright using BC for $\pi_{\theta}$ and FQE [20] for $Q_{\phi}$
4:  $\mathcal{D}_{\text{model}}\leftarrow\emptyset$
5:  while not converged do
6:     // Expand dataset using model rollouts
7:     $\mathbf{s}_{0}\sim\mathcal{D}_{\text{env}}$triangleright start dataset expansion from any state in $\mathcal{D}_{\text{env}}$
8:     for $t=0,\ldots,R-1$ do
9:        $\mathcal{D}_{\text{model}}\leftarrow\mathcal{D}_{\text{model}}\cup\{\mathbf{s}_{t}\}$
10:        $\mathbf{a}_{t}=\pi_{\theta}(\mathbf{s}_{t})$
11:        $\mathbf{s}_{t+1},r_{t}\sim p_{\psi}(\cdot\mid\mathbf{s}_{t},\mathbf{a}_{t})$, where $p_{\psi}\sim\{p_{\psi_{1}},\cdots,p_{\psi_{M}}\}$
12:     // Generate imaginary data, $\tau=\{(\mathbf{s}_{0},\mathbf{a}_{0},r_{0},\cdots,\mathbf{s}_{H-1},\mathbf{a}_{H-1},r_{H-1},\mathbf{s}_{H})_{i}\}$
13:     $\mathbf{s}_{0}\sim\mathcal{D}_{\text{model}}$triangleright start imaginary rollout from any state in $\mathcal{D}_{\text{model}}$
14:     for $t=0,\ldots,H-1$ do
15:        $\mathbf{a}_{t}=\pi_{\theta}(\mathbf{s}_{t})$
16:        $\mathbf{s}_{t+1},r_{t}\sim p_{\psi}(\cdot\mid\mathbf{s}_{t},\mathbf{a}_{t})$, where $p_{\psi}\sim\{p_{\psi_{1}},\cdots,p_{\psi_{M}}\}$
17:     // Update critic using both offline and model-generated data
18:     Update critic $Q_{\phi}$ to minimize $\mathcal{L}^{\lambda}_{Q}(\phi)$ in Eq. 7 using $\tau$ and $\{\mathbf{s},\mathbf{a},r,\mathbf{s}^{\prime}\}\sim\mathcal{D}_{\text{env}}$
19:     // Update actor using only model-generated data
20:     Update actor $\pi_{\theta}$ to minimize ^$\hat{\mathcal{L}}^{\lambda}_{\pi}(\theta)$ in Eq. 12 using $\tau$
4.1 Lower expectile Q-learning
Most offline RL algorithms primarily focus on learning a conservative value function for out-of-distribution actions. In this paper, we propose Lower Expectile Q-learning (LEQ), which learns a conservative Q-function via expectile regression with small $\tau$, avoiding unreliable uncertainty estimation and exhaustive Q-value estimation.
As illustrated in Figure 1, the target value for $Q_{\phi}(\mathbf{s},\mathbf{a})$, where $\mathbf{a}\leftarrow\pi_{\theta}(\mathbf{s})$, can be estimated by rolling out an ensemble of world models and averaging $r(\mathbf{s},\mathbf{a})+\gamma Q_{\phi}(\mathbf{s}^{\prime},\mathbf{a}^{\prime})$ over all possible $\mathbf{s}^{\prime}$:
$\hat{y}_{\text{model}}=\mathbb{E}_{\psi\sim\{\psi_{1},...,\psi_{M}\}}\mathbb{E}_{(\mathbf{s}^{\prime},r)\sim p_{\psi}(\cdot\mid\mathbf{s},\mathbf{a})}\left[r+\gamma Q_{\phi}(\mathbf{s}^{\prime},\pi_{\theta}(\mathbf{s}^{\prime}))\right].$ (2)
This target value has three error sources: the predicted future state and reward $\mathbf{s}^{\prime},r\sim p_{\psi}(\cdot\mid\mathbf{s},\mathbf{a})$ and future Q-value $Q_{\phi}(\mathbf{s}^{\prime},\pi_{\theta}(\mathbf{s}^{\prime}))$. Thus, the target value computed from model-generate data, $\hat{y}_{\text{model}}$, is more prone to overestimation than that of the original target Q-value, $\hat{y}_{\text{env}}$, computed from $(\mathbf{s},\mathbf{a},r,\mathbf{s}^{\prime})\sim D_{\text{env}}$:
$\hat{y}_{\text{env}}=r+\gamma Q_{\phi}(\mathbf{s}^{\prime},\pi_{\theta}(\mathbf{s}^{\prime})).$ (3)
To mitigate the overestimation problem in estimating the true Q-value from $H$-step inaccurate world model rollouts, we propose to use expectile regression on target Q-value estimation with small $\tau$; as illustrated in Figure 1, expectile regression with small $\tau$ tends to choose the target Q-value that is lower than the expectation, effectively providing a conservative estimate of target Q-value. Another advantage of using expectile regression is that we do not have to exhaustively evaluate Q-values to get $\tau$-expectiles as Jeong et al. [14]; instead, we can do conservative estimation using sampling:
$L_{Q,\text{model}}(\phi)=\mathbb{E}_{\mathbf{s}_{0}\in\mathcal{D}_{\text{model}},\tau\sim p_{\psi},\pi_{\theta}}\left[\frac{1}{H}\sum_{t=0}^{H}L_{2}^{\tau}(Q_{\phi}(\mathbf{s}_{t},\pi_{\theta}(\mathbf{s}_{t}))-\hat{y}_{\text{model}})\right].$ (4)
In addition, the Q-function is also trained on the offline data $\mathcal{D}_{\text{env}}$ with the standard Bellman update:
$\mathcal{L}_{Q,\text{env}}(\phi)=\mathbb{E}_{(\mathbf{s},\mathbf{a},r,\mathbf{s}^{\prime})\in\mathcal{D}_{\text{env}}}\left[\frac{1}{2}(Q_{\phi}(\mathbf{s},\mathbf{a})-\hat{y}_{\text{env}})^{2}\right].$ (5)
To stabilize training of the Q-function, we adopt EMA regularization [11], which prevents drastic change of Q-values by regularizing the difference between the Q-predictions and ones from the exponential moving average:
$\mathcal{L}_{Q,\text{EMA}}(\phi)=\mathbb{E}_{(\mathbf{s},\mathbf{a})\in\mathcal{D}_{\text{env}}}\left[(Q_{\phi}(\mathbf{s},\mathbf{a})-Q_{\bar{\phi}}(\mathbf{s},\mathbf{a}))^{2}\right],$ (6)
where $\bar{\phi}$ is an exponential moving average of $\phi$. Note that by using EMA regularization, we do not use the target Q-network for Equations 2 and 3.
Finally, by combining the three losses above, we define the critic loss as follows:
$\mathcal{L}_{Q}(\phi)=\beta\mathcal{L}_{Q,\text{model}}(\phi)+(1-\beta)\mathcal{L}_{Q,\text{env}}(\phi)+\beta_{\text{EMA}}\mathcal{L}_{Q,\text{EMA}}(\phi).$ (7)
4.2 Lower expectile Q-learning with $\lambda$-return
To further improve LEQ for long-horizon tasks, we use $\lambda$-return instead of $1$-step return for Q-learning. $\lambda$-return allows a Q-function and policy to learn from low-bias multi-step returns [28]. Reducing bias in value estimation with $\lambda$-return is especially important on long-horizon tasks where values for nearby states are similar to each other, as illustrated in Figure 2.
We first define $\lambda$-return of a trajectory $\tau$ in timestep $t$, $Q_{t}^{\lambda}(\tau)$, using $N$-step return, $G_{t:t+N}(\tau)$: [^2] Our $\lambda$-return is slightly different from [31, 11] that puts a high weight to the last $N$-step return, $G_{t:H}(\tau)$.
$G_{t:t+N}(\tau)=\sum_{i=0}^{N-1}\gamma^{i}r(\mathbf{s}_{t+i},\mathbf{a}_{t+i})+\gamma^{N}Q_{\phi}(\mathbf{s}_{t+N},\mathbf{a}_{t+N}),$ (8)
$Q_{t}^{\lambda}(\tau)=\frac{1-\lambda}{1-\lambda^{H-t-1}}\sum_{i=1}^{H-t}\lambda^{i-1}G_{t:t+i}(\tau).$ (9)
Then, we can rewrite the Q-learning loss in Equation 4 with $1$-step return to the one with $\lambda$-return:
$\mathcal{L}^{\lambda}_{Q,\text{model}}(\phi)=\mathbb{E}_{\mathbf{s}_{0}\in\mathcal{D}_{\text{model}},\tau\sim p_{\psi},\pi_{\theta}}\left[\sum_{t=0}^{H-1}L_{2}^{\tau}(Q_{\phi}(\mathbf{s}_{t},\pi_{\theta}(\mathbf{s}_{t}))-Q^{\lambda}_{t}(\tau))\right].$ (10)
Figure 2: $\lambda$-return with noisy critic. For long-horizon tasks, value estimates (blue for high values and white for low values) are likely to be flat for the states far from the reward signal shown with the yellow star. Since Q-values are small and similar to each other in this region, Q-values are easily dominated by noise, leading to a wrong learning signal for the policy, as illustrated in the middle row. $\lambda$-return can effectively reduce the critic noise by considering both short-term and long-term value estimates, leading to a better learning signal for the policy.
4.3 Lower expectile policy learning with $\lambda$-return
For policy optimization, we can use a deterministic policy $\mathbf{a}=\pi_{\theta}(\mathbf{s})$ and update the policy using the deterministic policy gradients similar to DDPG [22]. [^3] LEQ also works with a stochastic policy; but, a deterministic policy is easier to train in an offline setup. Instead of maximizing the immediate Q-value, $Q_{\phi}(\mathbf{s},\mathbf{a})$, we propose to directly maximize the lower expectile of $\lambda$-return, which is a more accurate learning target for a policy, analogous to the conservative critic target in Section 4.2:
$\mathcal{L}_{\pi}^{\lambda}(\theta)=-\mathbb{E}_{\mathbf{s}_{0}\in\mathcal{D}_{\text{model}},\tau\sim p_{\psi},\pi_{\theta}}\left[\sum_{t=0}^{H}\mathbb{E}^{\tau}_{\tau\sim p_{\psi},\pi_{\theta}}\left[Q_{t}^{\lambda}(\tau)\right]\right].$ (11)
However, due to the expectile term in Equation 11, computing the gradient of $\mathcal{L}_{\pi}^{\lambda}(\theta)$ is not trivial. To estimate this gradient, we propose a differentiable surrogate loss, approximating $\mathbb{E}^{\tau}_{\tau\sim p_{\psi},\pi_{\theta}}\left[Q^{\lambda}_{t}(\tau)\right]$ in Equation 11 with $Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})$:
^$\hat{\mathcal{L}}_{\pi}^{\lambda}(\theta)=-\mathbb{E}_{\mathbf{s}_{0}\in\mathcal{D}_{\text{model}},\tau\sim p_{\psi},\pi_{\theta}}\left[\sum_{t=0}^{H}\lvert\tau-\mathbbm{1}\left(Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})>Q^{\lambda}_{t}(\tau)\right)\rvert\cdot Q_{t}^{\lambda}(\tau)\right].$ (12)
Intuitively, this surrogate loss sets a higher weight ($1-\tau$) on a conservative $\lambda$-return estimation (i.e. $Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})>Q^{\lambda}_{t}(\tau)$), encouraging a policy to optimize for this conservative $\lambda$-return. On the other hand, an optimistic $\lambda$-return estimation (i.e. $Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})<Q^{\lambda}_{t}(\tau)$) has a less impact to the policy with a smaller weight ($\tau$). Thus, optimizing this surrogate loss leads to the policy maximizing the lower expectile of $\lambda$-return. We provide a proof in Appendix B saying that the proposed surrogate loss is a better approximation of Equation 11 than directly maximizing Q-values, $Q_{\phi}(\mathbf{s},\mathbf{a})$.
4.4 Expanding dataset with model rollouts
One of the problem of offline RL is that data distribution is limited to the offline dataset $\mathcal{D}_{\text{env}}$. To tackle this problem, we can simulate the current policy inside the model and use the generated trajectories to expand the dataset, similar to prior works [35, 30]. However, the state coverage will be identical when the policy converges, which might lead to catastrophic forgetting.
To prevent this issue, we expand the dataset using the model rollouts from a noisy exploration policy. Specifically, we execute the exploration policy $\pi_{\exp}(\cdot\mid\mathbf{s})$, which simply adds noise $\epsilon\sim N(0,\sigma_{\text{exp}}^{2})$ to the current policy $\pi_{\theta}(\mathbf{s})$, and generate a trajectory of length $R$ ($R=5$ in this paper). We refer this expanded dataset as $\mathcal{D}_{\text{model}}$. Note that we do not use off-policy actions and rewards for critic and policy updates; instead, we generate $H$-step model rollouts starting from $\mathbf{s}\sim\mathcal{D}_{\text{model}}$ and use them for training the policy and Q-function. Thus, we need to store only the states from the rollouts.
5 Experiments
In this paper, we propose a novel model-based offline RL method with simple and efficient yet accurate conservative value estimation. Through our experiments, we aim to answer the following questions: (1) Can LEQ solve long-horizon tasks? (2) How does LEQ perform in widely used offline RL benchmarks? (3) Which component enables model-based offline RL to learn the AntMaze tasks?
(a) Hopper
(b) Walker2d
(c) HalfCheetah
Figure 3: Locomotion tasks.
(d) Umaze
(e) Medium
(f) Large
(g) Ultra
Figure 4: AntMaze tasks.
5.1 Tasks
To show the strength of LEQ in solving long-horizon tasks, we use the AntMaze tasks, which aims to navigate a $8$-DOF ant robot to the desired goal position, as shown in Figure 4. Specifically, we use umaze, medium, large datasets from D4RL [6], and ultra dataset from Jiang et al. [15]. Moreover, we evaluate our method on MuJoCo locomotion tasks (Figure 4) with dense rewards with D4RL [6] and NeoRL [26] datasets. Please refer to Appendix A for more experimental details.
5.2 Compared offline RL algorithms
We compare the performance of LEQ with the state-of-the-art offline RL algorithms. Please note that LEQ uses the same hyperparameters across all tasks, except the expectile parameter, $\tau$.
Model-free offline RL.
We consider behavioral cloning (BC[25]; TD3+BC [7], which combines BC loss to TD3; CQL [19], which penalizes the actions out of data distribution; and IQL [17], which utilizes expectile regression to estimate the value function. For locomotion tasks, we also compare with EDAC [1], which penalizes the Q-values according to the uncertainty of Q-functions
Model-based offline RL.
We consider MOPO [34] and MOBILE [30], which penalize Q-values according to the transition uncertainty and the bellman uncertainty of a world model, respectively; COMBO [35], which combines CQL with MBPO; RAMBO [27], which trains an adversarial world model against the policy; and CBOP [14], which utilizes multi-step returns for critic updates.
5.3 Results on long-horizon AntMaze tasks
As shown in Table 1, LEQ significantly outperforms the prior model-based approaches for all $8$ datasets. LEQ achieves $58.6$ and $60.2$ for antmaze-large-play and antmaze-large-diverse, while the second best method, RAMBO [27], scores only $0.0$ and $2.4$, respectively. We believe these performance gains come from our conservative value estimation, which works more stable than the uncertainty-based penalization of prior works.
Moreover, LEQ even significantly outperforms the model-free approaches in antmaze-umaze, antmaze-large, and antmaze-ultra. Despite its superior performance, LEQ often shows high variance during training, resulting in worse performance on antmaze-medium. Over the course of training, LEQ mostly achieves high success rates, but the evaluation results sometimes drops to $0\%$ as shown in Appendix, Figure 5. We leave the problem of reducing the high variance of our method in certain environments as a future work.
Table 1: AntMaze results. Each number represents the average success rate on $100$ trials over different seeds. The results for LEQ, MOBILE, and CBOP are averaged over $5$ seeds. The results for other methods are reported following their respective papers.
Model-free Model-based
Dataset BC TD3+BC CQL IQL MOPO COMBO RAMBO MOBILE CBOP LEQ (ours)
antmaze-umaze $65.0$ $78.6$ $74.0$ $87.5$ $0.0$ $80.3$ $25.0$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $\mathbf{94.4}$$\mathbf{\pm 6.3}$
antmaze-umaze-diverse $55.0$ $71.4$ $84.0$ $62.2$ $0.0$ $57.3$ $0.0$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $\mathbf{71.0}$$\mathbf{\pm 12.3}$
antmaze-medium-play $0.0$ $3.0$ $61.2$ $\mathbf{71.2}$ $0.0$ $0.0$ $16.4$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $58.8$$\pm 33.0$
antmaze-medium-diverse $0.0$ $10.6$ $53.7$ $\mathbf{70.0}$ $0.0$ $0.0$ $23.2$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $46.2$$\pm 23.2$
antmaze-large-play $0.0$ $0.0$ $15.8$ $39.6$ $0.0$ $0.0$ $0.0$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $\mathbf{58.6}$$\mathbf{\pm 9.1}$
antmaze-large-diverse $0.0$ $0.2$ $14.9$ $47.5$ $0.0$ $0.0$ $2.4$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $\mathbf{60.2}$$\mathbf{\pm 18.3}$
antmaze-ultra-play -- -- -- $8.3$- -- -- $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $\mathbf{25.8}$$\mathbf{\pm 18.2}$
antmaze-ultra-diverse -- -- -- $15.6$- -- -- $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $\mathbf{55.8}$$\mathbf{\pm 18.3}$
Total w/o antmaze-ultra $120.0$ $163.8$ $303.6$ $354.1$ $0.0$ $137.6$ $67.0$ $0.0$ $0.0$ $\mathbf{388.8}$
Total -- -- -- $378.0$- -- -- $0.0$ $0.0$ $\mathbf{470.4}$
[…]
5.4 Results on MuJoCo Gym locomotion tasks
For D4RL MuJoCo Gym tasks in Table 3, LEQ achieves comparable results with the best score of prior works in $6$ out of $12$ tasks. Furthermore, in Table 2, LEQ outperforms most of the prior works in the NeoRL benchmark, especially in the Hopper and Walker2d domains. These results show that LEQ serves as a general offline RL algorithm, not limited to long-horizon tasks.
Similar to antmaze-medium, LEQ also suffers from the high variance problem. During training, LEQ often achieves high performance, but then, suddenly falls back to 00, as shown in Appendix, Figure 5. This is mainly because the learned models sometimes fail to capture failures (e.g. hopper and walker falling off) and predict an optimistic future (e.g. hopper and walker walking forward).
Table 2: NeoRL results. LEQ and IQL results are averaged over $5$ seeds. The results for prior works are reported following Sun et al. [30] and Qin et al. [26]. MOPO is an improved version of MOPO presented in Sun et al. [30]. We highlight the results that are better than $95\%$ of the best score.
Model-free Model-based
Dataset BC TD3+BC CQL EDAC IQL MOPO MOBILE LEQ (ours)
Hopper-L $15.1$ $15.8$ $16.0$ $18.3$ $16.7$ $6.2$ $17.4$ $\mathbf{24.2}$$\pm 2.3$
Hopper-M $51.3$ $70.3$ $64.5$ $44.9$ $28.4$ $1.0$ $51.1$ $\mathbf{104.3}$$\pm 5.2$
Hopper-H $43.1$ $75.3$ $76.6$ $52.5$ $22.3$ $11.5$ $87.8$ $\mathbf{95.5}$$\pm 13.9$
Walker2d-L $28.5$ $43.0$ $44.7$ $40.2$ $30.7$ $11.6$ $37.6$ $\mathbf{65.1}$$\pm 2.3$
Walker2d-M $48.7$ $58.5$ $57.3$ $57.6$ $51.8$ $39.9$ $\mathbf{62.2}$ $45.2$$\pm 19.4$
Walker2d-H $\mathbf{72.6}$ $69.6$ $\mathbf{75.3}$ $\mathbf{75.5}$ $\mathbf{76.3}$ $18.0$ $\mathbf{74.9}$ $\mathbf{73.7}$$\pm 1.1$
HalfCheetah-L $29.1$ $30.0$ $38.2$ $31.3$ $30.7$ $40.1$ $\mathbf{54.7}$ $33.4$$\pm 1.6$
HalfCheetah-M $49.0$ $52.3$ $54.6$ $54.9$ $51.8$ $62.3$ $\mathbf{77.8}$ $59.2$$\pm 3.9$
HalfCheetah-H $71.4$ $75.3$ $77.4$ $\mathbf{81.4}$ $76.3$ $65.9$ $\mathbf{83.0}$ $71.8$$\pm 8.0$
Total $408.8$ $490.1$ $504.6$ $456.6$ $385.0$ $256.5$ $\mathbf{546.5}$ $\mathbf{572.4}$
Table 3: D4RL MuJoCo Gym results. Each number is a normalized score averaged over $100$ trials [6]. Our results are averaged over $5$ seeds. The results for prior works are reported following their respective papers. MOPO is an improved version of MOPO, introduced in Sun et al. [30]. We highlight the results that are better than $95\%$ of the best score.
Model-free Model-based
Dataset BC TD3+BC CQL EDAC IQL MOPO COMBO RAMBO MOBILE CBOP LEQ (ours)
hopper-r $3.7$ $8.5$ $5.3$ $25.3$ $7.6$ $31.7$ $17.9$ $25.4$ $\mathbf{31.9}$ $\mathbf{32.8}$ $\mathbf{32.4}$$\pm 0.3$
hopper-m $54.1$ $59.3$ $61.9$ $\mathbf{101.6}$ $66.3$ $62.8$ $97.2$ $87.0$ $\mathbf{106.6}$ $\mathbf{102.6}$ $\mathbf{103.4}$$\pm 0.3$
hopper-mr $16.6$ $60.9$ $86.3$ $\mathbf{101.0}$ $94.7$ $\mathbf{99.4}$ $\mathbf{103.5}$ $89.5$ $99.5$ $\mathbf{104.3}$ $\mathbf{103.9}$$\pm 1.3$
hopper-me $53.9$ $98.0$ $96.9$ $\mathbf{110.7}$ $91.5$ $81.6$ $\mathbf{111.1}$ $88.2$ $\mathbf{112.6}$ $\mathbf{111.6}$ $\mathbf{109.4}$$\pm 1.8$
walker2d-r $1.3$ $1.6$ $5.4$ $16.6$ $5.2$ $7.4$ $7.0$ $0.0$ $17.9$ $17.8$ $\mathbf{21.5}$$\pm 0.1$
walker2d-m $70.9$ $83.7$ $79.5$ $\mathbf{92.5}$ $78.3$ $81.3$ $84.1$ $81.9$ $84.9$ $\mathbf{87.7}$ $74.9$$\pm 26.9$
walker2d-mr $20.3$ $81.8$ $76.8$ $87.1$ $73.9$ $85.6$ $56.0$ $89.2$ $89.9$ $92.7$ $\mathbf{98.7}$$\pm 6.0$
walker2d-me $90.1$ $110.1$ $109.1$ $\mathbf{114.7}$ $109.6$ $\mathbf{112.9}$ $103.3$ $56.7$ $\mathbf{115.2}$ $\mathbf{117.2}$ $108.2$$\pm 1.3$
halfcheetah-r $2.2$ $11.0$ $31.3$ $28.4$ $11.8$ $\mathbf{38.5}$ $\mathbf{38.8}$ $\mathbf{39.5}$ $\mathbf{39.3}$ $32.8$ $30.8$$\pm 3.3$
halfcheetah-m $43.2$ $48.3$ $46.9$ $65.9$ $47.4$ $73.0$ $54.2$ $\mathbf{77.9}$ $\mathbf{74.6}$ $\mathbf{74.3}$ $71.7$$\pm 4.4$
halfcheetah-mr $37.6$ $44.6$ $45.3$ $61.3$ $44.2$ $\mathbf{72.1}$ $55.1$ $\mathbf{68.7}$ $\mathbf{71.7}$ $66.4$ $65.5$$\pm 1.1$
halfcheetah-me $44.0$ $90.7$ $95.0$ $\mathbf{106.3}$ $86.7$ $90.8$ $90.0$ $95.4$ $\mathbf{108.2}$ $\mathbf{105.4}$ $\mathbf{102.8}$$\pm 0.4$
Total $437.9$ $698.5$ $739.7$ $911.4$ $717.2$ $844.0$ $802.0$ $812.4$ $\mathbf{959.5}$ $\mathbf{953.4}$ $\mathbf{923.2}$
5.5 Ablation studies
To understand why LEQ (LEQ-$\lambda$) works well in long-horizon tasks, we conduct ablation studies and answer to the following four questions: (1) Does using $\lambda$-return help? (2) Is LEQ better than prior uncertainty-based penalization methods? (3) Which factor enables LEQ to work in AntMaze? and (4) How do imagination length $H$ and data expansion length $R$ affect the performance?
(1) $\lambda$-returns.
To verify the effect of $\lambda$-return, we compare our method (LEQ-$\lambda$) with the versions with $1$-step return (LEQ-$1$) and $H$-step return (LEQ-$H$). Table 4 shows that using $\lambda$-return drastically improves the performance on AntMaze compared to using $1$-step return or $H$-step return. This result is coherent with the observations in prior online RL methods [28, 11].
(2) Lower expectile Q-learning.
We compare our lower expectile Q-learning with another conservative value estimator, MOBIP used in MOBILE [30], which penalizes Q-values with the standard deviation of Q-ensemble networks. The only difference between LEQ and MOBIP is their target Q-value computation for both critic and policy updates. Table 4 shows that using MOBIP not only deteriorates the success rates (in MOBIP-$1$) but also does not benefit from $\lambda$-return (in MOBIP-$\lambda$).
(3) What makes offline model-based RL work in AntMaze?
Prior to LEQ, none of offline model-based RL methods work in AntMaze, whereas our method even outperforms model-free methods. Thus, we investigate which changes in LEQ enable offline model-based RL work in AntMaze.
Table 4: Ablation study results on the AntMaze tasks. (1) We compare different Q-targets, LEQ-$\lambda$, LEQ-$1$, and LEQ-$H$. (2) We compare our lower expectile Q-learning strategy with another conservative Q-value estimation, MOBIP-$1$ and MOBIP-$\lambda$. (3) MOBILE is our re-implementation of MOBILE using MOBILE’s default hyperparameters, i.e., $\beta=0.95$, $\gamma=0.99$, and $R=5$. We note that lowering $\beta$ to $0.25$ is crucial for MOBILE to achieve meaningful scores in AntMaze.
Dataset umaze medium large ultra Total
umaze diverse play diverse play diverse play diverse
LEQ-$\lambda$ (ours) $\mathbf{94.4}$$\pm 6.3$ $\mathbf{71.0}$$\pm 12.3$ $58.8$$\pm 33.0$ $46.2$$\pm 23.2$ $\mathbf{58.6}$$\pm 9.1$ $\mathbf{60.2}$$\pm 18.3$ $25.8$$\pm 18.2$ $\mathbf{55.8}$$\pm 18.3$ $\mathbf{470.4}$
LEQ-$H$ $\mathbf{93.0}$$\pm 3.4$ $60.7$$\pm 10.4$ $46.3$$\pm 32.4$ $0.0$ ±$\pm 0.0$ $\mathbf{57.0}$$\pm 25.6$ $33.3$$\pm 43.0$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $290.3$
LEQ-$1$ $89.6$$\pm 4.8$ $37.0$$\pm 32.8$ $55.8$$\pm 28.7$ $29.8$$\pm 24.5$ $34.2$$\pm 13.4$ $49.3$$\pm 9.0$ $\mathbf{42.2}$$\pm 13.2$ $35.6$$\pm 13.0$ $373.5$
MOBIP-$\lambda$ $84.3$$\pm 3.5$ $40.3$$\pm 20.4$ $51.3$$\pm 9.0$ $39.7$$\pm 12.5$ $28.3$$\pm 21.5$ $33.7$$\pm 10.0$ $38.0$$\pm 27.1$ $23.3$$\pm 4.9$ $338.9$
MOBIP-1 $59.5$$\pm 3.5$ $46.5$$\pm 1.5$ $57.0$$\pm 11.0$ $\mathbf{54.0}$$\pm 9.0$ $23.5$$\pm 19.5$ $38.5$$\pm 1.5$ $39.5$$\pm 11.5$ $20.5$$\pm 20.5$ $339.0$
MOBILE $1.0$$\pm 2.0$ $0.0$ ±$\pm 0.0$ $6.4$$\pm 5.5$ $5.0$$\pm 5.0$ $0.8$$\pm 1.6$ $0.8$$\pm 1.2$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $14.0$
MOBILE ($\mathbf{\beta}$ = 0.25) $77.0$$\pm 6.4$ $20.4$$\pm 15.7$ $\mathbf{64.6}$$\pm 11.1$ $31.6$$\pm 16.9$ $2.6$$\pm 2.8$ $7.2$$\pm 8.9$ $4.6$$\pm 3.0$ $5.0$$\pm 4.6$ $213.0$
MOBILE ($\mathbf{\gamma}$ = 0.997) $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $7.2$$\pm 4.1$ $1.6$$\pm 2.1$ $9.6$$\pm 7.1$ $5.4$$\pm 4.9$ $0.0$ ±$\pm 0.0$ $1.8$$\pm 2.7$ $25.6$
MOBILE ($R$ = 10) $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $5.0$$\pm 5.1$ $0.6$$\pm 1.2$ $7.4$$\pm 14.8$ $1.6$$\pm 3.2$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $14.6$
We first re-implement MOBILE with some technical tricks used in LEQ: LayerNorm [3], SymLog [11], single Q-network, and no target Q-value clipping; but, MOBILE achieves a barely non-zero score, $14.0$. We found that the key to make MOBILE work is reducing $\beta$, the ratio for the loss calculated from imaginary rollouts and from dataset transitions. When we lower $\beta$ from $0.95$ to $0.25$ (used in LEQ), MOBILE shows meaningful performances in umaze and medium mazes, and achieves $213.0$ in total. We suggest that utilizing the true transition from the dataset is important in long-horizon tasks, which was undervalued in prior works.
(4) Imagination length $H$ and dataset expansion length $R$.
As shown in Table 5, the performance increases when it goes to $H=10$ from $H=5$, but it drops when $H=15$. This result shows the trade-off of using the world model: the further the agent imagines, more the agent becomes robust to the error of the critic, but more it becomes prone to the error from the model prediction.
We also evaluate LEQ without the dataset expansion ($H=10,R=1$). In AntMaze, the results with and without the dataset expansion are similar, as shown in Table 5. On the other hand, the dataset expansion makes the policy more stable and better in the D4RL MuJoCo tasks (Table 13).
Table 5: LEQ with different imagination length $H$ and data expansion length $R$. A longer $H$ can mitigate critic biases, while increasing model errors, which leads to poor performance. Each number is averaged over $5$ random seeds.
Dataset $\mathbf{H=10,R=5}$ (ours) $H=5,R=5$ $H=15,R=5$ $H=10,R=1$
antmaze-umaze $\mathbf{94.4}$$\pm 6.3$ $\mathbf{95.2}$$\pm 1.7$ $\mathbf{98.6}$$\pm 0.5$ $\mathbf{97.4}$$\pm 1.4$
antmaze-umaze-diverse $\mathbf{71.0}$$\pm 12.3$ $\mathbf{67.2}$$\pm 9.1$ $\mathbf{70.7}$$\pm 15.2$ $63.0$$\pm 23.2$
antmaze-medium-play $58.8$$\pm 33.0$ $46.4$$\pm 31.9$ $\mathbf{76.3}$$\pm 17.2$ $58.2$$\pm 28.0$
antmaze-medium-diverse $\mathbf{46.2}$$\pm 23.2$ $18.6$$\pm 28.7$ $30.3$$\pm 40.1$ $28.6$$\pm 33.7$
antmaze-large-play $58.6$$\pm 9.1$ $48.6$$\pm 15.4$ $\mathbf{62.0}$$\pm 9.9$ $56.0$$\pm 9.8$
antmaze-large-diverse $\mathbf{60.2}$$\pm 18.3$ $35.2$$\pm 8.7$ $33.0$$\pm 3.2$ $\mathbf{57.0}$$\pm 4.5$
antmaze-ultra-play $25.8$$\pm 18.2$ $\mathbf{54.2}$$\pm 10.8$ $0.0$ ±$\pm 0.0$ $39.2$$\pm 15.1$
antmaze-ultra-diverse $\mathbf{55.8}$$\pm 18.3$ $39.4$$\pm 6.1$ $0.0$ ±$\pm 0.0$ $36.0$$\pm 12.0$
Total $\mathbf{470.4}$ $404.8$ $371.0$ $435.4$
6 Conclusion
In this paper, we propose a novel offline model-based reinforcement learning method, LEQ, which uses expectile regression to get a conservative evaluation of a policy from model-generated trajectories. Expectile regression eases the pain of constructing the whole distribution of Q-targets and allows for estimating the conservative value via sampling. Combined with $\lambda$-returns in both critic and policy updates for the imaginary rollouts, the policy can receive learning signals that are more robust to both model errors and critic errors. We empirically show that LEQ improves the performance in various tasks – especially, achieving the state-of-the-art performance in the long-horizon AntMaze tasks.
6.1 Limitations
Following prior work on model-based offline RL [30, 14], we assume access to the ground-truth termination function of a task, different from online model-based RL approaches, which learn a termination function from interactions. However, since this termination function is conditioned on a state, a model requires to plan on a state space (or an observation space), which could be challenging in a high-dimensional state space (e.g. pixel observations). Extending the proposed approach to complex environments with high-dimensional observations would be an immediate next step.
6.2 Broader Impacts
Our method aims to increase the ability of autonomous agents, such as robots and self-driving cars, to learn from static, offline data without interacting with the world. This enables autonomous agents to utilize data with diverse qualities (not necessarily from experts). We believe that this paper does not have any immediate negative societal impact.
Acknowledgments and Disclosure of Funding
We would like to thank Junik Bae for helpful discussion. This work was supported in part by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant (RS-2020-II201361, Artificial Intelligence Graduate School Program (Yonsei University)) and the National Research Foundation of Korea (NRF) grant (RS-2024-00333634) funded by the Korean Government (MSIT). Kwanyoung Park was supported by Electronics and Telecommunications Research Institute (ETRI).
Appendix A Training Details
Computing resources.
All experiments are done on a single RTX 4090 GPU and $4$ AMD EPYC 9354 CPU cores. We use $5$ different random seeds for each experiment and report the mean and standard deviation. Each offline RL experiment takes $2$ hours for ours, $12$ hours for MOBILE, and $24$ hours for CBOP.
Environment details.
For the locomotion tasks, we use the dataset provided by D4RL [6] and NeoRL [26]. Following IQL [17], we normalize rewards using the maximum and minimum return of all trajectories. We use the true termination functions of the environments, implemented in MOBILE [30].
For the AntMaze tasks, we use the dataset provided by D4RL [6]. Following IQL [17], we subtract $1$ from the rewards in the datasets so that the agent receives $-1$ for each step and 00 on termination. We use the true termination functions of the environments. The termination functions of the AntMaze tasks are not deterministic because a goal of a maze is randomized every time the environment is reset. Nevertheless, we follow the implementation of CBOP [14], where the termination region is set to a circle around the mean of the goal distribution with the radius $0.5$.
Method implementation details.
For all compared methods, we use the results from their corresponding papers when available. For IQL [17], we run the official implementation with $5$ seeds to reproduce the results for the random datasets in D4RL and NeoRL. For the AntMaze tasks, we run the official implementation of MOBILE and CBOP with $5$ random seeds. Please note that the original MOBILE implementation does not use the true termination function, so we replace it with our termination function. For MOPO, COMBO, and RAMBO, we use the results reported in RAMBO [27].
World models.
For training world models, we use the architecture and training script from OfflineRL-Kit [29], matching the implementation of MOBILE [30]. Each world model is implemented as a $4$-layer MLPs with the hidden layer size of $200$. We construct an ensemble of world models by selecting $5$ out of $7$ models with the best validation scores. We pretrain the ensemble of world models for each of $5$ random seeds (i.e. training in total $35$ world models and using $25$ models), which takes approximately $5$ hours in average.
Policy and critic networks.
We use $3$-layer MLPs with size of $256$ both for the policy network and the critic network. We use layer normalization [3] to prevent catastrophic over/underestimation [4], and squash the state inputs using symlog to keep training stable from outliers in long-horizon model rollouts [11].
Pretraining.
For some environments, we found that a randomly initialized policy can lead to abnormal rewards or transition prediction from the world models in the early stage, leading to unstable training. Following CBOP [14], we pretrain a policy $\pi_{\theta}$ and a critic $Q_{\phi}$ using behavioral cloning and FQE [20], respectively. We use a slightly different implementation of FQE from the original implementation, where the $\arg\min$ operation is approximated with mini-batch gradient descent, similar to standard Q-learning as shown in Algorithm 2.
Algorithm 2 FQE: Fitted Q Evaluation [20]
0:  Offline dataset $\mathcal{D}_{\text{env}}$, policy $\pi_{\theta}$
1:  Randomly initialize Q-function $Q_{\phi}$
2:  while not converged do
3:     $\{\mathbf{s}_{i},\mathbf{a}_{i},r_{i},\mathbf{s}^{\prime}_{i}\}_{i=1}^{N}\sim\mathcal{D}_{\text{env}}$
4:     $y_{i}=\texttt{sg}(Q_{\phi}(\mathbf{s}^{\prime}_{i},\pi_{\theta}(\mathbf{s}^{\prime}_{i})))$triangleright $\texttt{sg}(\cdot)$ is stop-gradient operator
5:     $L_{\text{FQE}}(\phi)=\frac{1}{N}\sum_{i=1}^{N}(Q_{\phi}(\mathbf{s}_{i},\mathbf{a}_{i})-y_{i})^{2}$
6:     Update $Q_{\phi}$ using gradient descent to minimize $L_{\text{FQE}}(\phi)$
Comparisons with prior methods.
We provide a comparison of LEQ with the prior model-based approaches and the baseline methods used in our ablation studies in Table 6.
Table 6: Comparisons with the prior model-based approaches and the baseline methods introduced for our ablation studies. We color the hyperparameters in blue if they are the same with LEQ. Otherwise, we color them in red.
Components CBOP MOBILE MOBILE MOBIP LEQ (ours)
Training scheme MVE [5] MBPO [13] MBPO [13] Dyna [32] Dyna [32]
Conservatism Lower-confidence bound Lower-confidence bound Lower-confidence bound Lower-confidence bound Lower expectile
Policy Stochastic Stochastic Stochastic Deterministic Deterministic
Policy objective $Q(\mathbf{s},\mathbf{a})$ $Q(\mathbf{s},\mathbf{a})$ $Q(\mathbf{s},\mathbf{a})$ $\lambda$-returns $\lambda$-returns
Policy pretraining BC BC BC
# of critics 20-50 2 1 1 1
Critic objective Multi-step (adaptive weighting) One-step One-step $\lambda$-returns $\lambda$-returns
Critic pretraining FQE [20] FQE [20] FQE [20]
Horizon length ($H$) 10 1 1 10 10
Rollout length ($R$) 1 or 5 10 5 5
Discount rate ($\gamma$) 0.99 0.99 0.997 0.997 0.997
$\beta$ in Equation 7 1.0 0.95 0.25 0.25 0.25
Impl. tricks Clip Q-values with 00 LayerNorm + Symlog LayerNorm + Symlog LayerNorm + Symlog
Running time 24h 12h 40m 4h 2h
Hyperparameters of LEQ.
We report task-agnostic hyperparameters of our method in Table 7. We note that we use the same hyperparameters across all tasks, except $\tau$. We search the value of $\tau$ in $\{0.1,0.3,0.4,0.5\}$ and report the best value for the main experimental results. In addition, we report the exhaustive results in Tables 11 and 12, and summarize $\tau$ used in the main results in Table 8.
Table 7: Shared hyperparameters of LEQ.
Hyperparameters Value Description
$lr_{\text{actor}}$ 3e-5 Learning rate of actor
$lr_{\text{critic}}$ 1e-4 Learning rate of critic
Optimizer Adam Optimizer
$T_{\text{expand}}$ 5000 Interval of expanding dataset
$N_{\text{expand}}$ 50000 Number of data for each expansion of dataset
$R$ 5 Rollout length for dataset expansion
$\sigma_{\text{exp}}$ 1.0 Exploration noise for dataset expansion
$N_{\text{iter}}$ 1M Total number of gradient steps.
$B_{\text{env}}$ 256 Batch size from original dataset
$B_{\text{model}}$ 256 Batch size from expanded dataset
$\gamma$ 0.997 Discount factor
$\lambda$ 0.95 $\lambda$ value for $\lambda$-return
$H$ 10 Imagination length
$\beta_{\text{EMA}}$ 1 Weight for critic EMA regularization
$\epsilon_{\text{EMA}}$ 0.995 Critic EMA decay
Table 8: Task-specific hyperparameter $\tau$ of LEQ.
Domain Task $\tau$
AntMaze umaze $0.1$
umaze-diverse $0.1$
medium-play $0.3$
medium-diverse $0.1$
large-play $0.3$
large-diverse $0.3$
ultra-play $0.1$
ultra-diverse $0.1$
MuJoCo hopper-r $0.1$
hopper-m $0.1$
hopper-mr $0.3$
hopper-me $0.1$
walker2d-r $0.1$
walker2d-m $0.3$
walker2d-mr $0.5$
walker2d-me $0.1$
halfcheetah-r $0.3$
halfcheetah-m $0.3$
halfcheetah-mr $0.4$
halfcheetah-me $0.1$
NeoRL Hopper-L $0.1$
Hopper-M $0.1$
Hopper-H $0.1$
Walker2d-L $0.3$
Walker2d-M $0.1$
Walker2d-H $0.1$
HalfCheetah-L $0.1$
HalfCheetah-M $0.3$
HalfCheetah-H $0.3$
Task-specific hyperparameters of the compared methods.
We report the best hyperparameters of MOBILE for the AntMaze tasks in Tables 10 and 10. For MOBILE and MOBILE, we search the value of $c$ within $\{0.1,0.5,1.0,1.5\}$, as suggested in MOBILE [30], where $c$ is the coefficient of the penalized bellman operator:
$T\hat{Q}(\mathbf{s},\mathbf{a})=r(\mathbf{s},\mathbf{a})+\gamma Q(\mathbf{s}^{\prime},\mathbf{a}^{\prime})-c\cdot\text{Std}(Q(\mathbf{s}^{\prime},\mathbf{a}^{\prime})).$ (13)
For CBOP, we conduct hyperparameter search for $\psi$ in $\{0.5,2.0,3.0,5.0\}$, as suggested in the original paper, where $\psi$ is an LCB coefficient of CBOP. We do not report the best hyperparameter for MOBILE and CBOP because both methods score zero points for all hyperparameters in AntMaze.
Table 9: Task-specific hyperparameters in MOBILE.
Domain Task $c$
AntMaze umaze $1.0$
umaze-diverse $1.0$
medium-play $1.0$
medium-diverse $0.1$
large-play $0.1$
large-diverse $0.1$
ultra-play $1.0$
ultra-diverse $1.0$
Table 10: Task-specific hyperparameters in MOBILE with $\lambda$-returns.
Domain Task $c$
AntMaze umaze $1.0$
umaze-diverse $0.5$
medium-play $0.1$
medium-diverse $0.1$
large-play $0.1$
large-diverse $0.1$
ultra-play $1.0$
ultra-diverse $0.5$
Appendix B Proof of the Policy Objective
We show that the surrogate loss in Equation 12 leads to a better approximation for the expectile of $\lambda$-returns in Equation 11 than maximizing $Q_{\phi}(s,a)$. In other words, we show that optimizing the following policy objective:
$\hat{J}_{\lambda}(\theta)=\mathbb{E}_{\tau\sim p_{\psi},\pi_{\theta}}[(W^{\tau}(Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})>Q^{\lambda}_{t}(\tau))Q_{t}^{\lambda}(\tau)],$ (14)
leads to optimizing a lower-bias estimator of $\mathbb{E}^{\tau}_{\tau\sim p_{\psi},\pi_{\theta}}[Q_{t}^{\lambda}(\tau)]$ than $Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})$.
To show this, we first prove that $\hat{Y}_{\text{new}}=\frac{\mathbb{E}[W^{\tau}(Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})>Q_{t}^{\lambda}(\tau))\cdot Q_{t}^{\lambda}(\tau)]}{\mathbb{E}[W^{\tau}(Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})>Q_{t}^{\lambda}(\tau))]}$ is closer to $\mathbb{E}^{\tau}_{\tau\sim p_{\psi},\pi_{\theta}}[Q_{t}^{\lambda}(\tau)]$ than $Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})$. For deriving the proof, we generalize this situation to have an arbitrary distribution $X$ and estimation $\hat{Y}$, which corresponds to $X=Q_{t}^{\lambda}(\tau),\hat{Y}=Q_{\phi}(\mathbf{s},\mathbf{a})$.
Theorem 1.
Let $X$ be a distribution and $Y=E^{\tau}[X]$ be a lower expectile of $X$ (i.e. $0<\tau\leq 0.5$). Let $\hat{Y}$ be an arbitrary estimation of $Y$, and define $W^{\tau}(\cdot)=|\tau-\mathbbm{1}(\cdot)|$. If we let $\hat{Y}_{\text{new}}=\frac{\mathbb{E}[W^{\tau}(\hat{Y}>X)\cdot X]}{\mathbb{E}[W^{\tau}(\hat{Y}>X)]}$ be a new estimation of $Y$, then $|Y-\hat{Y}_{\text{new}}|\leq\frac{1-2\tau}{1-\tau}\cdot p(Y\leq X\leq\hat{Y})\cdot|Y-\hat{Y}|$.
Proof.
Without loss of generality, we assume $\hat{Y}\geq Y$. Then, we have $\hat{Y}_{\text{new}}\geq Y$. Thus,
$|\hat{Y}_{\text{new}}-Y|$
$=\hat{Y}_{\text{new}}-Y$
$=\frac{\mathbb{E}[W^{\tau}(\hat{Y}>X)\cdot X]}{\mathbb{E}[W^{\tau}(\hat{Y}>X)]}-\frac{\mathbb{E}[W^{\tau}(Y>X)\cdot X]}{\mathbb{E}[W^{\tau}(Y>X)]}\quad\text{\qquad\qquad\qquad\qquad\qquad\qquad\;\;\;($\because$ Def. of $\hat{Y}_{\text{new}}$ and $Y$)}$
$=\frac{\mathbb{E}[W^{\tau}(Y>X)\cdot X]+\mathbb{E}[(1-2\tau)\cdot\mathbbm{1}(Y\leq X\leq\hat{Y})X]}{\mathbb{E}[W^{\tau}(Y>X)]+\mathbb{E}[(1-2\tau)\cdot\mathbbm{1}(Y\leq X\leq\hat{Y})]}-\frac{\mathbb{E}[W^{\tau}(Y>X)\cdot X]}{\mathbb{E}[W^{\tau}(Y>X)]}\quad\text{($\because$ Def. of $W^{\tau}(\cdot)$)}$
$=\frac{\mathbb{E}[W^{\tau}(Y>X)]\mathbb{E}[(1-2\tau)\cdot\mathbbm{1}(Y\leq X\leq\hat{Y})X]-\mathbb{E}[W^{\tau}(Y>X)X]\mathbb{E}[(1-2\tau)\cdot\mathbbm{1}(Y\leq X\leq\hat{Y})]}{\mathbb{E}[W^{\tau}(Y>X)](\mathbb{E}[W^{\tau}(Y>X)]+\mathbb{E}[(1-2\tau)\cdot\mathbbm{1}(Y\leq X\leq\hat{Y})])}$
$\leq\frac{1-2\tau}{\mathbb{E}[W^{\tau}(Y>X)]^{2}}\cdot(\mathbb{E}[W^{\tau}(Y>X)]\mathbb{E}[\mathbbm{1}(Y\leq X\leq\hat{Y})X]-\mathbb{E}[W^{\tau}(Y>X)X]\mathbb{E}[\mathbbm{1}(Y\leq X\leq\hat{Y})]$
$=\frac{1-2\tau}{\mathbb{E}[W^{\tau}(Y>X)]}\cdot(\mathbb{E}[\mathbbm{1}(Y\leq X\leq\hat{Y})X]-Yp(Y\leq X\leq\hat{Y}))$
$=\frac{(1-2\tau)p(Y\leq X\leq\hat{Y})}{\mathbb{E}[W^{\tau}(Y>X)]}\cdot(\mathbb{E}_{Y\leq X\leq\hat{Y}}[X]-Y)$
$\leq\frac{(1-2\tau)p(Y\leq X\leq\hat{Y})}{\mathbb{E}[W^{\tau}(Y>X)]}\cdot(\hat{Y}-Y)$
$\leq\frac{1-2\tau}{1-\tau}\cdot p(Y\leq X\leq\hat{Y})\cdot|\hat{Y}-Y|$
Note that this theorem shows that the bias of the new estimation is always smaller than the original estimation, since $\frac{1-2\tau}{1-\tau}<1$ and $p(Y\leq X\leq\hat{Y})\leq 1$. If we plug in the distribution of $Q_{t}^{\lambda}(\tau)$ to $X$ and $\hat{Y}=Q_{\phi}(s_{t},\mathbf{a}_{t})$, then $Y=\mathbb{E}^{\tau}[X]=\mathbb{E}^{\tau}_{\tau\sim p_{\psi},\pi_{\theta}}[Q_{t}^{\lambda}(\tau)]$, and we can show the desired result using the theorem: $\hat{Y}_{\text{new}}=\frac{\mathbb{E}[W^{\tau}(Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})>Q_{t}^{\lambda}(\tau))\cdot Q_{t}^{\lambda}(\tau)]}{\mathbb{E}[W^{\tau}(Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})>Q_{t}^{\lambda}(\tau)]}$ is closer to $\mathbb{E}^{\tau}_{\tau\sim p_{\psi},\pi_{\theta}}[Q_{t}^{\lambda}(\tau)]$ than $Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})$.
Here, the normalizing factor $\mathbb{E}[W^{\tau}(Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})>Q_{t}^{\lambda}(\tau))]$ is non-differentiable with $\tau$. Specifically, the gradient is 0 everywhere (except $Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})=Q_{t}^{\lambda}(\tau)$). Thus, if we calculate the gradient of $\hat{Y}_{\text{new}}$, the gradient for the normalizing factor disappears. Therefore, we can omit the normalizing factor and get an equivalent formula $\mathbb{E}[W^{\tau}(Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})>Q_{t}^{\lambda}(\tau))\cdot Q_{t}^{\lambda}(\tau)]$ for gradient-based optimization.
Appendix C More Results
High variance in locomotion tasks.
When we train LEQ in locomotion tasks, we observe that our method often achieves $100$% success rates and then falls back to 00%, as shown in Figure 5. This is mainly because the learned models sometimes fail to capture failures (e.g. hopper and walker falling off) and predict an optimistic future (e.g. hopper and walker walking forward).
Figure 5: High variance during training. Our algorithm experiences oscillation on the performances due to optimistic imaginations near the initial states.
Results for all expectiles $\tau$.
To give insights how the expectile parameter $\tau$ affects the performance of LEQ, we report the performance of LEQ with all expectile values $\{0.1,0.3,0.4,0.5\}$. The expectile parameter $\tau$ has a trade-off – high expectile makes the model’s predictions less conservative while making a policy easily exploit the model. We recommend first trying $\tau=0.1$, which works well for most of the tasks, and increase $\tau$ until the performance starts to drop.
Table 11: Antmaze results of LEQ with different expectiles. We report the results in Antmaze task with expectiles value of 0.1, 0.3, 0.4, 0.5. The best value is highlighted.
Expectile 0.1 0.3 0.4 0.5
antmaze-umaze $\mathbf{94.4}$$\pm 6.3$ $39.0$$\pm 28.1$ $0.2$$\pm 0.4$ $3.0$$\pm 5.5$
antmaze-umaze-diverse $\mathbf{71.0}$$\pm 12.2$ $23.6$$\pm 21.7$ $4.0$$\pm 4.2$ $0.0$ ±$\pm 0.0$
antmaze-medium-play $50.2$$\pm 39.9$ $\mathbf{58.8}$$\pm 33.0$ $36.0$$\pm 21.8$ $0.6$$\pm 1.2$
antmaze-medium-diverse $\mathbf{46.2}$$\pm 23.2$ $13.2$$\pm 13.3$ $11.6$$\pm 14.8$ $10.6$$\pm 13.3$
antmaze-large-play $42.0$$\pm 30.6$ $\mathbf{58.6}$$\pm 9.1$ $52.2$$\pm 15.8$ $42.2$$\pm 7.3$
antmaze-large-diverse $60.6$$\pm 32.1$ $\mathbf{60.2}$$\pm 18.3$ $48.8$$\pm 5.8$ $36.8$$\pm 9.7$
antmaze-ultra-play $\mathbf{25.8}$$\pm 18.2$ $10.8$±$\pm 8.8$ $11.6$$\pm 12.5$ $9.2$$\pm 11.5$
antmaze-ultra-diverse $\mathbf{55.8}$$\pm 18.3$ $4.6$±$\pm 3.4$ $7.6$$\pm 7.3$ $0.6$$\pm 1.2$
Table 12: D4RL mujoco results of LEQ with different expectiles. We report the results in D4RL mujoco task with expectiles value of 0.1, 0.3, 0.4, 0.5. The best value is highlighted.
Expectile 0.1 0.3 0.4 0.5
hopper-r $\mathbf{32.4}$$\pm 0.3$ $13.7$$\pm 9.1$ $16.4$$\pm 9.3$ $12.5$$\pm 10.1$
hopper-m $\mathbf{103.4}$$\pm 0.3$ $102.7$$\pm 1.7$ $81.4$$\pm 24.8$ $38.6$$\pm 29.2$
hopper-mr $103.2$$\pm 1.0$ $\mathbf{103.9}$$\pm 1.3$ $71.5$$\pm 34.7$ $103.8$$\pm 1.9$
hopper-me $\mathbf{109.4}$$\pm 1.8$ $108.0$$\pm 8.7$ $64.2$$\pm 35.8$ $33.7$$\pm 0.5$
walker2d-r $\mathbf{21.5}$$\pm 0.1$ $21.5$$\pm 0.5$ $14.0$$\pm 8.8$ $8.7$$\pm 6.7$
walker2d-m $26.3$$\pm 37.4$ $\mathbf{74.9}$$\pm 26.9$ $60.3$$\pm 40.9$ $34.8$$\pm 34.3$
walker2d-mr $48.6$$\pm 19.5$ $60.5$$\pm 27.4$ $88.5$$\pm 3.5$ $\mathbf{98.7}$$\pm 6.0$
walker2d-me $\mathbf{108.2}$$\pm 1.3$ $98.8$$\pm 28.8$ $105.8$$\pm 25.9$ $33.7$$\pm 31.9$
halfcheetah-r $23.8$$\pm 1.8$ $\mathbf{30.8}$$\pm 3.3$ $29.0$$\pm 2.9$ $30.2$$\pm 2.5$
halfcheetah-m $65.3$$\pm 2.0$ $\mathbf{71.7}$$\pm 4.4$ $58.5$$\pm 23.8$ $55.5$$\pm 16.7$
halfcheetah-mr $60.6$$\pm 1.4$ $55.4$$\pm 27.3$ $\mathbf{65.5}$$\pm 1.1$ $52.4$$\pm 26.7$
halfcheetah-me $\mathbf{102.8}$$\pm 0.4$ $81.5$$\pm 19.6$ $58.1$$\pm 26.1$ $46.3$$\pm 17.7$
Ablation study on dataset expansion.
Table 13 shows the ablation results on the dataset expansion in D4RL MuJoCo tasks. The results show that the dataset expansion generally improves the performance, especially in Hopper environments.
Table 13: D4RL MuJoCo ablation results for dataset expansion. Results are averaged over $5$ random seeds. The dataset expansion generally improves the performance of LEQ.
Dataset LEQ (ours) LEQ w/o Dataset Expansion
hopper-r $\mathbf{32.4}$$\pm 0.3$ $17.6$$\pm 8.6$
hopper-m $\mathbf{103.4}$$\pm 0.3$ $52.7$$\pm 45.3$
hopper-mr $\mathbf{103.9}$$\pm 1.3$ $\mathbf{103.7}$$\pm 1.3$
hopper-me $\mathbf{109.4}$$\pm 1.8$ $79.7$$\pm 42.4$
walker2d-r $\mathbf{21.5}$$\pm 0.1$ $\mathbf{20.5}$$\pm 2.2$
walker2d-m $74.9$$\pm 26.9$ $\mathbf{87.2}$$\pm 4.3$
walker2d-mr $\mathbf{98.7}$$\pm 6.0$ $78.7$$\pm 35.5$
walker2d-me $\mathbf{108.2}$$\pm 1.3$ $\mathbf{110.4}$$\pm 0.8$
halfcheetah-r $\mathbf{30.8}$$\pm 3.3$ $\mathbf{27.7}$$\pm 2.2$
halfcheetah-m $\mathbf{71.7}$$\pm 4.4$ $\mathbf{71.6}$$\pm 3.8$
halfcheetah-mr $\mathbf{65.5}$$\pm 1.1$ $54.4$$\pm 26.3$
halfcheetah-me $\mathbf{102.8}$$\pm 0.4$ $83.9$$\pm 28.0$
Total $\mathbf{923.2}$ $788.2$
Ablation study on MOBILE in AntMaze.
In Table 14, we report the performance of MOBILE for all possible combination of the hyperparameters between the values of MOBILE and LEQ. Specifically, we use $\beta=\{0.25,0.95\}$, $\gamma=\{0.99,0.997\}$, and $R=\{5,10\}$. The result shows that $\beta=0.25$ is crucial. In addition, the configuration of LEQ yields the best result among all configurations for MOBILE.
Table 14: Complete hyperparameter search results of MOBILE on AntMaze. MOBILE uses the hyperparameters from MOBILE: $\beta=0.95$, $\gamma=0.99$, and $R=5$, whereas LEQ uses $\beta=0.25$, $\gamma=0.997$, and $R=10$. The results show that $\beta$ is the most critical hyperparameter that makes MOBILE work in AntMaze.
  Hyperparams. umaze medium large ultra Total
$\beta$ $\gamma$ $R$ umaze diverse play diverse play diverse play diverse
$0.25$ $0.997$ $10$ $53.8$$\pm 26.8$ $\mathbf{22.5}$$\pm 22.2$ $54.0$$\pm 5.8$ $\mathbf{49.5}$$\pm 6.2$ $\mathbf{28.3}$$\pm 6.0$ $\mathbf{28.0}$$\pm 11.4$ $\mathbf{25.5}$$\pm 6.9$ $\mathbf{23.8}$$\pm 15.8$ $\mathbf{285.3}$
$0.25$ $0.997$ $5$ $\mathbf{74.0}$$\pm 6.9$ $3.7$$\pm 2.6$ $54.7$$\pm 27.9$ $28.0$$\pm 9.6$ $18.7$$\pm 18.6$ $8.0$$\pm 9.3$ $9.7$$\pm 8.2$ $9.0$$\pm 3.7$ $205.7$
$0.25$ $0.99$ $10$ $39.7$$\pm 23.4$ $5.0$$\pm 7.1$ $39.3$$\pm 27.9$ $38.0$$\pm 15.0$ $0.0$ ±$\pm 0.0$ $3.7$$\pm 5.2$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $125.7$
$0.25$ $0.99$ $5$ $\mathbf{77.0}$$\pm 6.4$ $20.4$$\pm 15.7$ $\mathbf{64.6}$$\pm 11.1$ $31.6$$\pm 16.9$ $2.6$$\pm 2.8$ $7.2$$\pm 8.9$ $4.6$$\pm 3.0$ $5.0$$\pm 4.6$ $213.0$
$0.95$ $0.997$ $10$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $1.8$$\pm 3.0$ $0.5$$\pm 0.9$ $0.2$$\pm 0.4$ $2.2$ ±$\pm 2.3$ $1.0$$\pm 1.7$ $0.0$ ±$\pm 0.0$ $5.7$
$0.95$ $0.997$ $5$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $7.2$$\pm 4.1$ $1.6$$\pm 2.1$ $9.6$$\pm 7.1$ $5.4$$\pm 4.9$ $0.0$ ±$\pm 0.0$ $1.8$$\pm 2.7$ $25.6$
$0.95$ $0.99$ $10$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $5.0$$\pm 5.1$ $0.6$$\pm 1.2$ $7.4$$\pm 14.8$ $1.6$$\pm 3.2$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $14.6$
$0.95$ $0.99$ $5$ $1.0$$\pm 2.0$ $0.0$ ±$\pm 0.0$ $6.4$$\pm 5.5$ $5.0$$\pm 5.0$ $0.8$$\pm 1.6$ $0.8$$\pm 1.2$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $14.0$

### v2

@@
-Tackling Long-Horizon Tasks with Model-based Offline Reinforcement Learning
+Model-based Offline Reinforcement Learning with Lower Expectile Q-Learning
 Abstract
 Model-based offline reinforcement learning (RL) is a compelling approach that addresses the challenge of learning from limited, static data by generating imaginary trajectories using learned models.
-However, it falls short in solving long-horizon tasks due to high bias in value estimation from model rollouts.
-In this paper, we introduce a novel model-based offline RL method, Lower Expectile Q-learning (LEQ), which enhances long-horizon task performance by mitigating the high bias in model-based value estimation via expectile regression of $\lambda$-returns.
-Our empirical results show that LEQ significantly outperforms previous model-based offline RL methods on long-horizon tasks, such as the D4RL AntMaze tasks, matching or surpassing the performance of model-free approaches.
-Our experiments demonstrate that expectile regression, $\lambda$-returns, and critic training on offline data are all crucial for addressing long-horizon tasks.
-Additionally, LEQ achieves performance comparable to the state-of-the-art model-based and model-free offline RL methods on the NeoRL benchmark and the D4RL MuJoCo Gym tasks.
+However, these approaches often struggle with inaccurate value estimation from model rollouts.
+In this paper, we introduce a novel model-based offline RL method, Lower Expectile Q-learning (LEQ), which provides a low-bias model-based value estimation via lower expectile regression of $\lambda$-returns.
+Our empirical results show that LEQ significantly outperforms previous model-based offline RL methods on long-horizon tasks, such as the D4RL AntMaze tasks, matching or surpassing the performance of model-free approaches and sequence modeling approaches.
+Furthermore, LEQ matches the performance of state-of-the-art model-based and model-free methods in dense-reward environments across both state-based tasks (NeoRL and D4RL) and pixel-based tasks (V-D4RL), showing that LEQ works robustly across diverse domains.
+Our ablation studies demonstrate that lower expectile regression, $\lambda$-returns, and critic training on offline data are all crucial for LEQ.
 1 Introduction
-One of the major challenges in offline reinforcement learning (RL) is the overestimation of values for out-of-distribution actions due to the lack of environment interactions [21, 19].
-Model-based offline RL addresses this issue by generating additional (imaginary) training data using a learned model, thereby augmenting the given offline data with synthetic experiences that cover out-of-distribution states and actions [34, 16, 35, 2, 30].
-While these approaches have demonstrated strong performance in simple, short-horizon tasks, they struggle with noisy model predictions and value estimations, particularly in long-horizon tasks [23].
+One of the major challenges in offline reinforcement learning (RL) is the overestimation of values for out-of-distribution actions due to the lack of environment interactions (Levine et al., 2020; Kumar et al., 2020).
+Model-based offline RL addresses this issue by generating additional (imaginary) training data using a learned model, thereby augmenting the given offline data with synthetic experiences that cover out-of-distribution states and actions (Yu et al., 2020; Kidambi et al., 2020; Yu et al., 2021; Argenson and Dulac-Arnold, 2021; Sun et al., 2023).
+While these approaches have demonstrated strong performance in simple, short-horizon tasks, they struggle with noisy model predictions and value estimates, particularly in long-horizon tasks (Park et al., 2024).
 This challenge is evident in their poor performances (i.e.
-near zero) on the D4RL AntMaze tasks [6, 15].
-Typical model-based offline RL methods alleviate the inaccurate value estimation problem (mostly overestimation) by penalizing Q-values estimated from model rollouts with uncertainties in model predictions [34, 16] or value predictions [30, 14].
-While these penalization terms prevent a policy from exploiting erroneous value estimations, the policy now does not maximize the true value, but maximizes the value penalized by heuristically estimated uncertainties, which can lead to sub-optimal behaviors.
-This is especially problematic in long-horizon, sparse-reward tasks, where Q-values are similar across nearby states [23].
-Another way to reduce bias in value estimates is by using multi-step returns [31, 12].
-CBOP [14] constructs an explicit distribution of multi-step Q-values from thousands of model rollouts and uses this value as a target for training the Q-function.
-However, CBOP is computationally expensive for estimating a target value and uses multi-step returns solely for Q-learning, which provides insufficient learning signals for obtaining long-horizon behaviors.
-To tackle long-horizon tasks with model-based offline RL, we introduce a simple yet effective model-based offline RL algorithm, Lower Expectile Q-learning (LEQ).
+near zero) on the D4RL AntMaze tasks (Fu et al., 2020).
+Typical model-based offline RL methods alleviate the inaccurate value estimation problem (mostly overestimation) by penalizing Q-values estimated from model rollouts with uncertainties in model predictions (Yu et al., 2020; Kidambi et al., 2020) or value predictions (Sun et al., 2023; Jeong et al., 2023).
+While these penalization terms prevent a policy from exploiting erroneous value estimates, the policy now does not maximize the true value, but maximizes the value penalized by heuristically estimated uncertainties, which can lead to sub-optimal behaviors.
+This is especially problematic in long-horizon, sparse-reward tasks, where Q-values are similar across nearby states (Park et al., 2024).
+Another way to reduce bias in value estimates is using multi-step returns (Sutton, 1988; Hessel et al., 2018).
+CBOP (Jeong et al., 2023) constructs an explicit distribution of multi-step Q-values from thousands of model rollouts and uses this value as a target for training the Q-function.
+However, CBOP is computationally expensive for estimating a target value and uses multi-step returns solely for Q-learning, not for policy optimization.
+To tackle these issues in model-based offline RL, we introduce a simple yet effective model-based offline RL algorithm, Lower Expectile Q-learning (LEQ).
 As illustrated in Figure 1, LEQ uses expectile regression with a small $\tau$ for both policy and Q-function training, providing an efficient and elegant way to achieve conservative Q-value estimates.
-Moreover, to better handle long-horizon tasks, we propose to optimize a policy and Q-function using $\lambda$-returns (i.e.
-TD($\lambda$) targets) of long ($15$-step) model rollouts, allowing the policy to directly learn from low-bias multi-step returns [28].
-The experiments on the D4RL AntMaze and MuJoCo Gym tasks [6], as well as the NeoRL benchmark [26], demonstrate that our proposed conservative policy optimization with $\lambda$-return and critic training on offline data significantly improves offline RL policies in long-horizon tasks while achieving comparable performance in short-horizon, dense-reward tasks.
-Specifically, to the best of our knowledge, LEQ is the first model-based offline RL algorithm capable of matching or outperforming the performance of model-free offline RL algorithms on the long-horizon AntMaze tasks [6, 15].
+Moreover, we propose to optimize both policy and Q-function using $\lambda$-returns (i.e.
+TD($\lambda$) targets) of long ($10$-step) model rollouts, allowing the policy to directly learn from low-bias multi-step returns (Schulman et al., 2016).
+The experiments on the D4RL AntMaze, MuJoCo Gym (Fu et al., 2020), NeoRL (Qin et al., 2022), and V-D4RL (Lu et al., 2023) benchmarks show that LEQ improves model-based offline RL across diverse domains.
+To the best of our knowledge, LEQ is the first model-based offline RL algorithm capable of outperforming model-free offline RL algorithms on the long-horizon AntMaze tasks (Fu et al., 2020; Jiang et al., 2023).
+Moreover, LEQ matches the top scores across various benchmarks, while prior methods demonstrate superior performances only for a specific domain.
 Figure 1: Lower Expectile Q-learning (LEQ).
-(left) In offline model-based RL, an agent can generate imaginary trajectories using a world model.
-(right) For conservative Q-evaluation of the policy, LEQ learns the lower expectile of the target $Q$-distribution from a sampled individual rollout $\mathcal{T}_{i}$, without estimating the entire Q-distribution with exhaustive rollouts.
+(left) In model-based offline RL, an agent can generate imaginary trajectories using a world model.
+(right) For conservative Q-evaluation of the policy, LEQ learns the lower expectile of the target $Q$-distribution from a few sampled rollouts $\mathcal{T}_{i}$, without estimating the entire Q-distribution with exhaustive rollouts.
 2 Related Work
-Offline RL [21] aims to solve a reinforcement learning problem only with pre-collected datasets, better than behavioral cloning policies [25].
-One can simply apply off-policy RL algorithms on top of the fixed dataset.
-However, off-policy RL methods suffer from the overestimation of Q-values for actions unseen in the offline dataset [8, 18, 19], since an overestimated value function cannot get corrected through online environment interactions in offline RL.
-Model-free offline RL algorithms have addressed this value overestimation problem on out-of-distribution actions by (1) regularizing a policy to only output actions in the offline data [24, 17, 7] or (2) adopting a conservative value estimation for executing actions different from the dataset [19, 1].
+Offline RL (Levine et al., 2020) aims to solve an RL problem only with pre-collected datasets, outperforming behavioral cloning policies (Pomerleau, 1989).
+While it is possible to apply off-policy RL algorithms on fixed datasets, these algorithms suffer from the overestimation of Q-values for actions unseen in the offline dataset (Fujimoto et al., 2019; Kumar et al., 2019; 2020) since the overestimated values cannot be corrected through interactions with environments as in online RL.
+Model-free offline RL algorithms have addressed this value overestimation problem on out-of-distribution actions by (1) regularizing a policy to only output actions in the offline data (Peng et al., 2019; Kostrikov et al., 2022; Fujimoto and Gu, 2021) or (2) adopting a conservative value estimation for executing actions different from the dataset (Kumar et al., 2020; An et al., 2021).
 Despite their strong performances on the standard offline RL benchmarks, model-free offline RL policies tend to be constrained to the support of the data (i.e.
@@
 Model-based offline RL approaches have tried to overcome this limitation by suggesting a better use of the limited offline data – learning a world model and generating imaginary data with the learned model that covers out-of-distribution actions.
-Similar to Dyna-style online model-based RL [32, 9, 10, 11], an offline model-based RL policy can be trained on both offline data and model rollouts.
+Similar to Dyna-style online model-based RL (Sutton, 1991; Hafner et al., 2019; 2021; 2023), an offline model-based RL policy can be trained on both offline data and model rollouts.
 But, again, learned models may be inaccurate on states and actions outside the data support, making a policy easily exploit the learned models.
-Recent model-based offline RL algorithms have adopted the conservatism idea from model-free offline RL, penalizing policies incurring (1) uncertain transition dynamics [34, 16, 35] or (2) uncertain value estimation [30, 14].
-This conservative use of model-generated data enables model-based offline RL to outperform model-free offline RL in widely used offline RL benchmarks [30].
-However, uncertainty estimation is difficult and often inaccurate [35].
-Instead of relying on such heuristic [34, 16, 30] or expensive [14] uncertainty estimation, we propose to learn a conservative value function via expectile regression with a small $\tau$, which is simple, efficient, yet effective.
+Recent model-based offline RL algorithms have adopted the conservatism idea from model-free offline RL, penalizing policies incurring (1) uncertain transition dynamics (Yu et al., 2020; Kidambi et al., 2020; Yu et al., 2021) or (2) uncertain value estimation (Sun et al., 2023; Jeong et al., 2023).
+This conservative use of model-generated data enables model-based offline RL to outperform model-free offline RL in widely used offline RL benchmarks (Sun et al., 2023).
+However, uncertainty estimation is difficult and often inaccurate (Yu et al., 2021).
+Instead of relying on such heuristic (Yu et al., 2020; Kidambi et al., 2020; Sun et al., 2023) or expensive (Jeong et al., 2023) uncertainty estimation, we propose to learn a conservative value function via expectile regression with a small $\tau$, which is simple, efficient, yet effective, as illustrated in Figure 2.
+Expectile regression for offline RL has been first introduced by IQL (Kostrikov et al., 2022), which has been extended to model-based offline RL, such as IQL-TD-MPC (Chitnis et al., 2024).
+IQL uses “upper expectile” to approximate the “max operation” in $V(s)=\max_{a}Q(s,a)$ without querying out-of-distribution actions.
+On the other hand, our work fundamentally differs from IQL-like approaches in that our method uses “lower expectile” to get “conservative return estimates” from trajectories generated by potentially inaccurate model rollouts.
+Q-learning Lower Expectile Q-learning
+$\mathcal{L}_{Q}(\phi)=(Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})-Q_{t}^{\lambda}(\mathcal{T}))^{2}$Longrightarrow ${\color[rgb]{0.1875,0.39453125,0.6875}\definecolor[named]{pgfstrokecolor}{rgb}{0.1875,0.39453125,0.6875}\lvert\tau-\mathbbm{1}(Q_{t}^{\lambda}(\mathcal{T})>Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t}))\rvert}\cdot\mathcal{L}_{Q}(\phi)$
+$\nabla_{\theta}\mathcal{L}_{\pi}(\theta)=-\nabla_{\theta}\pi_{\theta}(\mathbf{s}_{t})\cdot\nabla_{\mathbf{a}_{t}}Q_{t}^{\lambda}(\mathcal{T})$Longrightarrow ${\color[rgb]{0.1875,0.39453125,0.6875}\definecolor[named]{pgfstrokecolor}{rgb}{0.1875,0.39453125,0.6875}\lvert\tau-\mathbbm{1}(Q_{t}^{\lambda}(\mathcal{T})>Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t}))\rvert}\cdot\nabla_{\theta}\mathcal{L}_{\pi}(\theta)$
+Figure 2: Comparison of standard Q-learning and Lower Expectile Q-learning (LEQ).
+LEQ generalizes standard Q-learning (with $\lambda$-returns $Q_{t}^{\lambda}(\mathcal{T})$) by multiplying a simple asymmetric weight $\lvert\tau-\mathbbm{1}(Q_{t}^{\lambda}(\mathcal{T})>Q_{\phi}(s_{t},a_{t}))\rvert$ to the Q-learning objectives.
+$\mathcal{T}=(\mathbf{s}_{0},\mathbf{a}_{0},r_{0},\mathbf{s}_{1},\mathbf{a}_{1},r_{1},\cdots,\mathbf{s}_{T})$ is a model-generated trajectory and $\tau\leq 0.5$ is the expectile hyperparameter controlling the degree of conservatism.
+When $\tau=0.5$, LEQ reduces to standard Q-learning.
 3 Preliminaries
@@
 […]
-In this paper, we consider the offline RL setup [21], where a policy $\pi$ is trained with a fixed given offline dataset, $\mathcal{D}_{\text{env}}=\{\tau_{1},\tau_{2},...,\tau_{N}\}$, without any additional online interactions.
+In this paper, we consider the offline RL setup (Levine et al., 2020), where a policy $\pi$ is trained with a fixed given offline dataset, $\mathcal{D}_{\text{env}}=\{\mathcal{T}_{1},\mathcal{T}_{2},...,\mathcal{T}_{N}\}$, without any additional online interactions.
 Model-based offline RL.
 As an offline RL policy is trained from a fixed dataset, one of the major challenges in offline RL is the limited data support; thus, lack of generalization to out-of-distribution states and actions.
-Model-based offline RL [16, 34, 35, 27, 30, 14] tackles this problem by augmenting the training data with imaginary training data (i.e.
-model rollouts) generated from the learned transition dynamics and reward model, $p_{\psi}(\mathbf{s}_{t+1},r\mid\mathbf{s}_{t},\mathbf{a}_{t})$.
-The typical process of model-based offline RL is as follows: (1) pretrain a model (or an ensemble of models) and an initial policy from the offline data, (2) generate short imaginary rollouts $\{\tau\}$ using the pretrained model and add them to the training dataset $\mathcal{D}_{\text{model}}\leftarrow\mathcal{D}_{\text{model}}\cup\{\tau\}$, (3) perform an offline RL algorithm on the augmented dataset $\mathcal{D}_{\text{model}}\cup\mathcal{D}_{\text{env}}$, and repeat (2) and (3).
+Model-based offline RL (Kidambi et al., 2020; Yu et al., 2020; 2021; Rigter et al., 2022; Sun et al., 2023; Jeong et al., 2023) tackles this problem by augmenting the training data with imaginary training data (i.e.
+model rollouts) generated from the learned transition dynamics and reward model, $p_{\psi}(\mathbf{s}_{t+1},r_{t}\mid\mathbf{s}_{t},\mathbf{a}_{t})$.
+The typical process of model-based offline RL is as follows: (1) pretrain a model (or an ensemble of models) and an initial policy from the offline data, (2) generate short imaginary rollouts $\{\mathcal{T}\}$ using the pretrained model and add them to the training dataset $\mathcal{D}_{\text{model}}\leftarrow\mathcal{D}_{\text{model}}\cup\{\mathcal{T}\}$, (3) perform an offline RL algorithm on the augmented dataset $\mathcal{D}_{\text{model}}\cup\mathcal{D}_{\text{env}}$, and repeat (2) and (3).
 Expectile regression.
 Expectile is a generalization of the expectation of a distribution $X$.
-While the expectation of $X$, $\mathbb{E}[X]$, can be viewed as a minimizer of the least-square objective, $L_{2}(y)=\mathbb{E}_{x\sim X}[(y-x)^{2}]$, $\tau$-expectile of $X$, $\mathbb{E}^{\tau}[X]$, can be defined as a minimizer of the asymmetric least-square objective:
-$L_{2}^{\tau}(y)=\mathbb{E}_{x\sim X}\left[\lvert\tau-\mathbbm{1}(y>x)\rvert\cdot(y-x)^{2}\right],$ (1)
-where $\lvert\tau-\mathbbm{1}(y>x)\rvert$ is an asymmetric weighting of $L_{2}$ and $0\leq\tau\leq 1$.
-We refer a $\tau$-expectile with $\tau<0.5$ as a lower expectile of $X$.
+While the expectation of $X$, $\mathbb{E}[X]$, can be viewed as a minimizer of the least-square objective, $\mathbb{E}_{x\sim X}[L_{2}(y-x)]=\mathbb{E}_{x\sim X}[\frac{1}{2}(y-x)^{2}]$, $\tau$-expectile of $X$, $\mathbb{E}^{\tau}[X]$, can be defined as a minimizer of the asymmetric least-square objective $\mathbb{E}_{x\sim X}[L_{2}^{\tau}(y-x)]$, where $L_{2}^{\tau}(\cdot)$ is defined as:
+$L_{2}^{\tau}(u)=\lvert\tau-\mathbbm{1}(u>0)\rvert\cdot u^{2},$ (1)
+where $\lvert\tau-\mathbbm{1}(u>0)\rvert$ is an asymmetric weighting of least-squared objective with $0\leq\tau\leq 1$.
+We refer to a $\tau$-expectile with $\tau<0.5$ as a lower expectile of $X$.
 When $\tau<0.5$, the objective assigns a high weight $1-\tau$ for smaller $x$ and a low weight $\tau$ for bigger $x$.
@@
 4 Approach
-The primary limitation for model-based offline RL in solving long-horizon tasks is inherent errors in a world model and critic outside the offline data support.
+The primary challenge of model-based offline RL is inherent errors in a world model and critic outside the support of offline data.
 Conservative value estimation can effectively handle such (falsely optimistic) errors.
-Prior approaches estimate conservative values through diverse uncertainty penalties; but they are either unreliable [35] or computationally expensive [14].
 In this paper, we introduce Lower Expectile Q-learning (LEQ), an efficient model-based offline RL method that achieves conservative value estimation via expectile regression of Q-values with lower expectiles when learning from model-generated data (Section 4.1).
-Additionally, we address the noisy value estimation problem in long-horizon tasks [23] using $\lambda$-returns on $10$-step imaginary rollouts (Section 4.2).
+Additionally, we address the noisy value estimation problem (Park et al., 2024) using $\lambda$-returns on $10$-step imaginary rollouts (Section 4.2).
 Finally, we train a deterministic policy conservatively by maximizing the lower expectile of $\lambda$-returns (Section 4.3).
@@
 2:  Pretrain $\{p_{\psi_{1}},\cdots,p_{\psi_{M}}\}$ on $\mathcal{D}_{\text{env}}$triangleright $\mathcal{L}_{\text{wm}}(\psi)=-\mathbb{E}_{(\mathbf{s},\mathbf{a},r,\mathbf{s}^{\prime})\in\mathcal{D}_{\text{env}}}\log p_{\psi}(\mathbf{s}^{\prime},r\mid\mathbf{s},\mathbf{a})$
-3:  Pretrain $\pi_{\theta}$ and $Q_{\phi}$ on $\mathcal{D}_{\text{env}}$triangleright using BC for $\pi_{\theta}$ and FQE [20] for $Q_{\phi}$
+3:  Pretrain $\pi_{\theta}$ and $Q_{\phi}$ on $\mathcal{D}_{\text{env}}$triangleright using BC for $\pi_{\theta}$ and FQE (Le et al., 2019) for $Q_{\phi}$
 4:  $\mathcal{D}_{\text{model}}\leftarrow\emptyset$
@@
 10:        $\mathbf{a}_{t}=\pi_{\theta}(\mathbf{s}_{t})$
-11:        $\mathbf{s}_{t+1},r_{t}\sim p_{\psi}(\cdot\mid\mathbf{s}_{t},\mathbf{a}_{t})$, where $p_{\psi}\sim\{p_{\psi_{1}},\cdots,p_{\psi_{M}}\}$
-12:     // Generate imaginary data, $\tau=\{(\mathbf{s}_{0},\mathbf{a}_{0},r_{0},\cdots,\mathbf{s}_{H-1},\mathbf{a}_{H-1},r_{H-1},\mathbf{s}_{H})_{i}\}$
+11:        $\mathbf{s}_{t+1},r_{t}\sim p_{\psi}(\cdot\mid\mathbf{s}_{t},\mathbf{a}_{t})$, where $p_{\psi}\sim\{p_{\psi_{1}},\cdots,p_{\psi_{M}}\}$triangleright sample $p_{\psi}$ every step
+12:     // Generate imaginary data, $\mathcal{T}=\{(\mathbf{s}_{0},\mathbf{a}_{0},r_{0},\cdots,\mathbf{s}_{H-1},\mathbf{a}_{H-1},r_{H-1},\mathbf{s}_{H})_{i}\}$
 13:     $\mathbf{s}_{0}\sim\mathcal{D}_{\text{model}}$triangleright start imaginary rollout from any state in $\mathcal{D}_{\text{model}}$
@@
 15:        $\mathbf{a}_{t}=\pi_{\theta}(\mathbf{s}_{t})$
-16:        $\mathbf{s}_{t+1},r_{t}\sim p_{\psi}(\cdot\mid\mathbf{s}_{t},\mathbf{a}_{t})$, where $p_{\psi}\sim\{p_{\psi_{1}},\cdots,p_{\psi_{M}}\}$
-17:     // Update critic using both offline and model-generated data
-18:     Update critic $Q_{\phi}$ to minimize $\mathcal{L}^{\lambda}_{Q}(\phi)$ in Eq. 7 using $\tau$ and $\{\mathbf{s},\mathbf{a},r,\mathbf{s}^{\prime}\}\sim\mathcal{D}_{\text{env}}$
+16:        $\mathbf{s}_{t+1},r_{t}\sim p_{\psi}(\cdot\mid\mathbf{s}_{t},\mathbf{a}_{t})$, where $p_{\psi}\sim\{p_{\psi_{1}},\cdots,p_{\psi_{M}}\}$triangleright sample $p_{\psi}$ every step
+17:     // Update critic using both model-generated data and offline data
+18:     Update critic $Q_{\phi}$ to minimize $\mathcal{L}^{\lambda}_{Q}(\phi)$ in Eq. 7 using $\mathcal{T}$ and $\{\mathbf{s},\mathbf{a},r,\mathbf{s}^{\prime}\}\sim\mathcal{D}_{\text{env}}$
 19:     // Update actor using only model-generated data
-20:     Update actor $\pi_{\theta}$ to minimize ^$\hat{\mathcal{L}}^{\lambda}_{\pi}(\theta)$ in Eq. 12 using $\tau$
+20:     Update actor $\pi_{\theta}$ to minimize ^$\hat{\mathcal{L}}^{\lambda}_{\pi}(\theta)$ in Eq. 11 using $\mathcal{T}$
 4.1 Lower expectile Q-learning
@@
 In this paper, we propose Lower Expectile Q-learning (LEQ), which learns a conservative Q-function via expectile regression with small $\tau$, avoiding unreliable uncertainty estimation and exhaustive Q-value estimation.
-As illustrated in Figure 1, the target value for $Q_{\phi}(\mathbf{s},\mathbf{a})$, where $\mathbf{a}\leftarrow\pi_{\theta}(\mathbf{s})$, can be estimated by rolling out an ensemble of world models and averaging $r(\mathbf{s},\mathbf{a})+\gamma Q_{\phi}(\mathbf{s}^{\prime},\mathbf{a}^{\prime})$ over all possible $\mathbf{s}^{\prime}$:
+As illustrated in Figure 1, the target value for $Q_{\phi}(\mathbf{s},\mathbf{a})$, where $\mathbf{a}\leftarrow\pi_{\theta}(\mathbf{s})$, can be estimated by rolling out an ensemble of world models and averaging $r+\gamma Q_{\phi}(\mathbf{s}^{\prime},\mathbf{a}^{\prime})$ over all possible $\mathbf{s}^{\prime}$:
 $\hat{y}_{\text{model}}=\mathbb{E}_{\psi\sim\{\psi_{1},...,\psi_{M}\}}\mathbb{E}_{(\mathbf{s}^{\prime},r)\sim p_{\psi}(\cdot\mid\mathbf{s},\mathbf{a})}\left[r+\gamma Q_{\phi}(\mathbf{s}^{\prime},\pi_{\theta}(\mathbf{s}^{\prime}))\right].$ (2)
 This target value has three error sources: the predicted future state and reward $\mathbf{s}^{\prime},r\sim p_{\psi}(\cdot\mid\mathbf{s},\mathbf{a})$ and future Q-value $Q_{\phi}(\mathbf{s}^{\prime},\pi_{\theta}(\mathbf{s}^{\prime}))$.
-Thus, the target value computed from model-generate data, $\hat{y}_{\text{model}}$, is more prone to overestimation than that of the original target Q-value, $\hat{y}_{\text{env}}$, computed from $(\mathbf{s},\mathbf{a},r,\mathbf{s}^{\prime})\sim D_{\text{env}}$:
+Thus, the target value from model-generated data, $\hat{y}_{\text{model}}$, is more prone to overestimation than the original target Q-value, $\hat{y}_{\text{env}}$, computed from $(\mathbf{s},\mathbf{a},r,\mathbf{s}^{\prime})\sim D_{\text{env}}$:
 $\hat{y}_{\text{env}}=r+\gamma Q_{\phi}(\mathbf{s}^{\prime},\pi_{\theta}(\mathbf{s}^{\prime})).$ (3)
-To mitigate the overestimation problem in estimating the true Q-value from $H$-step inaccurate world model rollouts, we propose to use expectile regression on target Q-value estimation with small $\tau$; as illustrated in Figure 1, expectile regression with small $\tau$ tends to choose the target Q-value that is lower than the expectation, effectively providing a conservative estimate of target Q-value.
-Another advantage of using expectile regression is that we do not have to exhaustively evaluate Q-values to get $\tau$-expectiles as Jeong et al.
-[14]; instead, we can do conservative estimation using sampling:
-$L_{Q,\text{model}}(\phi)=\mathbb{E}_{\mathbf{s}_{0}\in\mathcal{D}_{\text{model}},\tau\sim p_{\psi},\pi_{\theta}}\left[\frac{1}{H}\sum_{t=0}^{H}L_{2}^{\tau}(Q_{\phi}(\mathbf{s}_{t},\pi_{\theta}(\mathbf{s}_{t}))-\hat{y}_{\text{model}})\right].$ (4)
-In addition, the Q-function is also trained on the offline data $\mathcal{D}_{\text{env}}$ with the standard Bellman update:
+To mitigate the overestimation of $\hat{y}_{\text{model}}$ from inaccurate $H$-step model rollouts, we propose to use lower expectile regression on target Q-value estimates with small $\tau$.
+As illustrated in Figure 2, expectile regression with small $\tau$ learns a Q-function predicting Q-values lower than the expectation, i.e., a conservative estimate of a target Q-value.
+Another advantage of using lower expectile regression is that we do not have to exhaustively evaluate Q-values to get $\tau$-expectiles as Jeong et al.
+(2023); instead, we can learn a conservative Q-function with sampling:
+$L_{Q,\text{model}}(\phi)=\mathbb{E}_{\mathbf{s}_{0}\in\mathcal{D}_{\text{model}},\mathcal{T}\sim p_{\psi},\pi_{\theta}}\left[\frac{1}{H}\sum_{t=0}^{H}L_{2}^{\tau}(Q_{\phi}(\mathbf{s}_{t},\pi_{\theta}(\mathbf{s}_{t}))-\hat{y}_{\text{model}})\right].$ (4)
+Additionally, we train the Q-function with the transitions in the dataset $\mathcal{D}_{\text{env}}$, which do not have the risk of overestimation caused by the inaccurate model, using the standard Bellman update:
 $\mathcal{L}_{Q,\text{env}}(\phi)=\mathbb{E}_{(\mathbf{s},\mathbf{a},r,\mathbf{s}^{\prime})\in\mathcal{D}_{\text{env}}}\left[\frac{1}{2}(Q_{\phi}(\mathbf{s},\mathbf{a})-\hat{y}_{\text{env}})^{2}\right].$ (5)
-To stabilize training of the Q-function, we adopt EMA regularization [11], which prevents drastic change of Q-values by regularizing the difference between the Q-predictions and ones from the exponential moving average:
-$\mathcal{L}_{Q,\text{EMA}}(\phi)=\mathbb{E}_{(\mathbf{s},\mathbf{a})\in\mathcal{D}_{\text{env}}}\left[(Q_{\phi}(\mathbf{s},\mathbf{a})-Q_{\bar{\phi}}(\mathbf{s},\mathbf{a}))^{2}\right],$ (6)
-where $\bar{\phi}$ is an exponential moving average of $\phi$. Note that by using EMA regularization, we do not use the target Q-network for Equations 2 and 3.
-Finally, by combining the three losses above, we define the critic loss as follows:
-$\mathcal{L}_{Q}(\phi)=\beta\mathcal{L}_{Q,\text{model}}(\phi)+(1-\beta)\mathcal{L}_{Q,\text{env}}(\phi)+\beta_{\text{EMA}}\mathcal{L}_{Q,\text{EMA}}(\phi).$ (7)
+To stabilize training of the Q-function, we adopt EMA regularization (Hafner et al., 2023), which prevents drastic change of Q-values by regularizing the difference between the predictions from $Q_{\phi}$ and ones from its exponential moving average $Q_{\bar{\phi}}$:
+$\mathcal{L}_{Q,\text{EMA}}(\phi)=\mathbb{E}_{(\mathbf{s},\mathbf{a})\in\mathcal{D}_{\text{env}}}\left[(Q_{\phi}(\mathbf{s},\mathbf{a})-Q_{\bar{\phi}}(\mathbf{s},\mathbf{a}))^{2}\right].$ (6)
+Finally, by combining the three aforementioned losses, we define the critic loss as follows:
+$\mathcal{L}_{Q}(\phi)=\beta\mathcal{L}_{Q,\text{model}}(\phi)+(1-\beta)\mathcal{L}_{Q,\text{env}}(\phi)+\omega_{\text{EMA}}\mathcal{L}_{Q,\text{EMA}}(\phi).$ (7)
 4.2 Lower expectile Q-learning with $\lambda$-return
-To further improve LEQ for long-horizon tasks, we use $\lambda$-return instead of $1$-step return for Q-learning.
-$\lambda$-return allows a Q-function and policy to learn from low-bias multi-step returns [28].
-Reducing bias in value estimation with $\lambda$-return is especially important on long-horizon tasks where values for nearby states are similar to each other, as illustrated in Figure 2.
-We first define $\lambda$-return of a trajectory $\tau$ in timestep $t$, $Q_{t}^{\lambda}(\tau)$, using $N$-step return, $G_{t:t+N}(\tau)$: [^2] Our $\lambda$-return is slightly different from [31, 11] that puts a high weight to the last $N$-step return, $G_{t:H}(\tau)$.
-$G_{t:t+N}(\tau)=\sum_{i=0}^{N-1}\gamma^{i}r(\mathbf{s}_{t+i},\mathbf{a}_{t+i})+\gamma^{N}Q_{\phi}(\mathbf{s}_{t+N},\mathbf{a}_{t+N}),$ (8)
-$Q_{t}^{\lambda}(\tau)=\frac{1-\lambda}{1-\lambda^{H-t-1}}\sum_{i=1}^{H-t}\lambda^{i-1}G_{t:t+i}(\tau).$ (9)
-Then, we can rewrite the Q-learning loss in Equation 4 with $1$-step return to the one with $\lambda$-return:
-$\mathcal{L}^{\lambda}_{Q,\text{model}}(\phi)=\mathbb{E}_{\mathbf{s}_{0}\in\mathcal{D}_{\text{model}},\tau\sim p_{\psi},\pi_{\theta}}\left[\sum_{t=0}^{H-1}L_{2}^{\tau}(Q_{\phi}(\mathbf{s}_{t},\pi_{\theta}(\mathbf{s}_{t}))-Q^{\lambda}_{t}(\tau))\right].$ (10)
-Figure 2: $\lambda$-return with noisy critic.
-For long-horizon tasks, value estimates (blue for high values and white for low values) are likely to be flat for the states far from the reward signal shown with the yellow star.
-Since Q-values are small and similar to each other in this region, Q-values are easily dominated by noise, leading to a wrong learning signal for the policy, as illustrated in the middle row.
-$\lambda$-return can effectively reduce the critic noise by considering both short-term and long-term value estimates, leading to a better learning signal for the policy.
+To further improve LEQ, we use $\lambda$-return instead of $1$-step return for Q-learning.
+$\lambda$-return allows a Q-function and policy to learn from low-bias multi-step returns (Schulman et al., 2016).
+Using $N$-step returns $G_{t:t+N}(\mathcal{T})=\sum_{i=0}^{N-1}\gamma^{i}r_{t+i}+\gamma^{N}Q_{\phi}(\mathbf{s}_{t+N},\mathbf{a}_{t+N})$, we define $\lambda$-return of an $H$-step trajectory $\mathcal{T}$ in timestep $t$, $Q_{t}^{\lambda}(\mathcal{T})$ as: [^2] Our $\lambda$-return slightly differs from (Sutton, 1988) that puts a high weight to the last $N$-step return, $G_{t:H}(\mathcal{T})$.
+$Q_{t}^{\lambda}(\mathcal{T})=\frac{1-\lambda}{1-\lambda^{H-t-1}}\sum_{i=1}^{H-t}\lambda^{i-1}G_{t:t+i}(\mathcal{T}).$ (8)
+Then, we can rewrite the Q-learning loss in Equation 4 with $\lambda$-return targets:
+$\mathcal{L}^{\lambda}_{Q,\text{model}}(\phi)=\mathbb{E}_{\mathbf{s}_{0}\in\mathcal{D}_{\text{model}},\mathcal{T}\sim p_{\psi},\pi_{\theta}}\left[\sum_{t=0}^{H-1}L_{2}^{\tau}(Q_{\phi}(\mathbf{s}_{t},\pi_{\theta}(\mathbf{s}_{t}))-Q^{\lambda}_{t}(\mathcal{T}))\right].$ (9)
 4.3 Lower expectile policy learning with $\lambda$-return
-For policy optimization, we can use a deterministic policy $\mathbf{a}=\pi_{\theta}(\mathbf{s})$ and update the policy using the deterministic policy gradients similar to DDPG [22].
-[^3] LEQ also works with a stochastic policy; but, a deterministic policy is easier to train in an offline setup.
-Instead of maximizing the immediate Q-value, $Q_{\phi}(\mathbf{s},\mathbf{a})$, we propose to directly maximize the lower expectile of $\lambda$-return, which is a more accurate learning target for a policy, analogous to the conservative critic target in Section 4.2:
-$\mathcal{L}_{\pi}^{\lambda}(\theta)=-\mathbb{E}_{\mathbf{s}_{0}\in\mathcal{D}_{\text{model}},\tau\sim p_{\psi},\pi_{\theta}}\left[\sum_{t=0}^{H}\mathbb{E}^{\tau}_{\tau\sim p_{\psi},\pi_{\theta}}\left[Q_{t}^{\lambda}(\tau)\right]\right].$ (11)
-However, due to the expectile term in Equation 11, computing the gradient of $\mathcal{L}_{\pi}^{\lambda}(\theta)$ is not trivial.
-To estimate this gradient, we propose a differentiable surrogate loss, approximating $\mathbb{E}^{\tau}_{\tau\sim p_{\psi},\pi_{\theta}}\left[Q^{\lambda}_{t}(\tau)\right]$ in Equation 11 with $Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})$:
-^$\hat{\mathcal{L}}_{\pi}^{\lambda}(\theta)=-\mathbb{E}_{\mathbf{s}_{0}\in\mathcal{D}_{\text{model}},\tau\sim p_{\psi},\pi_{\theta}}\left[\sum_{t=0}^{H}\lvert\tau-\mathbbm{1}\left(Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})>Q^{\lambda}_{t}(\tau)\right)\rvert\cdot Q_{t}^{\lambda}(\tau)\right].$ (12)
-Intuitively, this surrogate loss sets a higher weight ($1-\tau$) on a conservative $\lambda$-return estimation (i.e.
-$Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})>Q^{\lambda}_{t}(\tau)$), encouraging a policy to optimize for this conservative $\lambda$-return.
-On the other hand, an optimistic $\lambda$-return estimation (i.e.
-$Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})<Q^{\lambda}_{t}(\tau)$) has a less impact to the policy with a smaller weight ($\tau$).
-Thus, optimizing this surrogate loss leads to the policy maximizing the lower expectile of $\lambda$-return.
-We provide a proof in Appendix B saying that the proposed surrogate loss is a better approximation of Equation 11 than directly maximizing Q-values, $Q_{\phi}(\mathbf{s},\mathbf{a})$.
+For policy optimization, we use a deterministic policy $\mathbf{a}=\pi_{\theta}(\mathbf{s})$ and update the policy using the deterministic policy gradients similar to DDPG (Lillicrap et al., 2016).
+[^3] LEQ also works with a stochastic policy; but, a deterministic policy is sufficient for our experiments.
+To provide more accurate learning targets for the policy, instead of maximizing the immediate Q-value, $Q_{\phi}(\mathbf{s},\mathbf{a})$, we maximize the lower expectile of $\lambda$-return, analogous to our conservative critic learning in Section 4.2:
+$\mathcal{L}_{\pi}^{\lambda}(\theta)=-\mathbb{E}_{\mathbf{s}_{0}\in\mathcal{D}_{\text{model}}}\left[\sum_{t=0}^{H}\mathbb{E}^{\tau}_{\mathcal{T}\sim p_{\psi},\pi_{\theta}}\left[Q_{t}^{\lambda}(\mathcal{T})\right]\right].$ (10)
+However, because of the expectile term, $\mathbb{E}^{\tau}[Q_{t}^{\lambda}]$, we cannot directly compute the gradient of $\mathcal{L}_{\pi}^{\lambda}(\theta)$.
+To change the expectile to expectation, we use the relationship $\mathbb{E}^{\tau}[Q_{t}^{\lambda}]=\frac{\mathbb{E}\left[\lvert\tau-\mathbbm{1}(\mathbb{E}^{\tau}[Q_{t}^{\lambda}]>Q_{t}^{\lambda})\rvert\cdot Q_{t}^{\lambda}\right]}{\mathbb{E}\left[\lvert\tau-\mathbbm{1}(\mathbb{E}^{\tau}[Q_{t}^{\lambda}]>Q_{t}^{\lambda})\rvert\right]}$, and optimize the unnormalized version, $\mathbb{E}\left[\lvert\tau-\mathbbm{1}(\mathbb{E}^{\tau}[Q_{t}^{\lambda}]>Q_{t}^{\lambda})\rvert\cdot Q_{t}^{\lambda}\right]$.
+By approximating $\mathbb{E}^{\tau}[Q_{t}^{\lambda}]$ with the learned Q-estimator $Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})$, we derive a differentiable surrogate loss of Equation 10:
+^$\hat{\mathcal{L}}_{\pi}^{\lambda}(\theta)=-\mathbb{E}_{\mathbf{s}_{0}\in\mathcal{D}_{\text{model}},\mathcal{T}\sim p_{\psi},\pi_{\theta}}\left[\sum_{t=0}^{H}\lvert\tau-\mathbbm{1}\left(Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})>Q^{\lambda}_{t}(\mathcal{T})\right)\rvert\cdot Q_{t}^{\lambda}(\mathcal{T})\right].$ (11)
+Intuitively, this surrogate loss sets a higher weight ($1-\tau$) on a conservative $\lambda$-return estimates (i.e.
+$Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})>Q^{\lambda}_{t}(\mathcal{T})$), encouraging a policy to optimize for this conservative $\lambda$-return.
+On the other hand, an optimistic $\lambda$-return estimates (i.e.
+$Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})<Q^{\lambda}_{t}(\mathcal{T})$) has a less impact to the policy with a smaller weight ($\tau$).
+We provide a proof in Appendix B showing that the proposed surrogate loss provides a better approximation of Equation 10 than the immediate Q-value, $Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})$.
 4.4 Expanding dataset with model rollouts
-One of the problem of offline RL is that data distribution is limited to the offline dataset $\mathcal{D}_{\text{env}}$.
-To tackle this problem, we can simulate the current policy inside the model and use the generated trajectories to expand the dataset, similar to prior works [35, 30].
-However, the state coverage will be identical when the policy converges, which might lead to catastrophic forgetting.
-To prevent this issue, we expand the dataset using the model rollouts from a noisy exploration policy.
-Specifically, we execute the exploration policy $\pi_{\exp}(\cdot\mid\mathbf{s})$, which simply adds noise $\epsilon\sim N(0,\sigma_{\text{exp}}^{2})$ to the current policy $\pi_{\theta}(\mathbf{s})$, and generate a trajectory of length $R$ ($R=5$ in this paper).
-We refer this expanded dataset as $\mathcal{D}_{\text{model}}$.
-Note that we do not use off-policy actions and rewards for critic and policy updates; instead, we generate $H$-step model rollouts starting from $\mathbf{s}\sim\mathcal{D}_{\text{model}}$ and use them for training the policy and Q-function.
-Thus, we need to store only the states from the rollouts.
+One of the problems of offline RL is that data distribution is limited to the offline dataset $\mathcal{D}_{\text{env}}$.
+To tackle this problem, we expand the dataset using simulated trajectories, which we refer to as $\mathcal{D}_{\text{model}}$ (Yu et al., 2021).
+To diversify the simulated trajectories in $\mathcal{D}_{\text{model}}$, we use a noisy exploration policy, which adds a noise $\epsilon\sim N(0,\sigma_{\text{exp}}^{2})$, to the current policy and generate a trajectory of length $R$.
 5 Experiments
-In this paper, we propose a novel model-based offline RL method with simple and efficient yet accurate conservative value estimation.
-Through our experiments, we aim to answer the following questions: (1) Can LEQ solve long-horizon tasks?
-(2) How does LEQ perform in widely used offline RL benchmarks?
-(3) Which component enables model-based offline RL to learn the AntMaze tasks?
-(a) Hopper
-(b) Walker2d
-(c) HalfCheetah
+In this paper, we propose a novel model-based offline RL method with simple, efficient, yet accurate conservative value estimation.
+Through our experiments, we aim to answer the following questions: (1) Can LEQ solve diverse domains of problems, including both dense-reward tasks and long-horizon sparse-reward tasks?
+(2) Can LEQ be applied to pixel-based environments?
+(3) How individual components of LEQ affect the performance?
+(a) Umaze
+(b) Medium
+(c) Large
+(d) Ultra
+Figure 3: AntMaze tasks.
+(e) Hopper
+(f) Walker2d
+(g) HalfCheetah
 Figure 4: Locomotion tasks.
-(d) Umaze
-(e) Medium
-(f) Large
-(g) Ultra
-Figure 4: AntMaze tasks.
 5.1 Tasks
-To show the strength of LEQ in solving long-horizon tasks, we use the AntMaze tasks, which aims to navigate a $8$-DOF ant robot to the desired goal position, as shown in Figure 4.
-Specifically, we use umaze, medium, large datasets from D4RL [6], and ultra dataset from Jiang et al.
-[15].
-Moreover, we evaluate our method on MuJoCo locomotion tasks (Figure 4) with dense rewards with D4RL [6] and NeoRL [26] datasets.
+To verify the strength of our low-bias model-based conservative value estimation in diverse domains, we test LEQ on four benchmarks: D4RL AntMaze, D4RL MuJoCo Gym (Fu et al., 2020), NeoRL (Qin et al., 2022), and V-D4RL (Lu et al., 2023).
+We first test on long-horizon AntMaze tasks: umaze, medium, large from D4RL, and ultra from Jiang et al.
+(2023), as shown in Figure 4.
+We also evaluate LEQ on locomotion tasks (Figure 4): state-based tasks from D4RL, NeoRL and pixel-based tasks from V-D4RL.
 Please refer to Appendix A for more experimental details.
 5.2 Compared offline RL algorithms
-We compare the performance of LEQ with the state-of-the-art offline RL algorithms. Please note that LEQ uses the same hyperparameters across all tasks, except the expectile parameter, $\tau$.
 Model-free offline RL.
-We consider behavioral cloning (BC[25]; TD3+BC [7], which combines BC loss to TD3; CQL [19], which penalizes the actions out of data distribution; and IQL [17], which utilizes expectile regression to estimate the value function.
-For locomotion tasks, we also compare with EDAC [1], which penalizes the Q-values according to the uncertainty of Q-functions
+We consider behavioral cloning (BC(Pomerleau, 1989); TD3+BC (Fujimoto and Gu, 2021), which combines BC loss to TD3; CQL (Kumar et al., 2020), which penalizes out-of-distribution actions; and IQL (Kostrikov et al., 2022), which uses upper-expectile regression to estimate the value function.
+For locomotion tasks, we also compare with EDAC (An et al., 2021), which penalizes Q-values based on its uncertainty.
+For pixel-based tasks, we compare with DrQ+BC (Lu et al., 2023), which combines BC loss to DrQ-v2 (Yarats et al., 2022); ACRO, which learns representations with a multi-step inverse dynamics model.
 Model-based offline RL.
-We consider MOPO [34] and MOBILE [30], which penalize Q-values according to the transition uncertainty and the bellman uncertainty of a world model, respectively; COMBO [35], which combines CQL with MBPO; RAMBO [27], which trains an adversarial world model against the policy; and CBOP [14], which utilizes multi-step returns for critic updates.
-5.3 Results on long-horizon AntMaze tasks
-As shown in Table 1, LEQ significantly outperforms the prior model-based approaches for all $8$ datasets.
-LEQ achieves $58.6$ and $60.2$ for antmaze-large-play and antmaze-large-diverse, while the second best method, RAMBO [27], scores only $0.0$ and $2.4$, respectively.
-We believe these performance gains come from our conservative value estimation, which works more stable than the uncertainty-based penalization of prior works.
-Moreover, LEQ even significantly outperforms the model-free approaches in antmaze-umaze, antmaze-large, and antmaze-ultra.
-Despite its superior performance, LEQ often shows high variance during training, resulting in worse performance on antmaze-medium.
-Over the course of training, LEQ mostly achieves high success rates, but the evaluation results sometimes drops to $0\%$ as shown in Appendix, Figure 5.
-We leave the problem of reducing the high variance of our method in certain environments as a future work.
+We consider MOPO (Yu et al., 2020) and MOBILE (Sun et al., 2023), which penalize Q-values according to the transition uncertainty and the Bellman uncertainty of a world model, respectively; COMBO (Yu et al., 2021), which combines CQL with MBPO; RAMBO (Rigter et al., 2022), which trains an adversarial world model against the policy; and CBOP (Jeong et al., 2023), which utilizes multi-step returns for critic updates; IQL-TD-MPC (Chitnis et al., 2024), which extends TD-MPC (Hansen et al., 2022) to offline setting with IQL.
+For pixel-based environments, we consider OfflineDV2 (Lu et al., 2023), which penalizes Q-values according to the dynamics errors, and ROSMO (Liu et al., 2023), which uses one-step model rollouts for policy improvement.
+LEQ follows MOBILE for most implementation details but implemented in JAX (Bradbury et al., 2018), which makes it $6$ times faster than the PyTorch versions of MOBILE and CBOP.
+Please refer to Table 8 for detailed comparison.
+Sequence modeling for offline RL.
+We consider TT (Janner et al., 2021), which trains a Transformer model (Vaswani, 2017) to predict offline trajectories and applies beam search to find the best trajectory; and TAP (Jiang et al., 2023), which improves TT by quantizing the action space with VQ-VAE (Van Den Oord et al., 2017).
 Table 1: AntMaze results.
@@
 The results for other methods are reported following their respective papers.
-Model-free Model-based
-Dataset BC TD3+BC CQL IQL MOPO COMBO RAMBO MOBILE CBOP LEQ (ours)
-antmaze-umaze $65.0$ $78.6$ $74.0$ $87.5$ $0.0$ $80.3$ $25.0$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $\mathbf{94.4}$$\mathbf{\pm 6.3}$
-antmaze-umaze-diverse $55.0$ $71.4$ $84.0$ $62.2$ $0.0$ $57.3$ $0.0$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $\mathbf{71.0}$$\mathbf{\pm 12.3}$
-antmaze-medium-play $0.0$ $3.0$ $61.2$ $\mathbf{71.2}$ $0.0$ $0.0$ $16.4$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $58.8$$\pm 33.0$
-antmaze-medium-diverse $0.0$ $10.6$ $53.7$ $\mathbf{70.0}$ $0.0$ $0.0$ $23.2$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $46.2$$\pm 23.2$
-antmaze-large-play $0.0$ $0.0$ $15.8$ $39.6$ $0.0$ $0.0$ $0.0$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $\mathbf{58.6}$$\mathbf{\pm 9.1}$
-antmaze-large-diverse $0.0$ $0.2$ $14.9$ $47.5$ $0.0$ $0.0$ $2.4$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $\mathbf{60.2}$$\mathbf{\pm 18.3}$
-antmaze-ultra-play -- -- -- $8.3$- -- -- $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $\mathbf{25.8}$$\mathbf{\pm 18.2}$
-antmaze-ultra-diverse -- -- -- $15.6$- -- -- $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $\mathbf{55.8}$$\mathbf{\pm 18.3}$
-Total w/o antmaze-ultra $120.0$ $163.8$ $303.6$ $354.1$ $0.0$ $137.6$ $67.0$ $0.0$ $0.0$ $\mathbf{388.8}$
-Total -- -- -- $378.0$- -- -- $0.0$ $0.0$ $\mathbf{470.4}$
+Model-free Seq. modeling Model-based
+Dataset CQL IQL TT TAP COMBO RAMBO MOBILE CBOP IQL-TD-MPC LEQ (ours)
+umaze $74.0$ $87.5$ $\mathbf{100.0}$ $81.5$ $80.3$ $25.0$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $52.0$ $94.4$$\pm 6.3$
+umaze-diverse $\mathbf{84.0}$ $62.2$ $21.5$ $68.5$ $57.3$ $0.0$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $72.6$ $71.0$$\pm 12.3$
+medium-play $61.2$ $71.2$ $\mathbf{93.3}$ $78.0$ $0.0$ $16.4$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $\mathbf{88.8}$ $58.8$$\pm 33.0$
+medium-diverse $53.7$ $\mathbf{70.0}$ $\mathbf{100.0}$ $85.0$ $0.0$ $23.2$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $40.3$ $46.2$$\pm 23.2$
+large-play $15.8$ $39.6$ $66.7$ $\mathbf{74.0}$ $0.0$ $0.0$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $66.6$ $58.6$$\pm 9.1$
+large-diverse $14.9$ $47.5$ $60.0$ $\mathbf{82.0}$ $0.0$ $2.4$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $4.0$ $60.2$$\pm 18.3$
+ultra-play -- $8.3$ $20.0$ $22.0$- -- $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $20.6$ $\mathbf{25.8}$$\mathbf{\pm 18.2}$
+ultra-diverse -- $15.6$ $33.3$ $26.0$- -- $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $3.6$ $\mathbf{55.8}$$\mathbf{\pm 18.3}$
+Total w/o ultra $303.6$ $354.1$ $441.5$ $\mathbf{469.0}$ $137.6$ $67.0$ $0.0$ $0.0$ $324.3$ $388.8$
+Total -- $378.0$ $\mathbf{494.8}$ $\mathbf{517.0}$- -- $0.0$ $0.0$ $348.5$ $470.4$
 […]
-5.4 Results on MuJoCo Gym locomotion tasks
-For D4RL MuJoCo Gym tasks in Table 3, LEQ achieves comparable results with the best score of prior works in $6$ out of $12$ tasks.
-Furthermore, in Table 2, LEQ outperforms most of the prior works in the NeoRL benchmark, especially in the Hopper and Walker2d domains.
-These results show that LEQ serves as a general offline RL algorithm, not limited to long-horizon tasks.
-Similar to antmaze-medium, LEQ also suffers from the high variance problem.
-During training, LEQ often achieves high performance, but then, suddenly falls back to 00, as shown in Appendix, Figure 5.
-This is mainly because the learned models sometimes fail to capture failures (e.g.
-hopper and walker falling off) and predict an optimistic future (e.g.
-hopper and walker walking forward).
 Table 2: NeoRL results.
@@
 The results for prior works are reported following Sun et al.
-[30] and Qin et al.
-[26].
+(2023) and Qin et al.
+(2022).
 MOPO is an improved version of MOPO presented in Sun et al.
-[30].
-We highlight the results that are better than $95\%$ of the best score.
+(2023).
+We highlight the results better than $95\%$ of the best score.
 Model-free Model-based
@@
 Total $408.8$ $490.1$ $504.6$ $456.6$ $385.0$ $256.5$ $\mathbf{546.5}$ $\mathbf{572.4}$
+5.3 Results on long-horizon AntMaze tasks
+As shown in Table 1, LEQ significantly outperforms the prior model-based approaches.
+For example, LEQ achieves $60.2$ and $55.8$ for large-diverse and ultra-diverse, while the second best method, IQL-TD-MPC (Chitnis et al., 2024), scores only $4.0$ and $3.6$, respectively.
+We believe these performance gains come from our conservative value estimation, which works more stable than the uncertainty-based penalization of prior works.
+Moreover, LEQ even significantly outperforms the model-free approaches in umaze, large, and ultra mazes, and outperforms sequence modeling methods, TT and TAP, which serve as strong baselines for AntMaze tasks, in the most challenging ultra mazes, showing the advantage of utilizing low-bias multi-step return on long-horizon tasks.
+Figure 5: Failure in medium mazes. The agent plans impossible trajectories on certain states (red circles).
+Despite its superior performance, LEQ shows high variance in the performance on antmaze-medium.
+We found that the medium mazes include states separated by walls that are very close to each other (denoted as red circles in Figure 5), such that all of the learned world models falsely believe the agent can pass through the walls.
+This incorrect prediction makes the agent to plan faster, but impossible trajectories as shown in Figure 5.
+We believe that this could be addressed by employing improved world models or increasing the number of ensembles for the world models.
 Table 3: D4RL MuJoCo Gym results.
-Each number is a normalized score averaged over $100$ trials [6].
+Each number is a normalized score averaged over $100$ trials (Fu et al., 2020).
 Our results are averaged over $5$ seeds.
@@
 MOPO is an improved version of MOPO, introduced in Sun et al.
-[30].
+(2023).
 We highlight the results that are better than $95\%$ of the best score.
+Model-free Seq. modeling Model-based
+Dataset CQL EDAC IQL TT TAP MOPO COMBO RAMBO MOBILE CBOP LEQ (ours)
+hopper-r $5.3$ $25.3$ $7.6$ $6.9$ $31.7$ $17.9$ $25.4$ $\mathbf{31.9}$ $\mathbf{32.8}$ $\mathbf{32.4}$$\pm 0.3$
+hopper-m $61.9$ $\mathbf{101.6}$ $66.3$ $67.4$ $63.4$ $62.8$ $97.2$ $87.0$ $\mathbf{106.6}$ $\mathbf{102.6}$ $\mathbf{103.4}$$\pm 0.3$
+hopper-mr $86.3$ $\mathbf{101.0}$ $94.7$ $99.4$ $87.3$ $\mathbf{99.4}$ $\mathbf{103.5}$ $89.5$ $99.5$ $\mathbf{104.3}$ $\mathbf{103.9}$$\pm 1.3$
+hopper-me $96.9$ $\mathbf{110.7}$ $91.5$ $\mathbf{110.0}$ $105.5$ $81.6$ $\mathbf{111.1}$ $88.2$ $\mathbf{112.6}$ $\mathbf{111.6}$ $\mathbf{109.4}$$\pm 1.8$
+walker2d-r $5.4$ $16.6$ $5.2$ $5.9$ $7.4$ $7.0$ $0.0$ $17.9$ $17.8$ $\mathbf{21.5}$$\pm 0.1$
+walker2d-m $79.5$ $\mathbf{92.5}$ $78.3$ $84.9$ $64.9$ $81.3$ $84.1$ $81.9$ $84.9$ $\mathbf{87.7}$ $74.9$$\pm 26.9$
+walker2d-mr $76.8$ $87.1$ $73.9$ $89.2$ $66.8$ $85.6$ $56.0$ $89.2$ $89.9$ $92.7$ $\mathbf{98.7}$$\pm 6.0$
+walker2d-me $109.1$ $\mathbf{114.7}$ $109.6$ $101.9$ $107.4$ $\mathbf{112.9}$ $103.3$ $56.7$ $\mathbf{115.2}$ $\mathbf{117.2}$ $108.2$$\pm 1.3$
+halfcheetah-r $31.3$ $28.4$ $11.8$ $6.1$ $\mathbf{38.5}$ $\mathbf{38.8}$ $\mathbf{39.5}$ $\mathbf{39.3}$ $32.8$ $30.8$$\pm 3.3$
+halfcheetah-m $46.9$ $65.9$ $47.4$ $46.9$ $45.0$ $73.0$ $54.2$ $\mathbf{77.9}$ $\mathbf{74.6}$ $\mathbf{74.3}$ $71.7$$\pm 4.4$
+halfcheetah-mr $45.3$ $61.3$ $44.2$ $44.1$ $40.8$ $\mathbf{72.1}$ $55.1$ $\mathbf{68.7}$ $\mathbf{71.7}$ $66.4$ $65.5$$\pm 1.1$
+halfcheetah-me $95.0$ $\mathbf{106.3}$ $86.7$ $95.0$ $91.8$ $90.8$ $90.0$ $95.4$ $\mathbf{108.2}$ $\mathbf{105.4}$ $\mathbf{102.8}$$\pm 0.4$
+Total $739.7$ $911.4$ $717.2$ $747.5$ $844.0$ $802.0$ $812.4$ $\mathbf{959.5}$ $\mathbf{953.4}$ $\mathbf{923.2}$
+5.4 Results on MuJoCo Gym locomotion tasks
+For the NeoRL benchmark in Table 2, LEQ outperforms most of the prior works, especially in the Hopper and Walker2d domains.
+Furthermore, for D4RL MuJoCo Gym tasks in Table 3, LEQ achieves comparable results with the best score of prior works in $7$ out of $12$ tasks, These results show that LEQ serves as a general offline RL algorithm, widely applicable to various domains.
+Figure 6: High variance during training. Our algorithm experiences oscillation due to optimistic imaginations near the initial states.
+Table 4: V-D4RL results.
+We report the mean and standard deviation of returns over $3$ seeds.
+For ROSMO, we replace the categorical distribution in their official code with Gaussian distribution to support continuous action spaces.
+For MOPO, we implement MOPO on top of DreamerV3.
+The results for other prior works are reported following Lu et al.
+(2023); Islam et al.
+(2023).
 Model-free Model-based
-Dataset BC TD3+BC CQL EDAC IQL MOPO COMBO RAMBO MOBILE CBOP LEQ (ours)
-hopper-r $3.7$ $8.5$ $5.3$ $25.3$ $7.6$ $31.7$ $17.9$ $25.4$ $\mathbf{31.9}$ $\mathbf{32.8}$ $\mathbf{32.4}$$\pm 0.3$
-hopper-m $54.1$ $59.3$ $61.9$ $\mathbf{101.6}$ $66.3$ $62.8$ $97.2$ $87.0$ $\mathbf{106.6}$ $\mathbf{102.6}$ $\mathbf{103.4}$$\pm 0.3$
-hopper-mr $16.6$ $60.9$ $86.3$ $\mathbf{101.0}$ $94.7$ $\mathbf{99.4}$ $\mathbf{103.5}$ $89.5$ $99.5$ $\mathbf{104.3}$ $\mathbf{103.9}$$\pm 1.3$
-hopper-me $53.9$ $98.0$ $96.9$ $\mathbf{110.7}$ $91.5$ $81.6$ $\mathbf{111.1}$ $88.2$ $\mathbf{112.6}$ $\mathbf{111.6}$ $\mathbf{109.4}$$\pm 1.8$
-walker2d-r $1.3$ $1.6$ $5.4$ $16.6$ $5.2$ $7.4$ $7.0$ $0.0$ $17.9$ $17.8$ $\mathbf{21.5}$$\pm 0.1$
-walker2d-m $70.9$ $83.7$ $79.5$ $\mathbf{92.5}$ $78.3$ $81.3$ $84.1$ $81.9$ $84.9$ $\mathbf{87.7}$ $74.9$$\pm 26.9$
-walker2d-mr $20.3$ $81.8$ $76.8$ $87.1$ $73.9$ $85.6$ $56.0$ $89.2$ $89.9$ $92.7$ $\mathbf{98.7}$$\pm 6.0$
-walker2d-me $90.1$ $110.1$ $109.1$ $\mathbf{114.7}$ $109.6$ $\mathbf{112.9}$ $103.3$ $56.7$ $\mathbf{115.2}$ $\mathbf{117.2}$ $108.2$$\pm 1.3$
-halfcheetah-r $2.2$ $11.0$ $31.3$ $28.4$ $11.8$ $\mathbf{38.5}$ $\mathbf{38.8}$ $\mathbf{39.5}$ $\mathbf{39.3}$ $32.8$ $30.8$$\pm 3.3$
-halfcheetah-m $43.2$ $48.3$ $46.9$ $65.9$ $47.4$ $73.0$ $54.2$ $\mathbf{77.9}$ $\mathbf{74.6}$ $\mathbf{74.3}$ $71.7$$\pm 4.4$
-halfcheetah-mr $37.6$ $44.6$ $45.3$ $61.3$ $44.2$ $\mathbf{72.1}$ $55.1$ $\mathbf{68.7}$ $\mathbf{71.7}$ $66.4$ $65.5$$\pm 1.1$
-halfcheetah-me $44.0$ $90.7$ $95.0$ $\mathbf{106.3}$ $86.7$ $90.8$ $90.0$ $95.4$ $\mathbf{108.2}$ $\mathbf{105.4}$ $\mathbf{102.8}$$\pm 0.4$
-Total $437.9$ $698.5$ $739.7$ $911.4$ $717.2$ $844.0$ $802.0$ $812.4$ $\mathbf{959.5}$ $\mathbf{953.4}$ $\mathbf{923.2}$
+Method BC CQL DrQ+BC ACRO OfflineDV2 MOPO ROSMO LEQ
+walker_walk-random $2.0$ $14.4$ $5.5$ $0.0$ $\mathbf{28.7}$ $3.2$$\pm 0.4$ $2.9$$\pm 0.4$ $22.4$$\pm 1.1$
+walker_walk-medium $40.9$ $14.8$ $\mathbf{46.8}$ $\mathbf{48.7}$ $34.1$ $37.1$$\pm 3.7$ $\mathbf{49.8}$$\pm 2.3$ $43.1$$\pm 3.2$
+walker_walk-medium_replay $16.5$ $11.4$ $28.7$ $27.8$ $\mathbf{56.5}$ $11.4$$\pm 9.2$ $28.1$$\pm 0.9$ $43.0$$\pm 7.3$
+walker_walk-medium_expert $47.7$ $56.4$ $\mathbf{86.4}$ $\mathbf{91.4}$ $43.9$ $46.6$$\pm 3.5$ $82.9$$\pm 0.7$ $\mathbf{87.2}$$\pm 2.4$
+cheetah_run-random $0.0$ $5.9$ $5.8$ $0.0$ $\mathbf{31.7}$ $3.2$$\pm 3.9$ $2.5$$\pm 0.5$ $14.8$$\pm 1.0$
+cheetah_run-medium $\mathbf{51.6}$ $40.9$ $\mathbf{53.0}$ $\mathbf{52.8}$ $17.2$ $34.7$$\pm 6.6$ $\mathbf{50.8}$$\pm 2.5$ $37.9$$\pm 8.0$
+cheetah_run-medium_replay $25.0$ $10.7$ $44.8$ $41.7$ $\mathbf{61.5}$ $30.3$$\pm 5.1$ $48.6$$\pm 1.5$ $34.3$$\pm 1.1$
+cheetah_run-medium_expert $\mathbf{57.5}$ $20.9$ $50.6$ $46.6$ $10.4$ $36.6$$\pm 5.6$ $45.4$$\pm 2.2$ $25.1$$\pm 6.6$
+Total $241.2$ $175.4$ $\mathbf{321.6}$ $\mathbf{309.0}$ $284.0$ $203.3$ $\mathbf{310.9}$ $\mathbf{307.6}$
+Similar to antmaze-medium, LEQ experiences high variance in MuJoCo tasks.
+During training, LEQ often achieves high performance, but then, suddenly falls back to 00, as shown in Figure 6.
+This is mainly because the learned models sometimes fail to capture failures (e.g.
+hopper and walker falling off) and predict an optimistic future (e.g.
+hopper and walker walking forward).
+5.5 Results with visual inputs
+As shown in Table 4, LEQ combined with DreamerV3 (Hafner et al., 2023) performs on par with the state-of-the-art methods on V-D4RL datasets, demonstrating its scalability to visual observations.
+Notably, LEQ achieves the highest score on the walker_walk-medium_expert dataset among model-based methods, where OfflineDV2 and MOPO struggles.
+LEQ also outperforms model-free approaches on random datasets and walker_walk-medium_replay dataset, highlighting the strength of model-based methods in datasets with diverse state-action distributions.
 5.6 Ablation studies
-To understand why LEQ (LEQ-$\lambda$) works well in long-horizon tasks, we conduct ablation studies and answer to the following four questions: (1) Does using $\lambda$-return help?
-(2) Is LEQ better than prior uncertainty-based penalization methods?
+To deeply understand how LEQ (LEQ-$\lambda$) work, we conduct ablation studies in AntMaze environments and answer to the following five questions: (1) Is lower expectile Q-learning better than prior uncertainty-based penalization methods?
+(2) Does lower expectile policy learning better than existing policy learning methods?
+(3) Does $\mathbf{\lambda}$-return help?
 (4) Which factor enables LEQ to work in AntMaze?
-and (4) How do imagination length $H$ and data expansion length $R$ affect the performance?
+and (5) How do imagination length $H$ and data expansion affect the performance?
+(1) Lower expectile Q-learning.
+We compare our lower expectile Q-learning with the conservative value estimator in MOBILE (Sun et al., 2023), which penalizes Q-values based on the standard deviation of Q-ensemble networks.
+In Table 5, replacing lower expectile Q-learning with MOBILE decreases the success rate, both with $\lambda$-returns ($461.8$ vs $338.9$) and without them ($373.5$ vs $339.0$).
+(2) Lower expectile policy learning.
+We also compare our lower expectile policy learning with AWR (Peng et al., 2019) and directly maximizing $Q(s,a)$.
+As shown in Table 5, LEQ shows better performance compared to AWR ($461.8$ vs $114.2$) and maximizing the $Q$-values ($461.8$ vs $410.3$).
 (3) $\lambda$-returns.
-To verify the effect of $\lambda$-return, we compare our method (LEQ-$\lambda$) with the versions with $1$-step return (LEQ-$1$) and $H$-step return (LEQ-$H$).
-Table 4 shows that using $\lambda$-return drastically improves the performance on AntMaze compared to using $1$-step return or $H$-step return.
-This result is coherent with the observations in prior online RL methods [28, 11].
-(2) Lower expectile Q-learning.
-We compare our lower expectile Q-learning with another conservative value estimator, MOBIP used in MOBILE [30], which penalizes Q-values with the standard deviation of Q-ensemble networks.
-The only difference between LEQ and MOBIP is their target Q-value computation for both critic and policy updates.
-Table 4 shows that using MOBIP not only deteriorates the success rates (in MOBIP-$1$) but also does not benefit from $\lambda$-return (in MOBIP-$\lambda$).
-(3) What makes offline model-based RL work in AntMaze?
-Prior to LEQ, none of offline model-based RL methods work in AntMaze, whereas our method even outperforms model-free methods.
-Thus, we investigate which changes in LEQ enable offline model-based RL work in AntMaze.
-Table 4: Ablation study results on the AntMaze tasks.
-(1) We compare different Q-targets, LEQ-$\lambda$, LEQ-$1$, and LEQ-$H$.
-(2) We compare our lower expectile Q-learning strategy with another conservative Q-value estimation, MOBIP-$1$ and MOBIP-$\lambda$.
-(3) MOBILE is our re-implementation of MOBILE using MOBILE’s default hyperparameters, i.e., $\beta=0.95$, $\gamma=0.99$, and $R=5$.
-We note that lowering $\beta$ to $0.25$ is crucial for MOBILE to achieve meaningful scores in AntMaze.
-Dataset umaze medium large ultra Total
-umaze diverse play diverse play diverse play diverse
-LEQ-$\lambda$ (ours) $\mathbf{94.4}$$\pm 6.3$ $\mathbf{71.0}$$\pm 12.3$ $58.8$$\pm 33.0$ $46.2$$\pm 23.2$ $\mathbf{58.6}$$\pm 9.1$ $\mathbf{60.2}$$\pm 18.3$ $25.8$$\pm 18.2$ $\mathbf{55.8}$$\pm 18.3$ $\mathbf{470.4}$
-LEQ-$H$ $\mathbf{93.0}$$\pm 3.4$ $60.7$$\pm 10.4$ $46.3$$\pm 32.4$ $0.0$ ±$\pm 0.0$ $\mathbf{57.0}$$\pm 25.6$ $33.3$$\pm 43.0$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $290.3$
-LEQ-$1$ $89.6$$\pm 4.8$ $37.0$$\pm 32.8$ $55.8$$\pm 28.7$ $29.8$$\pm 24.5$ $34.2$$\pm 13.4$ $49.3$$\pm 9.0$ $\mathbf{42.2}$$\pm 13.2$ $35.6$$\pm 13.0$ $373.5$
-MOBIP-$\lambda$ $84.3$$\pm 3.5$ $40.3$$\pm 20.4$ $51.3$$\pm 9.0$ $39.7$$\pm 12.5$ $28.3$$\pm 21.5$ $33.7$$\pm 10.0$ $38.0$$\pm 27.1$ $23.3$$\pm 4.9$ $338.9$
-MOBIP-1 $59.5$$\pm 3.5$ $46.5$$\pm 1.5$ $57.0$$\pm 11.0$ $\mathbf{54.0}$$\pm 9.0$ $23.5$$\pm 19.5$ $38.5$$\pm 1.5$ $39.5$$\pm 11.5$ $20.5$$\pm 20.5$ $339.0$
-MOBILE $1.0$$\pm 2.0$ $0.0$ ±$\pm 0.0$ $6.4$$\pm 5.5$ $5.0$$\pm 5.0$ $0.8$$\pm 1.6$ $0.8$$\pm 1.2$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $14.0$
-MOBILE ($\mathbf{\beta}$ = 0.25) $77.0$$\pm 6.4$ $20.4$$\pm 15.7$ $\mathbf{64.6}$$\pm 11.1$ $31.6$$\pm 16.9$ $2.6$$\pm 2.8$ $7.2$$\pm 8.9$ $4.6$$\pm 3.0$ $5.0$$\pm 4.6$ $213.0$
-MOBILE ($\mathbf{\gamma}$ = 0.997) $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $7.2$$\pm 4.1$ $1.6$$\pm 2.1$ $9.6$$\pm 7.1$ $5.4$$\pm 4.9$ $0.0$ ±$\pm 0.0$ $1.8$$\pm 2.7$ $25.6$
-MOBILE ($R$ = 10) $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $5.0$$\pm 5.1$ $0.6$$\pm 1.2$ $7.4$$\pm 14.8$ $1.6$$\pm 3.2$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $14.6$
-We first re-implement MOBILE with some technical tricks used in LEQ: LayerNorm [3], SymLog [11], single Q-network, and no target Q-value clipping; but, MOBILE achieves a barely non-zero score, $14.0$.
-We found that the key to make MOBILE work is reducing $\beta$, the ratio for the loss calculated from imaginary rollouts and from dataset transitions.
-When we lower $\beta$ from $0.95$ to $0.25$ (used in LEQ), MOBILE shows meaningful performances in umaze and medium mazes, and achieves $213.0$ in total.
-We suggest that utilizing the true transition from the dataset is important in long-horizon tasks, which was undervalued in prior works.
-(4) Imagination length $H$ and dataset expansion length $R$.
+The first three rows of Table 5 show the effect of $\lambda$-returns in LEQ.
+Substituting $\lambda$-returns with $H$-step returns ($461.8$ vs $290.3$) or $1$-step returns ($461.8$ vs $373.5$) significantly decreases the performance.
+Moreover, while LEQ shows better performance than MOBILE without $\lambda$-returns ($373.5$ vs $339.0$), the performance of LEQ gets significantly better with $\lambda$-returns, compared to MOBILE with $\lambda$-returns ($461.8$ vs $338.9$).
+Table 5: Impact of lower expectile Q-learning and $\lambda$-returns on AntMaze.
+We ablate the effects of lower expectile and $\lambda$-returns on the critic and policy updates in LEQ.
+The design choices from LEQ are colored in blue and other options are colored in red.
+The results are averaged over $5$ seeds.
+Design choices umaze medium large ultra Total
+conservatism critic update policy update umaze diverse play diverse play diverse play diverse
+Lower expectile $\lambda$-returns $\lambda$-returns $\mathbf{94.4}$$\pm 6.3$ $\mathbf{71.0}$$\pm 12.3$ $50.2$$\pm 39.9$ $46.2$$\pm 23.2$ $\mathbf{58.6}$$\pm 9.1$ $\mathbf{60.2}$$\pm 18.3$ $25.8$$\pm 18.2$ $\mathbf{55.8}$$\pm 18.3$ $\mathbf{461.8}$
+Lower expectile $H$-step $H$-step $\mathbf{93.0}$$\pm 3.4$ $60.7$$\pm 10.4$ $46.3$$\pm 32.4$ $0.0$ ±$\pm 0.0$ $\mathbf{57.0}$$\pm 25.6$ $33.3$$\pm 43.0$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $290.3$
+Lower expectile $1$-step $Q(\mathbf{s},\mathbf{a})$ $89.6$$\pm 4.8$ $37.0$$\pm 32.8$ $55.8$$\pm 28.7$ $29.8$$\pm 24.5$ $34.2$$\pm 13.4$ $49.3$$\pm 9.0$ $\mathbf{42.2}$$\pm 13.2$ $35.6$$\pm 13.0$ $373.5$
+Lower expectile $\lambda$-returns $Q(\mathbf{s},\mathbf{a})$ $81.0$$\pm 10.5$ $46.2$$\pm 16.8$ $\mathbf{61.8}$$\pm 12.4$ $40.6$$\pm 11.4$ $39.2$$\pm 12.5$ $40.5$$\pm 11.7$ $\mathbf{42.8}$$\pm 21.8$ $47.5$$\pm 5.9$ $410.3$
+Lower expectile $\lambda$-returns AWR $69.2$$\pm 7.5$ $44.4$$\pm 18.4$ $0.6$$\pm 0.6$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $114.2$
+MOBILE $\lambda$-returns $\lambda$-returns $84.3$$\pm 3.5$ $40.3$$\pm 20.4$ $51.3$$\pm 9.0$ $39.7$$\pm 12.5$ $28.3$$\pm 21.5$ $33.7$$\pm 10.0$ $38.0$$\pm 27.1$ $23.3$$\pm 4.9$ $338.9$
+MOBILE $1$-step $Q(\mathbf{s},\mathbf{a})$ $59.5$$\pm 3.5$ $46.5$$\pm 1.5$ $57.0$$\pm 11.0$ $\mathbf{54.0}$$\pm 9.0$ $23.5$$\pm 19.5$ $38.5$$\pm 1.5$ $39.5$$\pm 11.5$ $20.5$$\pm 20.5$ $339.0$
+(4) What makes model-based offline RL work in AntMaze?
+LEQ shows outstanding performance compared to previous offline model-based RL methods, especially in large and ultra mazes.
+To understand which aspects of LEQ enabled this success, we applied its changes to MOBILE and analyzed the impact.
+The results are detailed in Table 6.
+Table 6: Impact of hyperparameters in MOBILE on AntMaze.
+MOBILE uses the hyperparameters from MOBILE: $\beta=0.95$, $\gamma=0.99$, and $R=5$, whereas LEQ uses $\beta=0.25$, $\gamma=0.997$, and $R=10$.
+The results show that $\beta$ is the most critical hyperparameter that makes MOBILE work in AntMaze.
+  Hyperparams. umaze medium large ultra Total
+$\beta$ $\gamma$ $R$ umaze diverse play diverse play diverse play diverse
+$0.25$ $0.997$ $10$ $53.8$$\pm 26.8$ $\mathbf{22.5}$$\pm 22.2$ $54.0$$\pm 5.8$ $\mathbf{49.5}$$\pm 6.2$ $\mathbf{28.3}$$\pm 6.0$ $\mathbf{28.0}$$\pm 11.4$ $\mathbf{25.5}$$\pm 6.9$ $\mathbf{23.8}$$\pm 15.8$ $\mathbf{285.3}$
+$0.25$ $0.997$ $5$ $\mathbf{74.0}$$\pm 6.9$ $3.7$$\pm 2.6$ $54.7$$\pm 27.9$ $28.0$$\pm 9.6$ $18.7$$\pm 18.6$ $8.0$$\pm 9.3$ $9.7$$\pm 8.2$ $9.0$$\pm 3.7$ $205.7$
+$0.25$ $0.99$ $10$ $39.7$$\pm 23.4$ $5.0$$\pm 7.1$ $39.3$$\pm 27.9$ $38.0$$\pm 15.0$ $0.0$ ±$\pm 0.0$ $3.7$$\pm 5.2$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $125.7$
+$0.25$ $0.99$ $5$ $\mathbf{77.0}$$\pm 6.4$ $20.4$$\pm 15.7$ $\mathbf{64.6}$$\pm 11.1$ $31.6$$\pm 16.9$ $2.6$$\pm 2.8$ $7.2$$\pm 8.9$ $4.6$$\pm 3.0$ $5.0$$\pm 4.6$ $213.0$
+$0.95$ $0.997$ $10$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $1.8$$\pm 3.0$ $0.5$$\pm 0.9$ $0.2$$\pm 0.4$ $2.2$ ±$\pm 2.3$ $1.0$$\pm 1.7$ $0.0$ ±$\pm 0.0$ $5.7$
+$0.95$ $0.997$ $5$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $7.2$$\pm 4.1$ $1.6$$\pm 2.1$ $9.6$$\pm 7.1$ $5.4$$\pm 4.9$ $0.0$ ±$\pm 0.0$ $1.8$$\pm 2.7$ $25.6$
+$0.95$ $0.99$ $10$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $5.0$$\pm 5.1$ $0.6$$\pm 1.2$ $7.4$$\pm 14.8$ $1.6$$\pm 3.2$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $14.6$
+$0.95$ $0.99$ $5$ $1.0$$\pm 2.0$ $0.0$ ±$\pm 0.0$ $6.4$$\pm 5.5$ $5.0$$\pm 5.0$ $0.8$$\pm 1.6$ $0.8$$\pm 1.2$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $14.0$
+We first re-implement MOBILE with some technical tricks used in LEQ (denoted as MOBILE): LayerNorm (Ba et al., 2016), SymLog (Hafner et al., 2023), single Q-network, and no target Q-value clipping.
+However, MOBILE still achieves a barely non-zero score, $14.0$.
+We found that reducing the ratio $\beta$ between the losses from imaginary rollouts and dataset transitions is key to make MOBILE work (i.e.
+achieving meaningful performances in umaze and medium mazes, with a total score of $213.0$).
+This adjustment also allows for a higher discount rate and longer imagination horizon, yielding the best results for MOBILE.
+We suggest that utilizing the true transition from the dataset is important in long-horizon tasks, which has been undervalued in prior works.
+We provide additional extensive ablation results on LEQ in Appendix D.
+(5) Imagination length $H$ and dataset expansion.
 As shown in Table 7, the performance increases when it goes to $H=10$ from $H=5$, but it drops when $H=15$.
-This result shows the trade-off of using the world model: the further the agent imagines, more the agent becomes robust to the error of the critic, but more it becomes prone to the error from the model prediction.
-We also evaluate LEQ without the dataset expansion ($H=10,R=1$).
+This result shows the trade-off of using the world model – the further the agent imagines, the more the agent becomes robust to the error of the critic, but the more it becomes prone to the error from the model prediction.
+We also evaluate LEQ without the dataset expansion.
 In AntMaze, the results with and without the dataset expansion are similar, as shown in Table 7.
-On the other hand, the dataset expansion makes the policy more stable and better in the D4RL MuJoCo tasks (Table 13).
-Table 5: LEQ with different imagination length $H$ and data expansion length $R$.
+On the other hand, the dataset expansion makes the policy more stable and better in the D4RL MuJoCo tasks (in Appendix, Table 20).
+Table 7: LEQ with different imagination length $H$ and data expansion.
 A longer $H$ can mitigate critic biases, while increasing model errors, which leads to poor performance.
 Each number is averaged over $5$ random seeds.
-Dataset $\mathbf{H=10,R=5}$ (ours) $H=5,R=5$ $H=15,R=5$ $H=10,R=1$
+Dataset $\mathbf{H=10}$ (ours) $H=5$ $H=15$ w/o dataset expansion
 antmaze-umaze $\mathbf{94.4}$$\pm 6.3$ $\mathbf{95.2}$$\pm 1.7$ $\mathbf{98.6}$$\pm 0.5$ $\mathbf{97.4}$$\pm 1.4$
@@
 Total $\mathbf{470.4}$ $404.8$ $371.0$ $435.4$
+6 Limitations
+Following prior work on model-based offline RL (Sun et al., 2023; Jeong et al., 2023), we assume access to the ground-truth termination function of a task, different from online model-based RL approaches, which learn a termination function from interactions.
+As shown in Table 22, Using a learned termination function instead of the ground-truth termination function results in a significant performance drop ($461.8\rightarrow 232.6$), particularly in diverse datasets ($233.2\rightarrow 66.8$), where termination signals are limited because the dataset is collected by navigating to randomly selected goals.
+While relying on the ground-truth termination function simplifies the problem, it limits the applicability of the method to scenarios where this information is readily available.
+Extending the proposed approach to learn terminal signals from the dataset would be an immediate next step.
 7 Conclusion
 In this paper, we propose a novel offline model-based reinforcement learning method, LEQ, which uses expectile regression to get a conservative evaluation of a policy from model-generated trajectories.
-Expectile regression eases the pain of constructing the whole distribution of Q-targets and allows for estimating the conservative value via sampling.
+Expectile regression eases the pain of constructing the whole distribution of Q-targets and allows for learning a conservative Q-function via sampling.
 Combined with $\lambda$-returns in both critic and policy updates for the imaginary rollouts, the policy can receive learning signals that are more robust to both model errors and critic errors.
-We empirically show that LEQ improves the performance in various tasks – especially, achieving the state-of-the-art performance in the long-horizon AntMaze tasks.
-6.1 Limitations
-Following prior work on model-based offline RL [30, 14], we assume access to the ground-truth termination function of a task, different from online model-based RL approaches, which learn a termination function from interactions.
-However, since this termination function is conditioned on a state, a model requires to plan on a state space (or an observation space), which could be challenging in a high-dimensional state space (e.g.
-pixel observations).
-Extending the proposed approach to complex environments with high-dimensional observations would be an immediate next step.
-6.2 Broader Impacts
+We empirically show that LEQ robustly improves the performance of model-based approaches in various domains, including state-based locomotion, long-horizon navigation, and visual control tasks.
+Acknowledgments
+We would like to thank Junik Bae for helpful discussion.
+This work was supported in part by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant (RS-2020-II201361, Artificial Intelligence Graduate School Program (Yonsei University)) and the National Research Foundation of Korea (NRF) grant (RS-2024-00333634) funded by the Korean Government (MSIT).
+Kwanyoung Park was supported by the Electronics and Telecommunications Research Institute (ETRI) grant (24ZR1100) and IITP grant (RS-2024-00509279, Global AI Frontier Lab) funded by the Korea government (MSIT).
+Ethics Statement
 Our method aims to increase the ability of autonomous agents, such as robots and self-driving cars, to learn from static, offline data without interacting with the world.
 This enables autonomous agents to utilize data with diverse qualities (not necessarily from experts).
-We believe that this paper does not have any immediate negative societal impact.
-Acknowledgments and Disclosure of Funding
-We would like to thank Junik Bae for helpful discussion.
-This work was supported in part by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant (RS-2020-II201361, Artificial Intelligence Graduate School Program (Yonsei University)) and the National Research Foundation of Korea (NRF) grant (RS-2024-00333634) funded by the Korean Government (MSIT).
-Kwanyoung Park was supported by Electronics and Telecommunications Research Institute (ETRI).
+We believe that this paper does not have any immediate negative ethical concerns.
+Reproducibility Statement
+To ensure the reproducibility of our work, we provide the full code of LEQ in the supplementary materials, along with instructions to replicate the experiments presented in the paper.
+We provide the experimental details in Appendix A and the proof on the derivation of our surrogate policy objective in Appendix B.
 Appendix A Training Details
 Computing resources.
-All experiments are done on a single RTX 4090 GPU and $4$ AMD EPYC 9354 CPU cores.
-We use $5$ different random seeds for each experiment and report the mean and standard deviation.
-Each offline RL experiment takes $2$ hours for ours, $12$ hours for MOBILE, and $24$ hours for CBOP.
+All experiments are done on a single RTX 4090 GPU and $8$ AMD EPYC 9354 CPU cores.
+For state-based environments, we use $5$ different random seeds for each experiment and report the mean and standard deviation, while each offline RL experiment takes $2$ hours for ours, $12$ hours for MOBILE, and $24$ hours for CBOP.
+For pixel-based environments, we use $3$ different random seeds and each experiment takes $2$ hours for ours and $8$ hours for ROSMO.
 Environment details.
-For the locomotion tasks, we use the dataset provided by D4RL [6] and NeoRL [26].
-Following IQL [17], we normalize rewards using the maximum and minimum return of all trajectories.
-We use the true termination functions of the environments, implemented in MOBILE [30].
-For the AntMaze tasks, we use the dataset provided by D4RL [6].
-Following IQL [17], we subtract $1$ from the rewards in the datasets so that the agent receives $-1$ for each step and 00 on termination.
+For state-based locomotion tasks, we use the datasets provided by D4RL [Fu et al., 2020] and NeoRL [Qin et al., 2022].
+Following IQL [Kostrikov et al., 2022], we normalize rewards using the maximum and minimum returns of all trajectories.
+We use the true termination functions of the environments, implemented in MOBILE [Sun et al., 2023].
+For pixel-based environments, we do not normalize the rewards.
+For AntMaze tasks, we use the datasets provided by D4RL [Fu et al., 2020].
+Following IQL [Kostrikov et al., 2022], we subtract $1$ from the rewards in the datasets so that the agent receives $-1$ for each step and 00 on termination.
 We use the true termination functions of the environments.
 The termination functions of the AntMaze tasks are not deterministic because a goal of a maze is randomized every time the environment is reset.
-Nevertheless, we follow the implementation of CBOP [14], where the termination region is set to a circle around the mean of the goal distribution with the radius $0.5$.
-Method implementation details.
+Nevertheless, we follow the implementation of CBOP [Jeong et al., 2023], where the termination region is set to a circle around the mean of the goal distribution with the radius $0.5$.
+Implementation details of compared methods.
 For all compared methods, we use the results from their corresponding papers when available.
-For IQL [17], we run the official implementation with $5$ seeds to reproduce the results for the random datasets in D4RL and NeoRL.
+For IQL [Kostrikov et al., 2022], we run the official implementation with $5$ seeds to reproduce the results for the random datasets in D4RL and NeoRL.
 For the AntMaze tasks, we run the official implementation of MOBILE and CBOP with $5$ random seeds.
 Please note that the original MOBILE implementation does not use the true termination function, so we replace it with our termination function.
-For MOPO, COMBO, and RAMBO, we use the results reported in RAMBO [27].
+For MOPO, COMBO, and RAMBO, we use the results reported in RAMBO [Rigter et al., 2022].
+For DMControl tasks, we replace the categorical distribution of the policy with gaussian distribution of the official ROSMO codebase, and run the experiments with sampling hyperparameter $N=4$.
 World models.
-For training world models, we use the architecture and training script from OfflineRL-Kit [29], matching the implementation of MOBILE [30].
+For state-based environments, we use the architecture and training script from OfflineRL-Kit [Sun, 2023], matching the implementation of MOBILE [Sun et al., 2023].
 Each world model is implemented as a $4$-layer MLPs with the hidden layer size of $200$.
@@
 training in total $35$ world models and using $25$ models), which takes approximately $5$ hours in average.
+For pixel-based environments, we use the 12M model of DreamerV3 [Hafner et al., 2023] and pretrain the world model using its loss function.
+We follow the implementation of OfflineDV2 [Lu et al., 2023], training 7 ensemble of world models for stochastic latent prediction, which takes $4$ hours in average.
+We select one model for each step, following the design choice of MOBILE, as we found that randomly choosing a model every step can make imaginary rollouts more robust to model biases, leading to better performance.
 Policy and critic networks.
-We use $3$-layer MLPs with size of $256$ both for the policy network and the critic network.
-We use layer normalization [3] to prevent catastrophic over/underestimation [4], and squash the state inputs using symlog to keep training stable from outliers in long-horizon model rollouts [11].
-Pretraining.
-For some environments, we found that a randomly initialized policy can lead to abnormal rewards or transition prediction from the world models in the early stage, leading to unstable training.
-Following CBOP [14], we pretrain a policy $\pi_{\theta}$ and a critic $Q_{\phi}$ using behavioral cloning and FQE [20], respectively.
+For state-based environments, we use $3$-layer MLPs with size of $256$ both for the policy network and the critic network.
+We use layer normalization [Ba et al., 2016] to prevent catastrophic over/underestimation [Ball et al., 2023], and squash the state inputs using symlog to keep training stable from outliers in long-horizon model rollouts [Hafner et al., 2023].
+For pixel-based environments, we use the architecture of the 12M model in DreamerV3 models.
+Pretraining policy and critic networks.
+For some environments, we found that a randomly initialized policy can lead to abnormal rewards or transition prediction from the world models in the early stage, leading to unstable training [Jelley et al., 2024].
+Following CBOP [Jeong et al., 2023], we pretrain a policy $\pi_{\theta}$ and a critic $Q_{\phi}$ using behavioral cloning and FQE [Le et al., 2019], respectively for state-based experiments.
 We use a slightly different implementation of FQE from the original implementation, where the $\arg\min$ operation is approximated with mini-batch gradient descent, similar to standard Q-learning as shown in Algorithm 2.
-Algorithm 2 FQE: Fitted Q Evaluation [20]
+Algorithm 2 FQE: Fitted Q Evaluation [Le et al., 2019]
 0:  Offline dataset $\mathcal{D}_{\text{env}}$, policy $\pi_{\theta}$
@@
 3:     $\{\mathbf{s}_{i},\mathbf{a}_{i},r_{i},\mathbf{s}^{\prime}_{i}\}_{i=1}^{N}\sim\mathcal{D}_{\text{env}}$
-4:     $y_{i}=\texttt{sg}(Q_{\phi}(\mathbf{s}^{\prime}_{i},\pi_{\theta}(\mathbf{s}^{\prime}_{i})))$triangleright $\texttt{sg}(\cdot)$ is stop-gradient operator
+4:     $y_{i}=\texttt{sg}(r_{i}+Q_{\phi}(\mathbf{s}^{\prime}_{i},\pi_{\theta}(\mathbf{s}^{\prime}_{i})))$triangleright $\texttt{sg}(\cdot)$ is stop-gradient operator
 5:     $L_{\text{FQE}}(\phi)=\frac{1}{N}\sum_{i=1}^{N}(Q_{\phi}(\mathbf{s}_{i},\mathbf{a}_{i})-y_{i})^{2}$
@@
 We provide a comparison of LEQ with the prior model-based approaches and the baseline methods used in our ablation studies in Table 8.
-Table 6: Comparisons with the prior model-based approaches and the baseline methods introduced for our ablation studies.
-We color the hyperparameters in blue if they are the same with LEQ.
-Otherwise, we color them in red.
-Components CBOP MOBILE MOBILE MOBIP LEQ (ours)
-Training scheme MVE [5] MBPO [13] MBPO [13] Dyna [32] Dyna [32]
-Conservatism Lower-confidence bound Lower-confidence bound Lower-confidence bound Lower-confidence bound Lower expectile
-Policy Stochastic Stochastic Stochastic Deterministic Deterministic
-Policy objective $Q(\mathbf{s},\mathbf{a})$ $Q(\mathbf{s},\mathbf{a})$ $Q(\mathbf{s},\mathbf{a})$ $\lambda$-returns $\lambda$-returns
-Policy pretraining BC BC BC
-# of critics 20-50 2 1 1 1
-Critic objective Multi-step (adaptive weighting) One-step One-step $\lambda$-returns $\lambda$-returns
-Critic pretraining FQE [20] FQE [20] FQE [20]
-Horizon length ($H$) 10 1 1 10 10
-Rollout length ($R$) 1 or 5 10 5 5
-Discount rate ($\gamma$) 0.99 0.99 0.997 0.997 0.997
-$\beta$ in Equation 7 1.0 0.95 0.25 0.25 0.25
-Impl. tricks Clip Q-values with 00 LayerNorm + Symlog LayerNorm + Symlog LayerNorm + Symlog
-Running time 24h 12h 40m 4h 2h
+Table 8: Comparisons with the prior model-based methods and the baseline method. The hyperparameters same with LEQ are colored in blue; others are colored in red.
+Components CBOP MOBILE MOBILE LEQ (ours)
+Training scheme MVE [Feinberg et al., 2018] MBPO [Janner et al., 2019] MBPO [Janner et al., 2019] Dreamer [Hafner et al., 2023]
+Conservatism Lower-confidence bound Lower-confidence bound Lower-confidence bound Lower expectile
+Policy Stochastic Stochastic Stochastic Deterministic
+Policy objective $Q(\mathbf{s},\mathbf{a})$ $Q(\mathbf{s},\mathbf{a})$ $Q(\mathbf{s},\mathbf{a})$ $\lambda$-returns
+Policy pretraining BC BC
+# of critics 20-50 2 1 1
+Critic objective Multi-step (adaptive weighting) One-step One-step $\lambda$-returns
+Critic pretraining FQE [Le et al., 2019] FQE [Le et al., 2019]
+Horizon length ($H$) 10 1 1 10
+Expansion length ($R$) 1 or 5 10 5
+Discount rate ($\gamma$) 0.99 0.99 0.997 0.997
+$\beta$ in Equation 7 1.0 0.95 0.25 0.25
+Impl. tricks Clip Q-values with 00 LayerNorm + Symlog LayerNorm + Symlog
+Running time 24h 12h 40m 2h
 Hyperparameters of LEQ.
-We report task-agnostic hyperparameters of our method in Table 7.
-We note that we use the same hyperparameters across all tasks, except $\tau$.
+For state-based experiments, we report task-agnostic hyperparameters of our method in Table 9.
+We note that we use the same hyperparameters across all state-based tasks, except $\tau$.
 We search the value of $\tau$ in $\{0.1,0.3,0.4,0.5\}$ and report the best value for the main experimental results.
-In addition, we report the exhaustive results in Tables 11 and 12, and summarize $\tau$ used in the main results in Table 8.
-Table 7: Shared hyperparameters of LEQ.
+In addition, we report the exhaustive results in Tables 15 and 16, and summarize $\tau$ used in the main results in Table 10.
+For pixel-based experiments, we decrease the horizon length of DreamerV3 from $15$ to $5$, since we do not observe performance improvement with the longer imagination horizon (Appendix D), consistent with the finding from Lu et al.
+[2023].
+Moreover, we remove the entropy bonus, as exploration is not required in offline RL.
+We search the value of $\tau$ in $\{0.1,0.3,0.4,0.5\}$ and report the best value in the main experimental results.
+All other hyperparameters follow the default settings of DreamerV3.
+We also report the exhaustive results in Table 17, and summarize $\tau$ in Table 13.
+Table 9: Shared hyperparameters of LEQ in state-based experiments.
 Hyperparameters Value Description
@@
 $H$ 10 Imagination length
-$\beta_{\text{EMA}}$ 1 Weight for critic EMA regularization
+$\omega_{\text{EMA}}$ 1 Weight for critic EMA regularization
 $\epsilon_{\text{EMA}}$ 0.995 Critic EMA decay
-Table 8: Task-specific hyperparameter $\tau$ of LEQ.
+Table 10: Task-specific hyperparameter $\tau$ of LEQ in state-based experiments.
 Domain Task $\tau$
@@
 Task-specific hyperparameters of the compared methods.
-We report the best hyperparameters of MOBILE for the AntMaze tasks in Tables 10 and 10.
-For MOBILE and MOBILE, we search the value of $c$ within $\{0.1,0.5,1.0,1.5\}$, as suggested in MOBILE [30], where $c$ is the coefficient of the penalized bellman operator:
+We report the best hyperparameters of MOBILE for the AntMaze tasks in Tables 12 and 12.
+For MOBILE and MOBILE, we search the value of $c$ within $\{0.1,0.5,1.0,1.5\}$, as suggested in MOBILE [Sun et al., 2023], where $c$ is the coefficient of the penalized bellman operator:
 $T\hat{Q}(\mathbf{s},\mathbf{a})=r(\mathbf{s},\mathbf{a})+\gamma Q(\mathbf{s}^{\prime},\mathbf{a}^{\prime})-c\cdot\text{Std}(Q(\mathbf{s}^{\prime},\mathbf{a}^{\prime})).$ (12)
@@
 We do not report the best hyperparameter for MOBILE and CBOP because both methods score zero points for all hyperparameters in AntMaze.
+For MOPO in V-D4RL experiments, we search for $\lambda$ in $\{3,10\}$, as suggested in Lu et al.
+[2023], where $\lambda$ is the penalization coefficient in MOPO.
+Then, we report the best value in Table 14.
+For ROSMO, we use the hyperparameter specified in the official code.
 Table 11: Task-specific hyperparameters in MOBILE.
@@
 ultra-diverse $0.5$
+Table 13: Task-specific hyperparameter $\tau$ of LEQ in V-D4RL experiments.
+Domain Task $\tau$
+walker_walk random $0.5$
+medium $0.3$
+medium_replay $0.3$
+medium_expert $0.1$
+cheetah_run random $0.3$
+medium $0.1$
+medium_replay $0.1$
+medium_expert $0.5$
+Table 14: Task-specific hyperparameter $\lambda$ of MOPO in V-D4RL experiments.
+Domain Task $\lambda$
+walker_walk random $3.0$
+medium $3.0$
+medium_replay $3.0$
+medium_expert $3.0$
+cheetah_run random $3.0$
+medium $3.0$
+medium_replay $10.0$
+medium_expert $10.0$
 Appendix B Proof of the Policy Objective
@@
 In other words, we show that optimizing the following policy objective:
-$\hat{J}_{\lambda}(\theta)=\mathbb{E}_{\tau\sim p_{\psi},\pi_{\theta}}[(W^{\tau}(Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})>Q^{\lambda}_{t}(\tau))Q_{t}^{\lambda}(\tau)],$ (14)
-leads to optimizing a lower-bias estimator of $\mathbb{E}^{\tau}_{\tau\sim p_{\psi},\pi_{\theta}}[Q_{t}^{\lambda}(\tau)]$ than $Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})$.
-To show this, we first prove that $\hat{Y}_{\text{new}}=\frac{\mathbb{E}[W^{\tau}(Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})>Q_{t}^{\lambda}(\tau))\cdot Q_{t}^{\lambda}(\tau)]}{\mathbb{E}[W^{\tau}(Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})>Q_{t}^{\lambda}(\tau))]}$ is closer to $\mathbb{E}^{\tau}_{\tau\sim p_{\psi},\pi_{\theta}}[Q_{t}^{\lambda}(\tau)]$ than $Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})$.
-For deriving the proof, we generalize this situation to have an arbitrary distribution $X$ and estimation $\hat{Y}$, which corresponds to $X=Q_{t}^{\lambda}(\tau),\hat{Y}=Q_{\phi}(\mathbf{s},\mathbf{a})$.
+$\hat{J}_{\lambda}(\theta)=\mathbb{E}_{\mathcal{T}\sim p_{\psi},\pi_{\theta}}[(W^{\tau}(Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})>Q^{\lambda}_{t}(\mathcal{T}))Q_{t}^{\lambda}(\mathcal{T})],$ (13)
+leads to optimizing a lower-bias estimator of $\mathbb{E}^{\tau}_{\mathcal{T}\sim p_{\psi},\pi_{\theta}}[Q_{t}^{\lambda}(\mathcal{T})]$ than $Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})$.
+To show this, we first prove that $\hat{Y}_{\text{new}}=\frac{\mathbb{E}[W^{\tau}(Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})>Q_{t}^{\lambda}(\mathcal{T}))\cdot Q_{t}^{\lambda}(\mathcal{T})]}{\mathbb{E}[W^{\tau}(Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})>Q_{t}^{\lambda}(\mathcal{T}))]}$ is closer to $\mathbb{E}^{\tau}_{\mathcal{T}\sim p_{\psi},\pi_{\theta}}[Q_{t}^{\lambda}(\mathcal{T})]$ than $Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})$ for most of the situations.
+For deriving the proof, we generalize the problem by considering an arbitrary distribution $X$ and its estimate $\hat{Y}$, which correspond to $X=Q_{t}^{\lambda}(\tau),\hat{Y}=Q_{\phi}(\mathbf{s},\mathbf{a})$.
+Then, we show $\hat{Y}_{\text{new}}=\frac{\mathbb{E}[W^{\tau}(\hat{Y}>X)\cdot X]}{\mathbb{E}[W^{\tau}(\hat{Y}>X)]}$ is closer to $Y$ than $\hat{Y}$.
+We split the case to two cases where $\hat{Y}\geq Y$ (Lemma 1) and $\hat{Y}\leq Y$ (Lemma 2) and show the effectiveness of $\hat{Y}_{\text{new}}$ instead of $\hat{Y}$ for each case.
+Lemma 1.
+Let $X$ be a distribution and $Y=E^{\tau}[X]$ be a lower expectile of $X$ (i.e.
+$0<\tau\leq 0.5$).
+Let $\hat{Y}$ be an arbitrary optimistic estimate of $Y$ (i.e., $\hat{Y}\geq Y$), and define $W^{\tau}(\cdot)=|\tau-\mathbbm{1}(\cdot)|$.
+If we let $\hat{Y}_{\text{new}}=\frac{\mathbb{E}[W^{\tau}(\hat{Y}>X)\cdot X]}{\mathbb{E}[W^{\tau}(\hat{Y}>X)]}$ be a new estimate of $Y$, then $|\hat{Y}_{\text{new}}-Y|\leq|\hat{Y}-Y|$.
+Proof.
+$|\hat{Y}_{\text{new}}-Y|$
+$=\left\lvert\frac{\mathbb{E}[W^{\tau}(\hat{Y}>X)\cdot X]}{\mathbb{E}[W^{\tau}(\hat{Y}>X)]}-\frac{\mathbb{E}[W^{\tau}(Y>X)\cdot X]}{\mathbb{E}[W^{\tau}(Y>X)]}\right\rvert\quad\text{\qquad\qquad\qquad\qquad\qquad\qquad\;\;\;($\because$ Def.
+of $\hat{Y}_{\text{new}}$ and $Y$)}$
+$=\left\lvert\frac{\mathbb{E}[W^{\tau}(Y>X)\cdot X]+\mathbb{E}[(1-2\tau)\cdot\mathbbm{1}(Y\leq X\leq\hat{Y})\cdot X]}{\mathbb{E}[W^{\tau}(Y>X)]+\mathbb{E}[(1-2\tau)\cdot\mathbbm{1}(Y\leq X\leq\hat{Y})]}-\frac{\mathbb{E}[W^{\tau}(Y>X)\cdot X]}{\mathbb{E}[W^{\tau}(Y>X)]}\right\rvert\quad\text{($\because$ Def.
+of $W^{\tau}(\cdot)$)}$
+$=\left\lvert\frac{\mathbb{E}[W^{\tau}(Y>X)]\mathbb{E}[(1-2\tau)\cdot\mathbbm{1}(Y\leq X\leq\hat{Y})\cdot X]-\mathbb{E}[W^{\tau}(Y>X)\cdot X]\mathbb{E}[(1-2\tau)\cdot\mathbbm{1}(Y\leq X\leq\hat{Y})]}{\mathbb{E}[W^{\tau}(Y>X)](\mathbb{E}[W^{\tau}(Y>X)]+\mathbb{E}[(1-2\tau)\cdot\mathbbm{1}(Y\leq X\leq\hat{Y})])}\right\rvert$
+$=(1-2\tau)\cdot\left\lvert\frac{\mathbb{E}[W^{\tau}(Y>X)]\mathbb{E}[\mathbbm{1}(Y\leq X\leq\hat{Y})\cdot X]-\mathbb{E}[W^{\tau}(Y>X)\cdot X]\mathbb{E}[\mathbbm{1}(Y\leq X\leq\hat{Y})]}{\mathbb{E}[W^{\tau}(Y>X)](\mathbb{E}[W^{\tau}(Y>X)]+\mathbb{E}[(1-2\tau)\cdot\mathbbm{1}(Y\leq X\leq\hat{Y})])}\right\rvert$
+$=(1-2\tau)\cdot\left\lvert\frac{\mathbb{E}[\mathbbm{1}(Y\leq X\leq\hat{Y})\cdot X]-Yp(Y\leq X\leq\hat{Y})}{\mathbb{E}[W^{\tau}(Y>X)]+\mathbb{E}[(1-2\tau)\cdot\mathbbm{1}(Y\leq X\leq\hat{Y})]}\right\rvert$
+$=(1-2\tau)\cdot\left\lvert\frac{p(Y\leq X\leq\hat{Y})(\mathbb{E}_{Y\leq X\leq\hat{Y}}[X]-Y)}{\mathbb{E}[W^{\tau}(Y>X)]+\mathbb{E}[(1-2\tau)\cdot\mathbbm{1}(Y\leq X\leq\hat{Y})]}\right\rvert$
+$=(1-2\tau)\cdot\left\lvert\frac{p(Y\leq X\leq\hat{Y})(\mathbb{E}_{Y\leq X\leq\hat{Y}}[X]-Y)}{(1-\tau)-(1-2\tau)\cdot p(Y\leq X)+(1-2\tau)\cdot p(Y\leq X\leq\hat{Y})}\right\rvert$
+$=(1-2\tau)\cdot\left\lvert\frac{p(Y\leq X\leq\hat{Y})(\mathbb{E}_{Y\leq X\leq\hat{Y}}[X]-Y)}{(1-\tau)-(1-2\tau)\cdot p(\hat{Y}\leq X)}\right\rvert$
+$=(1-2\tau)\cdot\left\lvert\frac{p(Y\leq X\leq\hat{Y})(\mathbb{E}_{Y\leq X\leq\hat{Y}}[X]-Y)}{\tau+(1-2\tau)\cdot p(X\leq\hat{Y})}\right\rvert$
+$\leq\frac{p(Y\leq X\leq\hat{Y})(\mathbb{E}_{Y\leq X\leq\hat{Y}}[X]-Y)}{p(X\leq\hat{Y})}$
+$\leq\mathbb{E}_{Y\leq X\leq\hat{Y}}[X]-Y$
+$\leq\hat{Y}-Y=|\hat{Y}-Y|$
+Lemma 1 shows that when the estimated value is optimistic ($\hat{Y}\geq Y$), the bias of the new estimate is always smaller than that of the original estimate.
+In the context of LEQ algorithm, the lemma tells that if the critic network ($\hat{Y}$) overestimates the lower-expectile of the target returns ($Y=\mathbb{E}^{\tau}[X]$), the surrogate loss ($Y_{\text{new}})$ compensates the overestimation of the critic values.
+Unfortunately, when the estimated value is pessimistic ($\hat{Y}\leq Y$), there are some exceptional cases that the surrogate loss overcompensates the underestimation, resulting in an even larger error.
+The boundary for these cases is characterized in Lemma 2:
+Lemma 2.
+Let $X$ be a distribution and $Y=E^{\tau}[X]$ be a lower expectile of $X$ (i.e.
+$0<\tau\leq 0.5$).
+Let $\hat{Y}$ be an arbitrary conservative estimate of $Y$ (i.e., $\hat{Y}\leq Y$), and define $W^{\tau}(\cdot)=|\tau-\mathbbm{1}(\cdot)|$.
+If we let $\hat{Y}_{\text{new}}=\frac{\mathbb{E}[W^{\tau}(\hat{Y}>X)\cdot X]}{\mathbb{E}[W^{\tau}(\hat{Y}>X)]}$ be a new estimate of $Y$, then $|\hat{Y}_{\text{new}}-Y|\leq|\hat{Y}-Y|$, when $p(X\leq\hat{Y})\geq\frac{1}{2}(p(X\leq Y)-\frac{\tau}{1-2\tau})$.
+Proof.
+$|\hat{Y}_{\text{new}}-Y|$
+$=\left\lvert\frac{\mathbb{E}[W^{\tau}(\hat{Y}>X)\cdot X]}{\mathbb{E}[W^{\tau}(\hat{Y}>X)]}-\frac{\mathbb{E}[W^{\tau}(Y>X)\cdot X]}{\mathbb{E}[W^{\tau}(Y>X)]}\right\rvert\quad\text{\qquad\qquad\qquad\qquad\qquad\qquad\;\;($\because$ Def.
+of $\hat{Y}_{\text{new}}$ and $Y$)}$
+$=\left\lvert\frac{\mathbb{E}[W^{\tau}(Y>X)\cdot X]-\mathbb{E}[(1-2\tau)\cdot\mathbbm{1}(\hat{Y}\leq X\leq Y)X]}{\mathbb{E}[W^{\tau}(Y>X)]-\mathbb{E}[(1-2\tau)\cdot\mathbbm{1}(\hat{Y}\leq X\leq Y)]}-\frac{\mathbb{E}[W^{\tau}(Y>X)\cdot X]}{\mathbb{E}[W^{\tau}(Y>X)]}\quad\text{($\because$ Def.
+of$W^{\tau}(\cdot)$)}\right\rvert$
+$=\left\lvert\frac{\mathbb{E}[W^{\tau}(Y>X)]\mathbb{E}[(1-2\tau)\cdot\mathbbm{1}(\hat{Y}\leq X\leq Y)\cdot X]-\mathbb{E}[W^{\tau}(Y>X)\cdot X]\mathbb{E}[(1-2\tau)\cdot\mathbbm{1}(\hat{Y}\leq X\leq Y)]}{\mathbb{E}[W^{\tau}(Y>X)](\mathbb{E}[W^{\tau}(Y>X)]-\mathbb{E}[(1-2\tau)\cdot\mathbbm{1}(\hat{Y}\leq X\leq Y)])}\right\rvert$
+$=(1-2\tau)\cdot\left\lvert\frac{\mathbb{E}[W^{\tau}(Y>X)X]\mathbb{E}[\mathbbm{1}(\hat{Y}\leq X\leq Y)]-\mathbb{E}[W^{\tau}(Y>X)]\mathbb{E}[\mathbbm{1}(\hat{Y}\leq X\leq Y)\cdot X]}{\mathbb{E}[W^{\tau}(Y>X)](\mathbb{E}[W^{\tau}(Y>X)]-\mathbb{E}[(1-2\tau)\cdot\mathbbm{1}(\hat{Y}\leq X\leq Y)])}\right\rvert$
+$=(1-2\tau)\cdot\left\lvert\frac{Yp(\hat{Y}\leq X\leq Y)-\mathbb{E}[\mathbbm{1}(\hat{Y}\leq X\leq Y)\cdot X]}{\mathbb{E}[W^{\tau}(Y>X)]-\mathbb{E}[(1-2\tau)\cdot\mathbbm{1}(\hat{Y}\leq X\leq Y)]}\right\rvert$
+$=(1-2\tau)\cdot\left\lvert\frac{p(\hat{Y}\leq X\leq Y)(Y-\mathbb{E}_{\hat{Y}\leq X\leq Y}[X])}{\mathbb{E}[W^{\tau}(Y>X)]-\mathbb{E}[(1-2\tau)\cdot\mathbbm{1}(\hat{Y}\leq X\leq Y)]}\right\rvert$
+$=(1-2\tau)\cdot\left\lvert\frac{p(\hat{Y}\leq X\leq Y)(Y-\mathbb{E}_{\hat{Y}\leq X\leq Y}[X])}{(1-\tau)-(1-2\tau)\cdot p(Y\leq X)-(1-2\tau)\cdot p(\hat{Y}\leq X\leq Y)}\right\rvert$
+$=(1-2\tau)\cdot\left\lvert\frac{p(\hat{Y}\leq X\leq Y)(Y-\mathbb{E}_{\hat{Y}\leq X\leq Y}[X])}{(1-\tau)-(1-2\tau)\cdot p(\hat{Y}\leq X)}\right\rvert$
+$=(1-2\tau)\cdot\left\lvert\frac{p(\hat{Y}\leq X\leq Y)(Y-\mathbb{E}_{\hat{Y}\leq X\leq Y}[X])}{\tau+(1-2\tau)\cdot p(X\leq\hat{Y})}\right\rvert$
+$=\left\lvert\frac{p(\hat{Y}\leq X\leq Y)(Y-\mathbb{E}_{\hat{Y}\leq X\leq Y}[X])}{\frac{\tau}{1-2\tau}+p(X\leq\hat{Y})}\right\rvert$
+$=\frac{p(X\leq Y)-p(X\leq\hat{Y})}{\frac{\tau}{1-2\tau}+p(X\leq\hat{Y})}\cdot(Y-\mathbb{E}_{\hat{Y}\leq X\leq Y}[X])$
+$\leq Y-\mathbb{E}_{\hat{Y}\leq X\leq Y}[X]$
+$\leq Y-\hat{Y}=|\hat{Y}-Y|$
+By combining Lemma 1, Lemma 2, we get Theorem 1 as below:
 Theorem 1.
@@
 $0<\tau\leq 0.5$).
-Let $\hat{Y}$ be an arbitrary estimation of $Y$, and define $W^{\tau}(\cdot)=|\tau-\mathbbm{1}(\cdot)|$.
-If we let $\hat{Y}_{\text{new}}=\frac{\mathbb{E}[W^{\tau}(\hat{Y}>X)\cdot X]}{\mathbb{E}[W^{\tau}(\hat{Y}>X)]}$ be a new estimation of $Y$, then $|Y-\hat{Y}_{\text{new}}|\leq\frac{1-2\tau}{1-\tau}\cdot p(Y\leq X\leq\hat{Y})\cdot|Y-\hat{Y}|$.
+Let $\hat{Y}$ be an arbitrary estimate of $Y$, and define $W^{\tau}(\cdot)=|\tau-\mathbbm{1}(\cdot)|$.
+If we let $\hat{Y}_{\text{new}}=\frac{\mathbb{E}[W^{\tau}(\hat{Y}>X)\cdot X]}{\mathbb{E}[W^{\tau}(\hat{Y}>X)]}$ be a new estimate of $Y$, then $|\hat{Y}_{\text{new}}-Y|\leq|\hat{Y}-Y|$ if the following condition holds:
+$p(X\leq\hat{Y})\geq\frac{1}{2}(p(X\leq Y)-\frac{\tau}{1-2\tau}).$ (14)
 Proof.
-Without loss of generality, we assume $\hat{Y}\geq Y$. Then, we have $\hat{Y}_{\text{new}}\geq Y$. Thus,
-$|\hat{Y}_{\text{new}}-Y|$
-$=\hat{Y}_{\text{new}}-Y$
-$=\frac{\mathbb{E}[W^{\tau}(\hat{Y}>X)\cdot X]}{\mathbb{E}[W^{\tau}(\hat{Y}>X)]}-\frac{\mathbb{E}[W^{\tau}(Y>X)\cdot X]}{\mathbb{E}[W^{\tau}(Y>X)]}\quad\text{\qquad\qquad\qquad\qquad\qquad\qquad\;\;\;($\because$ Def.
-of $\hat{Y}_{\text{new}}$ and $Y$)}$
-$=\frac{\mathbb{E}[W^{\tau}(Y>X)\cdot X]+\mathbb{E}[(1-2\tau)\cdot\mathbbm{1}(Y\leq X\leq\hat{Y})X]}{\mathbb{E}[W^{\tau}(Y>X)]+\mathbb{E}[(1-2\tau)\cdot\mathbbm{1}(Y\leq X\leq\hat{Y})]}-\frac{\mathbb{E}[W^{\tau}(Y>X)\cdot X]}{\mathbb{E}[W^{\tau}(Y>X)]}\quad\text{($\because$ Def.
-of $W^{\tau}(\cdot)$)}$
-$=\frac{\mathbb{E}[W^{\tau}(Y>X)]\mathbb{E}[(1-2\tau)\cdot\mathbbm{1}(Y\leq X\leq\hat{Y})X]-\mathbb{E}[W^{\tau}(Y>X)X]\mathbb{E}[(1-2\tau)\cdot\mathbbm{1}(Y\leq X\leq\hat{Y})]}{\mathbb{E}[W^{\tau}(Y>X)](\mathbb{E}[W^{\tau}(Y>X)]+\mathbb{E}[(1-2\tau)\cdot\mathbbm{1}(Y\leq X\leq\hat{Y})])}$
-$\leq\frac{1-2\tau}{\mathbb{E}[W^{\tau}(Y>X)]^{2}}\cdot(\mathbb{E}[W^{\tau}(Y>X)]\mathbb{E}[\mathbbm{1}(Y\leq X\leq\hat{Y})X]-\mathbb{E}[W^{\tau}(Y>X)X]\mathbb{E}[\mathbbm{1}(Y\leq X\leq\hat{Y})]$
-$=\frac{1-2\tau}{\mathbb{E}[W^{\tau}(Y>X)]}\cdot(\mathbb{E}[\mathbbm{1}(Y\leq X\leq\hat{Y})X]-Yp(Y\leq X\leq\hat{Y}))$
-$=\frac{(1-2\tau)p(Y\leq X\leq\hat{Y})}{\mathbb{E}[W^{\tau}(Y>X)]}\cdot(\mathbb{E}_{Y\leq X\leq\hat{Y}}[X]-Y)$
-$\leq\frac{(1-2\tau)p(Y\leq X\leq\hat{Y})}{\mathbb{E}[W^{\tau}(Y>X)]}\cdot(\hat{Y}-Y)$
-$\leq\frac{1-2\tau}{1-\tau}\cdot p(Y\leq X\leq\hat{Y})\cdot|\hat{Y}-Y|$
-Note that this theorem shows that the bias of the new estimation is always smaller than the original estimation, since $\frac{1-2\tau}{1-\tau}<1$ and $p(Y\leq X\leq\hat{Y})\leq 1$.
-If we plug in the distribution of $Q_{t}^{\lambda}(\tau)$ to $X$ and $\hat{Y}=Q_{\phi}(s_{t},\mathbf{a}_{t})$, then $Y=\mathbb{E}^{\tau}[X]=\mathbb{E}^{\tau}_{\tau\sim p_{\psi},\pi_{\theta}}[Q_{t}^{\lambda}(\tau)]$, and we can show the desired result using the theorem: $\hat{Y}_{\text{new}}=\frac{\mathbb{E}[W^{\tau}(Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})>Q_{t}^{\lambda}(\tau))\cdot Q_{t}^{\lambda}(\tau)]}{\mathbb{E}[W^{\tau}(Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})>Q_{t}^{\lambda}(\tau)]}$ is closer to $\mathbb{E}^{\tau}_{\tau\sim p_{\psi},\pi_{\theta}}[Q_{t}^{\lambda}(\tau)]$ than $Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})$.
-Here, the normalizing factor $\mathbb{E}[W^{\tau}(Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})>Q_{t}^{\lambda}(\tau))]$ is non-differentiable with $\tau$.
-Specifically, the gradient is 0 everywhere (except $Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})=Q_{t}^{\lambda}(\tau)$).
+We combine the two cases dealt in Lemma 1, Lemma 2.
+Here, $\hat{Y}\geq Y$ from Lemma 1 can be omitted, since the condition of Equation 14 includes the case of $\hat{Y}\geq Y$.
+[…]
+The condition of Equation 14 is a conservative bound applicable to any distribution, and $|\hat{Y}_{\text{new}}-Y|\leq|\hat{Y}-Y|$ holds across much broader regions in general.
+For example, Figure 8 shows that the inequality holds for 100%, 99.9% of the cases for normal, uniform distributions, respectively.
+(a) Normal distribution
+(b) Uniform distribution
+Figure 8: Region of exceptions for the surrogate loss in realistic distributions.
+We visualize the cases where the surrogate loss is worse than the direct policy optimization, for two common distributions for $X$: (a) a normal distribution $\mathcal{N}(0,1)$ and (b) a uniform distribution $\mathcal{U}(0,1)$.
+We sample $\tau$ and $\hat{\tau}$ from arange($0,0.5,0.0025$) and compute $Y=\mathbb{E}^{\tau}[X]$, $\hat{Y}=\mathbb{E}^{\hat{\tau}}[X]$.
+The plots show the difference of errors $|\hat{Y}_{\text{new}}-Y|-|\hat{Y}-Y|$: positive values (red to yellow) indicate that the surrogate loss performs worse than the estimate, while negative values (green to blue) indicate that the surrogate loss outperforms the estimate.
+There is no exception for normal distribution, and only 39 out of 40000 cases are exceptional for uniform distribution.
+Here, we illustrate how we optimize $\mathbb{E}[W^{\tau}(Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})>Q_{t}^{\lambda}(\mathcal{T}))\cdot Q_{t}^{\lambda}(\mathcal{T})]$ instead of $\frac{\mathbb{E}[W^{\tau}(Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})>Q_{t}^{\lambda}(\mathcal{T}))\cdot Q_{t}^{\lambda}(\mathcal{T})]}{\mathbb{E}[W^{\tau}(Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})>Q_{t}^{\lambda}(\mathcal{T})]}$.
+The normalizing factor $\mathbb{E}[W^{\tau}(Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})>Q_{t}^{\lambda}(\mathcal{T}))]$ is non-differentiable with $\mathcal{T}$ and the gradient is 0 everywhere (except $Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})=Q_{t}^{\lambda}(\mathcal{T})$).
 Thus, if we calculate the gradient of $\hat{Y}_{\text{new}}$, the gradient for the normalizing factor disappears.
-Therefore, we can omit the normalizing factor and get an equivalent formula $\mathbb{E}[W^{\tau}(Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})>Q_{t}^{\lambda}(\tau))\cdot Q_{t}^{\lambda}(\tau)]$ for gradient-based optimization.
-Appendix C More Results
-High variance in locomotion tasks.
-When we train LEQ in locomotion tasks, we observe that our method often achieves $100$% success rates and then falls back to 00%, as shown in Figure 5.
-This is mainly because the learned models sometimes fail to capture failures (e.g.
-hopper and walker falling off) and predict an optimistic future (e.g.
-hopper and walker walking forward).
-Figure 5: High variance during training. Our algorithm experiences oscillation on the performances due to optimistic imaginations near the initial states.
-Results for all expectiles $\tau$.
+Therefore, we omit the normalizing factor and get an equivalent formula $\mathbb{E}[W^{\tau}(Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})>Q_{t}^{\lambda}(\mathcal{T}))\cdot Q_{t}^{\lambda}(\mathcal{T})]$ for gradient-based optimization.
+Appendix C Results for All Expectiles $\tau$
 To give insights how the expectile parameter $\tau$ affects the performance of LEQ, we report the performance of LEQ with all expectile values $\{0.1,0.3,0.4,0.5\}$.
@@
 antmaze-large-play $42.0$$\pm 30.6$ $\mathbf{58.6}$$\pm 9.1$ $52.2$$\pm 15.8$ $42.2$$\pm 7.3$
-antmaze-large-diverse $60.6$$\pm 32.1$ $\mathbf{60.2}$$\pm 18.3$ $48.8$$\pm 5.8$ $36.8$$\pm 9.7$
+antmaze-large-diverse $\mathbf{60.6}$$\pm 32.1$ $60.2$$\pm 18.3$ $48.8$$\pm 5.8$ $36.8$$\pm 9.7$
 antmaze-ultra-play $\mathbf{25.8}$$\pm 18.2$ $10.8$±$\pm 8.8$ $11.6$$\pm 12.5$ $9.2$$\pm 11.5$
@@
 halfcheetah-me $\mathbf{102.8}$$\pm 0.4$ $81.5$$\pm 19.6$ $58.1$$\pm 26.1$ $46.3$$\pm 17.7$
+Table 17: V-D4RL results of LEQ with different expectiles. We report the results in Antmaze task with expectiles value of 0.1, 0.3, 0.4, 0.5. The best value is highlighted.
+Expectile 0.1 0.3 0.4 0.5
+walker_walk-random $14.5$$\pm 1.1$ $20.2$$\pm 3.6$ $\mathbf{22.4}$$\pm 1.1$ $21.8$$\pm 0.8$
+walker_walk-medium $\mathbf{43.1}$$\pm 3.2$ $37.2$$\pm 5.3$ $35.0$$\pm 3.7$ $26.6$$\pm 1.9$
+walker_walk-medium_replay $40.6$$\pm 6.6$ $40.8$$\pm 3.4$ $41.3$$\pm 6.3$ $\mathbf{43.0}$$\pm 7.3$
+walker_walk-medium_expert $\mathbf{87.2}$$\pm 2.4$ $82.7$$\pm 5.0$ $77.0$$\pm 9.2$ $85.0$$\pm 1.6$
+cheetah_run-random $12.1$$\pm 1.9$ $\mathbf{14.8}$$\pm 1.0$ $14.2$$\pm 1.6$ $14.6$$\pm 2.9$
+cheetah_run-medium $25.0$$\pm 5.9$ $\mathbf{37.9}$$\pm 8.0$ $23.6$$\pm 11.9$ $32.2$$\pm 6.2$
+cheetah_run-medium_replay $\mathbf{34.3}$$\pm 1.1$ $32.3$$\pm 2.1$ $31.2$$\pm 3.5$ $31.2$$\pm 0.4$
+cheetah_run-medium_expert $23.9$$\pm 4.7$ $18.9$$\pm 3.9$ $\mathbf{25.1}$$\pm 6.6$ $24.3$$\pm 7.6$
+Total $280.7$ $284.8$ $269.9$ $278.6$
+Appendix D More Ablation Results
+Ablation of each component in LEQ.
+In Table 18, we ablate LEQ’s design choices on AntMaze tasks to investigate their importance.
+Notably, the performance of LEQ drops significantly if we remove the Q-learning loss ($461.8$ to $312.8$) or apply expectile regression for real transitions ($461.8$ to $249.2$), showing the importance of utilizing the real transitions in long-horizon tasks.
+Moreover, the expectile loss from the policy learning is crucial ($461.8$ to $357.0$), since simply maximizing $\lambda$-returns, even if $Q(\mathbf{s},\mathbf{a})$ is conservative, can lead to a suboptimal policy due to the lack of conservatism in the rewards from model rollouts.
+Minor design choices, such as EMA regularization and dataset expansion, have a small impact on the performance in AntMaze.
+Table 18: Ablation studies about various components in LEQ on AntMaze.
+We conduct ablation studies on (1) EMA regularization, (2) dataset expansion, (3) utilizing data transitions, (4) lower expectile policy optimization, (5) lower expectile regression for data transitions
+Method umaze medium large ultra Total
+umaze diverse play diverse play diverse play diverse
+LEQ $\mathbf{94.4}$$\pm 6.3$ $71.0$$\pm 12.3$ $50.2$$\pm 39.9$ $\mathbf{46.2}$$\pm 23.2$ $58.6$$\pm 9.1$ $60.2$$\pm 18.3$ $25.8$$\pm 18.2$ $\mathbf{55.8}$$\pm 18.3$ $\mathbf{461.8}$
+- EMA regularization $\mathbf{96.0}$$\pm 4.0$ $65.6$$\pm 5.1$ $33.8$$\pm 31.4$ $21.2$$\pm 26.1$ $58.8$$\pm 8.2$ $62.2$$\pm 15.0$ $8.0$$\pm 13.9$ $33.5$$\pm 35.1$ $379.1$
+- dataset expansion $\mathbf{97.4}$$\pm 1.4$ $63.0$$\pm 23.2$ $\mathbf{58.2}$$\pm 28.0$ $28.6$$\pm 33.7$ $56.0$$\pm 9.8$ $57.0$$\pm 4.5$ $\mathbf{39.2}$$\pm 15.1$ $36.0$$\pm 12.0$ $435.4$
+- data transitions $63.0$$\pm 19.8$ $\mathbf{75.3}$$\pm 4.2$ $31.8$$\pm 34.0$ $28.0$$\pm 27.0$ $50.5$$\pm 20.9$ $57.5$$\pm 8.6$ $0.8$$\pm 1.3$ $6.0$$\pm 5.1$ 312.8
+- expectile in policy update $93.4$$\pm 4.2$ $43.4$$\pm 26.1$ $45.8$$\pm 21.8$ $11.4$$\pm 18.6$ $53.8$$\pm 7.5$ $54.0$$\pm 6.9$ $33.8$$\pm 11.6$ $21.4$$\pm 13.2$ $357.0$
++ expectile in data transitions $53.4$$\pm 18.8$ $4.8$$\pm 7.7$ $27.2$$\pm 37.0$ $19.2$$\pm 11.9$ $\mathbf{69.0}$$\pm 17.4$ $\mathbf{75.4}$$\pm 6.4$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $249.2$
+Ablation on pretraining.
+Table 19 shows the effect of BC and FQE pretraining for LEQ and MOBILE in AntMaze tasks.
+Without pretraining, the performance of LEQ decreases (461.8 $\rightarrow$ 322.2), especially in medium mazes (96.4 $\rightarrow$ 4.6) and ultra mazes (81.6 $\rightarrow$ 12.4) due to early instability of training, matching the observation in CBOP [Jeong et al., 2023].
+However, for MOBILE, pretraining worsens the performance (285.3 $\rightarrow$ 232.2).
+Table 19: Ablation results for pretraining on AntMaze. Results are averaged over 5 random seeds.
+Method umaze medium large ultra Total
+conservatism pretrain umaze diverse play diverse play diverse play diverse
+LEQ O $\mathbf{94.4}$$\pm 6.3$ $71.0$$\pm 12.3$ $50.2$$\pm 39.9$ $\mathbf{46.2}$$\pm 23.2$ $58.6$$\pm 9.1$ $60.2$$\pm 18.3$ $25.8$$\pm 18.2$ $\mathbf{55.8}$$\pm 18.3$ $\mathbf{461.8}$
+LEQ X $\mathbf{94.0}$$\pm 1.9$ $65.6$$\pm 6.8$ $0.8$$\pm 1.6$ $3.8$$\pm 7.6$ $60.4$$\pm 13.4$ $59.8$$\pm 12.5$ $11.6$$\pm 12.4$ $26.2$$\pm 22.4$ $322.2$
+MOBILE O $73.4$$\pm 12.6$ $46.0$$\pm 9.4$ $24.6$$\pm 23.2$ $11.6$$\pm 9.9$ $31.0$$\pm 8.4$ $33.2$$\pm 10.6$ $12.4$$\pm 6.9$ $0.0$ ±$\pm 0.0$ $232.2$
+MOBILE X $53.8$$\pm 26.8$ $22.5$$\pm 22.2$ $54.0$$\pm 5.8$ $\mathbf{49.5}$$\pm 6.2$ $28.3$$\pm 6.0$ $28.0$$\pm 11.4$ $\mathbf{25.5}$$\pm 6.9$ $23.8$$\pm 15.8$ $\mathbf{285.3}$
 Ablation study on dataset expansion.
@@
 Total $\mathbf{923.2}$ $788.2$
-Ablation study on MOBILE in AntMaze.
-In Table 14, we report the performance of MOBILE for all possible combination of the hyperparameters between the values of MOBILE and LEQ.
-Specifically, we use $\beta=\{0.25,0.95\}$, $\gamma=\{0.99,0.997\}$, and $R=\{5,10\}$.
-The result shows that $\beta=0.25$ is crucial.
-In addition, the configuration of LEQ yields the best result among all configurations for MOBILE.
-Table 14: Complete hyperparameter search results of MOBILE on AntMaze.
-MOBILE uses the hyperparameters from MOBILE: $\beta=0.95$, $\gamma=0.99$, and $R=5$, whereas LEQ uses $\beta=0.25$, $\gamma=0.997$, and $R=10$.
-The results show that $\beta$ is the most critical hyperparameter that makes MOBILE work in AntMaze.
-  Hyperparams. umaze medium large ultra Total
-$\beta$ $\gamma$ $R$ umaze diverse play diverse play diverse play diverse
-$0.25$ $0.997$ $10$ $53.8$$\pm 26.8$ $\mathbf{22.5}$$\pm 22.2$ $54.0$$\pm 5.8$ $\mathbf{49.5}$$\pm 6.2$ $\mathbf{28.3}$$\pm 6.0$ $\mathbf{28.0}$$\pm 11.4$ $\mathbf{25.5}$$\pm 6.9$ $\mathbf{23.8}$$\pm 15.8$ $\mathbf{285.3}$
-$0.25$ $0.997$ $5$ $\mathbf{74.0}$$\pm 6.9$ $3.7$$\pm 2.6$ $54.7$$\pm 27.9$ $28.0$$\pm 9.6$ $18.7$$\pm 18.6$ $8.0$$\pm 9.3$ $9.7$$\pm 8.2$ $9.0$$\pm 3.7$ $205.7$
-$0.25$ $0.99$ $10$ $39.7$$\pm 23.4$ $5.0$$\pm 7.1$ $39.3$$\pm 27.9$ $38.0$$\pm 15.0$ $0.0$ ±$\pm 0.0$ $3.7$$\pm 5.2$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $125.7$
-$0.25$ $0.99$ $5$ $\mathbf{77.0}$$\pm 6.4$ $20.4$$\pm 15.7$ $\mathbf{64.6}$$\pm 11.1$ $31.6$$\pm 16.9$ $2.6$$\pm 2.8$ $7.2$$\pm 8.9$ $4.6$$\pm 3.0$ $5.0$$\pm 4.6$ $213.0$
-$0.95$ $0.997$ $10$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $1.8$$\pm 3.0$ $0.5$$\pm 0.9$ $0.2$$\pm 0.4$ $2.2$ ±$\pm 2.3$ $1.0$$\pm 1.7$ $0.0$ ±$\pm 0.0$ $5.7$
-$0.95$ $0.997$ $5$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $7.2$$\pm 4.1$ $1.6$$\pm 2.1$ $9.6$$\pm 7.1$ $5.4$$\pm 4.9$ $0.0$ ±$\pm 0.0$ $1.8$$\pm 2.7$ $25.6$
-$0.95$ $0.99$ $10$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $5.0$$\pm 5.1$ $0.6$$\pm 1.2$ $7.4$$\pm 14.8$ $1.6$$\pm 3.2$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $14.6$
-$0.95$ $0.99$ $5$ $1.0$$\pm 2.0$ $0.0$ ±$\pm 0.0$ $6.4$$\pm 5.5$ $5.0$$\pm 5.0$ $0.8$$\pm 1.6$ $0.8$$\pm 1.2$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $14.0$
+Ablation on using deterministic policy.
+We parameterize the stochastic policy $\pi(\cdot|s)$ as $\pi(a|s)=\tanh(N(\mu(s),\sigma^{2}(s)))$ as  Haarnoja et al.
+[2018] and run LEQ with this configuration.
+However, we found that the policy quickly becomes deterministic, because LEQ inadvertently penalizes the stochasticity of the policy while penalizing the uncertainty of the model rollouts.
+Specifically, when we use a stochastic policy, the stochasticity of the policy contributes to increasing the variance of $\lambda$-returns, which are therefore heavily penalized by LEQ.
+To compensate for this effect, we use an entropy bonus coefficient $\alpha=0.0003$.
+As demonstrated in Table 21, stochastic policy shows slightly worse performance compared to deterministic policy (461.8 $\rightarrow$ 380.4).
+However, we believe that LEQ can be extended to stochastic policies with further hyperparameter tuning on the stochasticity of the policy.
+Table 21: Ablation results on using deterministic policy. Results are averaged over 5 random seeds.
+Policy umaze medium large ultra Total
+umaze diverse play diverse play diverse play diverse
+Deterministic $\mathbf{94.4}$$\pm 6.3$ $\mathbf{71.0}$$\pm 12.3$ $\mathbf{50.2}$$\pm 39.9$ $\mathbf{46.2}$$\pm 23.2$ $\mathbf{58.6}$$\pm 9.1$ $\mathbf{60.2}$$\pm 18.3$ $25.8$$\pm 18.2$ $\mathbf{55.8}$$\pm 18.3$ $\mathbf{461.8}$
+Stochastic $\mathbf{91.0}$$\pm 7.9$ $\mathbf{70.4}$$\pm 6.7$ $35.8$$\pm 29.7$ $0.0$ ±$\pm 0.0$ $41.8$$\pm 16.5$ $50.2$$\pm 7.8$ $\mathbf{43.4}$$\pm 25.0$ $47.8$$\pm 23.5$ $380.4$
+Ablation on using learned terminal function.
+We conduct an ablation study to evaluate the impact of using learned terminal functions instead of the ground-truth terminal function.
+For the terminal prediction network, we use a 3-layer MLP with a hidden size of 256, consistent with the architecture of the policy and critic networks.
+As shown in Table 21, replacing the true terminal function with a learned terminal function leads to a significant drop in performance.
+Table 22: Ablation results on using learned terminal function. Results are averaged over 5 random seeds.
+Terminal umaze medium large ultra Total
+umaze diverse play diverse play diverse play diverse
+Groundtruth $\mathbf{94.4}$$\pm 6.3$ $\mathbf{71.0}$$\pm 12.3$ $50.2$$\pm 39.9$ $\mathbf{46.2}$$\pm 23.2$ $\mathbf{58.6}$$\pm 9.1$ $\mathbf{60.2}$$\pm 18.3$ $\mathbf{25.8}$$\pm 18.2$ $\mathbf{55.8}$$\pm 18.3$ $\mathbf{461.8}$
+Learned $66.8$$\pm 5.2$ $44.0$$\pm 22.3$ $\mathbf{54.0}$$\pm 31.4$ $2.0$$\pm 4.0$ $29.4$$\pm 8.1$ $0.0$ ±$\pm 0.0$ $15.6$$\pm 5.9$ $20.8$$\pm 5.2$ $232.6$
+Ablation study on horizon length in V-D4RL.
+Table 23 shows that when we use the default hyperparameter of DreamerV3, $H=15$, the performance drops in V-D4RL.
+The result suggests that we need to use a shorter imagination horizon for offline model-based RL.
+Table 23: V-D4RL ablation results for horizon length. Results are averaged over $3$ random seeds. $H=5$ generally improves the performance of LEQ.
+Dataset H=5 H=15
+walker_walk-random $\mathbf{23.2}$$\pm 1.1$ $15.4$$\pm 1.5$
+walker_walk-medium $\mathbf{50.0}$$\pm 3.6$ $40.0$$\pm 3.6$
+walker_walk-medium_replay $\mathbf{44.0}$$\pm 28.5$ $\mathbf{45.4}$$\pm 7.5$
+walker_walk-medium_expert $\mathbf{90.3}$$\pm 1.8$ $79.9$$\pm 1.4$
+cheetah_run-random $\mathbf{15.3}$$\pm 1.7$ $\mathbf{15.5}$$\pm 3.0$
+cheetah_run-medium $\mathbf{40.1}$$\pm 14.6$ $32.2$$\pm 4.3$
+cheetah_run-medium_replay $\mathbf{39.9}$$\pm 2.0$ $36.4$$\pm 2.1$
+cheetah_run-medium_expert $27.7$$\pm 12.6$ $\mathbf{29.3}$$\pm 4.1$
+Total $\mathbf{330.5}$ $294.1$
+Appendix E Video Predictions on V-D4RL
+Figure 9 presents the imagined trajectory generated by the action sequence during evaluation in the V-D4RL datasets: walker_walk-medium_replay, walker_walk-medium_expert, cheetah_run-medium_replay, cheetah_run-medium_expert, from the top to the bottom.
+Overall, world models trained in medium_replay datasets show better prediction compared to medium_expert dataset, likely due to their broader state distribution of the dataset.
+Nevertheless, LEQ achieves high performance on the walker_walk-medium_expert dataset, despite inaccurate predictions, highlighting robustness of LEQ in handling imperfect world models.
+(a) walker_walk-medium_replay
+(b) walker_walk-medium_expert
+(c) cheetah_run-medium_replay
+(d) cheetah_run-medium_expert
+Figure 9: Video prediction results in V-D4RL.
+The world model receives initial $5$ frames and simulates $64$ additional frames based on the action sequence during evaluation.
+For each trajectory, each row displays the true image, model prediction, and difference between these two, respectively.